### v1

# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

Abstract
Performance analysis of first-order algorithms with inexact oracle has gained recent attention due to various emerging applications in which obtaining exact gradients is impossible or computationally expensive. Previous research has demonstrated that the performance of accelerated first-order methods is more sensitive to gradient errors compared with non-accelerated ones. This paper investigates the nonasymptotic convergence bound of two accelerated methods with inexact gradients to solve deterministic smooth convex problems. Performance Estimation Problem (PEP) is used as the primary tool to analyze the convergence bounds of the underlying algorithms. By finding an analytical solution to PEP, we derive novel convergence bounds for Inexact Optimized Gradient Method (OGM) [39] and Inexact Fast Gradient Method (FGM) [16] with variable inexactness along iterations. Under the absolute error assumption, we derive bounds in which the accumulated errors are independent of the initial conditions and the trajectory of the sequences generated by the algorithms. Furthermore, we analyze the tradeoff between the convergence rate and accumulated error that guides finding the optimal stepsize. Finally, we determine the optimal strategy to set the gradient inexactness along iterations (if possible), ensuring that the accumulated error remains subordinate to the convergence rate.
1 Introduction
We consider the optimization problem
[…]
where $f$ is convex and has a Lipschitz continuous gradient. We assume the optimal value of (1) to be lower bounded, i.e., $f^{*}>-\infty$ with $f^{*}\triangleq\min_{\mathbf{x}}f(\mathbf{x})$ and $\mathcal{X}_{*}\triangleq\operatorname*{argmin}_{\mathbf{x}}f(\mathbf{x})$. We use $\mathcal{F}_{\mu,L}$ with $\mu\geq 0$ to denote the class of $\mu$ strongly convex functions with Lipschitz continuous gradient with constant $L$. This paper focuses on the class of merely convex functions with Lipschitz continuous gradient $f\in\mathcal{F}_{0,L}$. For this class of functions, the gradient descent (GD) method has the convergence rate $f(\mathbf{x}_{k})-f^{*}=\mathcal{O}(k^{-1})$. This rate can be improved to $\mathcal{O}(k^{-2})$ with Nesterovs’s fast gradient method (FGM) [51, 52, 50, 53]. Recently, using computer-aided analysis, the Optimized Gradient Method (OGM) has been proposed [22, 39] that improves the convergence rate of FGM by $\sqrt{2}$ constant.
The above rates are based on having access to the exact gradient at any point, which is not the case in many applications. The convergence rate deteriorates when there is an error in the gradient estimate. This paper considers gradient error in the absolute form as
[…]
where the error $\mathbf{e}$ can be controlled by a parameter $\boldsymbol{\eta}\in\mathbb{R}^{r}$ through a positive function $h(\cdot)$ and the bound holds either deterministically, which is the focus of this work, or with high probability in stochastic settings. The parameter $\boldsymbol{\eta}$ represents, e.g., sampling size, computational effort, or accuracy of an approximation to achieve a desired gradient inexactness.
Below, the gradient oracle inexactness is illustrated through a few examples.
1.1 Examples of inexact gradient oracles
Gradient inexactness is generally the result of the oracle not being evaluated at the desired point. For instance, if some or all coordinates of the desired point are the solution of an optimization problem that is only solved to a suboptimal solution, the resulting gradient will be inexact, e.g., bilevel optimization. In some other scenarios, the desired point results from an expectation that can only be estimated resulting in gradient inexactness see, e.g., nested composition optimization.
Example 1.1 (Bilevel optimization [29]).
Consider the bilevel optimization problem
$\min_{\mathbf{x}}$ $\quad f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
$\operatorname{s.t.}$ $\quad\mathbf{y}^{*}(\mathbf{x})=\operatorname*{argmin}_{\mathbf{y}}g(\mathbf{x},\mathbf{y}),$
with $f$ being continuously differentiable and $g$ being two times continuously differentiable and strongly convex functions. To solve the problem with a gradient-based method, we first obtain the gradient with respect to $\mathbf{x}$ as
$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$ $=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
$=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla\mathbf{y}^{*}(\mathbf{x})^{\top}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x})).$
This requires solving the lower-level problem to optimality, which is unattainable in many scenarios. When the lower-level problem is solved to a suboptimal point $\tilde{\mathbf{y}}(\mathbf{x})$ and the gradient is evaluated at this point instead of $\mathbf{y}^{*}(\mathbf{x})$, the resulting gradient is inexact:
$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\neq\tilde{\nabla}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))\triangleq\nabla_{\mathbf{x}}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x})).$
Under some continuity assumptions, the gradient error is shown to be bounded as
~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left\|\tilde{\mathbf{y}}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|,$
where $C\triangleq L_{f_{\mathbf{x}}}+\frac{L_{f_{\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}}+C_{f_{\mathbf{y}}}\left(\frac{L_{g_{\mathbf{x}}\mathbf{y}}}{\mu_{g}}+\frac{L_{g_{\mathbf{y}\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}^{2}}\right)$, with $L_{f_{\mathbf{x}}},L_{f_{\mathbf{y}}}$ are the Lipschitz continuity constants of $\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y})$ and $\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})$, respectively; $L_{g_{\mathbf{x}\mathbf{y}}},L_{g_{\mathbf{y}\mathbf{y}}}$ are the Lipschitz constants of $\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}),\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y})$ with respect to (w.r.t.) $\mathbf{y}$, respectively; $C_{f_{\mathbf{y}}}$, $C_{g_{\mathbf{x}\mathbf{y}}}$ are the upper bounds on $\left\|\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})\right\|$ and $\left\|\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y})\right\|$, respectively; and $\mu_{g}$ is the strong convexity constant of $g(\mathbf{x},\mathbf{y})$ w.r.t. $\mathbf{y}$– see [29] for details.
Now let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations. From the convergence rate of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
where $Q_{g}\triangleq L_{g}/\mu_{g}$. The above bound follows an exponential decay and represents the error bound in (2) with $\eta$ being the iteration number $k$.
Example 1.2 (Composition optimization).
Consider the nested composition optimization in the form of
$\min_{\mathbf{x}}\ f(\mathbf{x})\triangleq h(\mathbf{g}(\mathbf{x}))\quad\text{where}\quad\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{N}\mathbf{g}_{i}(\mathbf{x}),$
where $h:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\mathbf{g}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ are continuously differentiable, and $N$ is a large positive integer. From the chain rule, the exact gradient is $\nabla f(\mathbf{x})=\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))$. However, assume one can only approximate $\mathbf{g}$ and $\nabla\mathbf{g}$ through sampling as
$\tilde{\mathbf{g}}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\mathbf{g}}\right|}\mathbf{g}_{\mathcal{S}_{\mathbf{g}}[i]}(\mathbf{x})\quad\tilde{\nabla}\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\nabla}\right|}\nabla\mathbf{g}_{\mathcal{S}_{\nabla}[i]}(\mathbf{x}),$
where $\mathcal{S}_{\mathbf{g}}$, $\mathcal{S}_{\nabla}$ are two subsets of $\{1,2,\cdots,N\}$.
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Furthermore, assume $h(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$, and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Then with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right).$
The proof of the above bound is presented in Lemma A.1. This bound represents the error bound in (2) with $\boldsymbol{\eta}=(\left|\mathcal{S}_{\nabla}\right|,\left|\mathcal{S}_{\mathbf{g}}\right|)$. Furthermore, if $\mathcal{S}_{\nabla}=\mathcal{S}_{\mathbf{g}}$, for any fixed $\epsilon$, the bound follows a power law decay.
Another common scenario is when $\nabla f$ is unattainable, but, they can be approximated by (inexact) function values, i.e., zero-order information, denoted by $\tilde{f}$. In many such scenarios, it can be assumed that $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq b_{f}$. When $f(\mathbf{x})$ has Lipschitz continuous gradient, the zero-order methods generate inexact gradient estimates with quantifiable error bounds (see [6] and references therein).
Example 1.3 (Gradient estimation via forward finite difference).
Let $\vec{\mathbf{v}}_{i}\in\mathbb{R}^{d}$ denote the unit vector with the i-th element equal to 1, and $l>0$ be the finite difference interval. Define the i-th element of the gradient estimator as
[~$[\tilde{\nabla}f(\mathbf{x})]_{i}=\frac{\tilde{f}(\mathbf{x}+l\vec{\mathbf{v}}_{i})-\tilde{f}(\mathbf{x})}{l}.$
The error of the gradient estimate is shown by [6] to be bounded as
~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\frac{\sqrt{d}Ll}{2}+\frac{2\sqrt{d}b_{f}}{l},$
which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$. Note that the forward finite difference requires querying the zero-order oracle $d+1$ times to approximate the gradient at $\mathbf{x}$.
Another common approach for gradient approximation from (inexact) function values is based on Gaussian smoothing. This method randomly samples directions for gradient estimation; hence, the sample size can be some arbitrary number.
Example 1.4 (Gradient estimation via Gaussian smoothing).
Let $\{\mathbf{v}_{i}\}_{i=1}^{n}$ be i.i.d. random directions following the standard multivariate normal distribution, and define the gradient estimator as
~$\tilde{\nabla}f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{f}(\mathbf{x}+l\mathbf{v}_{i})-\tilde{f}(\mathbf{x})}{l}\mathbf{v}_{i}.$
It is shown in [6] that
$\left\|\mathbb{E}_{\{\mathbf{v}_{i}\}_{i=1}^{n}}[\tilde{\nabla}f(\mathbf{x})]-\nabla f(\mathbf{x})\right\|\leq\sqrt{d}Ll+\frac{\sqrt{d}b_{f}}{l},$
which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$. Furthermore,
~$\left\|\tilde{\nabla}f(\mathbf{x})-\mathbb{E}_{\{\mathbf{v}_{i}\}_{i=1}^{n}}[\tilde{\nabla}f(\mathbf{x})]\right\|^{2}\leq\frac{3d}{\epsilon n}\left(3\left\|\nabla f(\mathbf{x})\right\|^{2}+\frac{L^{2}l^{2}}{4}(d+2)(d+4)+\frac{4b_{f}^{2}}{l^{2}}\right)$
holds with probability at least $1-\frac{1}{\epsilon}$. In the later probabilistic bound, $h(\boldsymbol{\eta})$ is determined by $l$, $n$ and $b_{f}$ as $\boldsymbol{\eta}=(l,b_{f},n)$. This bound contains a constant, a relative error, and an absolute error terms.
1.2 Related works
The study on inexact gradient methods (without acceleration) dates back to 1987 [56], where two notions of inexactness were defined, i.e.,
~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|$ $\leq\delta,\qquad$ (absolute error), (3)
[…]
Below, we review the literature in stochastic setups for non-accelerated methods and (in more detail) in deterministic setups for accelerated gradient-based methods.
Stochastic non-accelerated methods under biased oracles.
The above gradient inexactness assumptions have specifically been considered in stochastic setups where ~$\tilde{\nabla}f(\mathbf{x})$ is a biased stochastic estimate of $\nabla f(\mathbf{x})$. [1] studies the performance of the biased stochastic gradient descent (SGD) method under both absolute and relative bias. [36] considers setups where lower bias and variance require higher computational cost and investigates their tradeoff in stochastic optimization. Similarly, [35] analyzes bias-variance tradeoff in the convex bandit setting. [7] analyzes biased stochastic gradient-based methods under the absolute bias oracles with and without bias-variance tradeoff. [14] (and references therein) summarizes different bias assumptions and introduces a hybrid approach integrating absolute and relative biases. [47] considers biased gradient estimates with absolute error and proposes adaptive algorithms that adaptively set the bias throughout the iterations of the algorithms.
Accelerated gradient methods with inexact gradient oracles.
Nesterov-type acceleration techniques [51] with inexact gradient oracles have been analyzed under different inexact assumptions. Below, these assumptions and their connections with the absolute error (which is the focus of this paper) are reviewed; moreover, the convergence performance of deterministic accelerated methods with inexact gradients are analyzed under the corresponding assumptions.
An inexact generalization of Nesterov’s Fast Gradient Method (FGM), namely, iFGM, is presented in Algorithm 1 which obtains accelerated rates of convergence in convex setups. To simplify the exposition, we only consider the merely convex problems and we set the dual update step as gradient descent with projection.
Algorithm 1 Inexact Fast Gradient Method (iFGM)
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $0<\alpha_{0}\leq 1$, $A_{0}=\alpha_{0}$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k}))$
3:     $\mathbf{z}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}^{2}\leq A_{k+1}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for
d’Aspremont [11] analyzes the performance of iFGM to minimize Liptshitz smooth objective function over compact convex sets $\mathcal{Q}$, with the inexact gradient satisfying
|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\delta,\quad\forall\mathbf{x},\mathbf{y},\mathbf{z}\in\mathcal{Q},$
and shows the sequence generated by iFGM satisfies
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{A_{K}}+3\delta.$
This inexactness assumption can be derived from the absolute error assumption for the bounded feasible region with diameter $D$ as
|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\bar{b}\left\|\mathbf{y}-\mathbf{z}\right\|\leq\bar{b}\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|=\bar{b}D\triangleq\delta,$
which shows the dependence of the error term on the diameter of the feasible region.
The notion of first-order $(\delta,L)$-oracle is proposed in [16], which removes the requirements of boundedness of the feasible set and differentiability of the objective function with a generalized definition of convexity and Lipschitz smoothness. More specifically, [16] assumes that the oracle provides an inexact function and subgradient pair $(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x}),\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}))$ that satisfies
$0\leq f(\mathbf{y})-(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}+\delta_{\mathbf{x}}.$
With $(\delta,L)$-oracle, [16] shows the sequence generated by iFGM satisfies
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{A_{K}}+\frac{\sum_{i=0}^{K}A_{i}\delta_{\mathbf{x}_{i}}}{A_{K}}.$
This result allows different inexactness levels for different iterations. However, to use the above results for an oracle that satisfies the absolute error assumption, the boundedness of the feasible region is needed [16], i.e., an oracles with absolute error assumption $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq\Delta_{1}$ and ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\Delta_{2}$, $\tilde{f}_{\delta_{\mathbf{x}}}=\tilde{f}(\mathbf{x})-\Delta_{1}-\Delta_{2}D$, ~$\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x})=\tilde{\nabla}f(\mathbf{x})$, and $\delta_{\mathbf{x}}=2\Delta_{1}+2\Delta_{2}D$ (with $D$ being the feasible region diameter) is a $(\delta,L)-$oracle.
To overcome the boundedness drawback of the $(\delta,L)$-oracle, [49] proposes the inexact first-order oracle of degree $q$ as
$0\leq f(\mathbf{y})-(f(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{y}-\mathbf{x}\right\|^{2}+\delta\left\|\mathbf{x}-\mathbf{y}\right\|^{q},$
with $q\in[0,2)$. With the appropriate selection of the $q$ and $\delta$ parameters, any inexact gradient that satisfies the absolute error condition satisfies the right-hand-side inequality above without the requirement of the bounded feasible region. Under this new oracle assumption, the sequence generated by iFGM satisfies
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4(L+q\rho)\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{(K+3)(2-q)\delta^{\frac{2}{2-q}}}{2\rho^{\frac{q}{2-q}}},$
with $\rho>0$ be any fixed number. Minimizing the right-hand-side with respect to $\rho$, the convergence rate becomes
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{8^{\frac{q}{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{q}(K+3)}{((K+1)(K+2)(K+3))^{\frac{q}{2}}}\delta.$
This latter condition allows unbounded regions; however, its left-hand-side inequality requires the inexact gradient to be a subgradient, which can not be guaranteed under the absolute error assumption. Furthermore, in the convergence bound, the error term depends on the initial condition. Finally, the analysis is provided for fixed inexactness levels.
Another accelerated algorithm studied under inexact oracles is the Similar Triangles Method (STM) [27]– see the inexact version in Algorithm 2. Compared to FGM, it only requires one projection for each iteration and, hence, it is preferable for constrained problems. The convergence properties of iSTM are studied under both absolute and relative errors.
Algorithm 2 Inexact Similar Triangles Method (iSTM)
0:  $\mathbf{z}_{-1}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $\alpha_{-1}=A_{-1}=0$.
1:  for $k=0,\cdots,K$ do
2:     $A_{k}=A_{k}+\alpha_{k}$ with $\alpha_{k}^{2}\leq A_{k}$
3:     $\mathbf{x}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k-1})$
4:     $\mathbf{z}_{k}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
5:     $\mathbf{y}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k})$
6:  end for
When the feasible region $\mathcal{Q}=\mathbb{R}^{d}$, i.e., no projection is required, iSTM can be formulated as iFGM. From $A_{k}=A_{k-1}+\alpha_{k}$, we can see that the $\mathbf{x}_{k}$ updates in iFGM and iSTM are the same with different indices since iFGM updates $\mathbf{x}_{k}$ at the end of each iteration while iSTM updates it at the beginning of each iteration. The $\mathbf{z}_{k}$ updates for the two algorithms are the same. Hence, we only need to consider the $\mathbf{y}_{k}$ updates. From step 3 of iSTM, $A_{k-1}\mathbf{y}_{k-1}=A_{k}\mathbf{x}_{k}-\alpha_{k}\mathbf{z}_{k-1}$. Inserting this identity into step 5, and by the update of step 4, we have
$\mathbf{y}_{k}=\mathbf{x}_{k}+\frac{\alpha_{k}}{A_{k}}(\mathbf{z}_{k}-\mathbf{z}_{k-1})=\mathbf{x}_{k}-\frac{\alpha_{k}^{2}}{A_{k}L}\tilde{\nabla}f(\mathbf{x}_{k}).$
When $\alpha_{k}^{2}=A_{k}$, the above update is equivalent to the step 2 in iFGM.
Furthermore, [2] analyzes iSTM under the inexact gradient with absolute error bound. The sequence generated by iSTM with $\alpha_{k}^{2}=A_{k},\forall k$, satisfies
$f(\mathbf{y}_{K})-f_{*}\leq\frac{8L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{K^{2}}+3\tilde{R}_{K}\delta+\frac{K\delta^{2}}{2L},$
with $\tilde{R}_{K}\triangleq\max_{0\leq j\leq K}\{\left\|\mathbf{z}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{x}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{y}_{j}-\mathbf{x}_{*}\right\|\}$. This convergence bound is only given under fixed inexactness levels and restricts the stepsize as $\alpha_{k}^{2}=A_{k}$. Another drawback is that $\tilde{R}_{K}$ is not exactly quantifiable; indeed, [2] only shows its boundedness by $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ under the exact gradient oracle.
The boundedness of $\tilde{R}_{K}$ is shown by the same group under the relative error assumption with a relaxed stepsize [44]. It is shown that the sequence generated by iSTM with $\alpha_{k}=\frac{(k+2)^{p-1}}{2\alpha}$, $p\in[1,2]$, and $\alpha=\mathcal{O}(\max\{1,K^{\frac{p}{4}}\sqrt{\delta},K^{\frac{p}{2}}\delta,K^{p}\delta^{2}\})$ satisfies
$f(\mathbf{y}_{K})-f_{*}\leq\frac{16\alpha L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+2)^{p}}.$
The paper shows $\tilde{R}_{K}$ is bounded by $2\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ when the stepsize is properly chosen and, hence, there is no separate error term in the convergence rate.
Besides the abovementioned works, several researchers have proposed alternative error assumptions and results. The result of iFGM for the strongly convex setup is provided in [17]. [8] considers a variation of iFGM with the absolute error setting. To exploit the tradeoff between the convergence rate and accumulated error, the intermediate gradient method is proposed in [18] which is closely related to iFGM. This method is further generalized for other setups in [25, 19, 60, 28, 61].
[26] and [3] analyze nonaccelerated and accelerated algorithms under relative and absolute errors, respectively, with the latter focusing on the tradeoff between robustness to error and convergence rate–see also [32]. [24, 45] assume the error to satisfy the ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq B\left\|\mathbf{x}-\mathbf{x}_{*}\right\|$ condition, which can be derived from the relative error bound. [24] provides results for stochastic proximal inexact gradient methods with and without acceleration for strongly convex problems, while [45] studies generalized monotone variational inequality problems. Other notable contributions include the analysis of absolute and relative inaccuracy for proximal point methods in [57], the study of proximal gradient methods with absolute error in [59], the inexact proximal gradient method for weakly convex functions with absolute error in [37], and minimization of the accumulated error when the total allocation of error-control resources is fixed [15].
1.3 Contributions
This paper employs Performance Estimation Problem (PEP) techniques to primarily analyze the convergence rate of the Generalized Optimized Gradient Method with an inexact gradient oracle (iGOGM). Our key contributions are as follows:
  • We derive an analytical solution to the relaxed PEP optimization problem for iGOGM under variable oracle inexactness levels. The approach used to find this solution inspires our theoretical proof of the convergence. Notably, this proof can be understood without prior knowledge of PEP techniques.
  • The derived convergence bound consists of two components: the exact convergence rate and the accumulated error. Significantly, the error term is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$, and is determined solely by the Lipschitz constant and the step size.
  • We provide a comprehensive analysis of the convergence bound from two perspectives: i) We investigate the tradeoff between the convergence rate and the accumulated error. ii) We establish optimal inexactness levels that preserve the method’s convergence rate.
  • We also provide the PEP-inspired convergence bound for the Generalized Fast Gradient Method with inexact gradient oracle (iGFGM).
We also provide the codes for our numerical experiments as well as the verification of the algebraic steps in our proofs. These codes are available at https://github.com/Yin-LIU/Inexact-Acceleration-PEP.
Notations
[…]
2 Performance Estimation Problems (PEP)
2.1 PEP overview
The tool used for analyzing the convergence rate of the inexact accelerated methods is known as Performance Estimation Problem (PEP). This technique is proposed by Drori and Teboulle [22] which uses an optimization problem to find the worst-case performance when a given algorithm is applied to a class of functions.
To solve the convex and Lipschitz smooth problems with the general first-order method using an exact gradient oracle, the PEP is defined as
$\max_{\begin{subarray}{c}d,f\\ \mathbf{x}_{*},\mathbf{x}_{0},\cdots,\mathbf{x}_{k}\in\mathbb{R}^{d}\\ \nabla f(\mathbf{x}_{0}),\cdots,\nabla f(\mathbf{x}_{K})\in\mathbb{R}^{d}\end{subarray}}\quad$ $f(\mathbf{x}_{K})-f(\mathbf{x}_{*})$
$\operatorname{s.t.}\qquad\quad$ $f\in\mathcal{F}_{0,L},\quad\mathbf{x}_{*}\in\operatorname*{argmin}_{\mathbf{x}}f(\mathbf{x})$
$\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}\leq R^{2}$
$\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\nabla f(\mathbf{x}_{i})\quad k=1,\cdots,K,$
where $\mathcal{F}_{0,L}$ denotes the class of merely convex functions with Lipschitz continuous gradients. The general first-order algorithm (GFO) performs the update
$\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\nabla f(\mathbf{x}_{i})\quad k=1,\cdots,K,$
which covers a wide range of first-order methods, including GD, FGM, Polyak’s Heavy Ball Method [55], and Optimized Gradient Method (see Algorithms 4 and 3[22].
The above problem is an infinite dimensional optimization as the variable $f$ is a function. [22] proposed a notion of $\mathcal{F}_{\mu,L}$-interpolation that only considers function values and gradients at discrete points $\{\mathbf{x}_{k}\}$, which are the terms that play a role in the optimization problem.
Definition 2.1 ($\mathcal{F}_{\mu,L}$-interpolation (Definition 2 in [63])).
Let $\mathcal{I}$ be an index set, and consider the set of triples $\mathcal{S}=\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in\mathcal{I}}$ where $\mathbf{x}_{i},\mathbf{g}_{i}\in\mathbb{R}^{d}$ for all $i\in\mathcal{I}$. Set $\mathcal{S}$ is $\mathcal{F}_{\mu,L}$-interpolatable if and only if there exists a function $f\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ such that we have both $\mathbf{g}_{i}=\nabla f(\mathbf{x}_{i})$ and $f(x_{i})=f_{i}$ for all $i\in\mathcal{I}$.
Theorem 2.2 ($\mathcal{F}_{\mu,L}$-interpretability (Theorem 4 in [63]) ).
Set $\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in I}$ is $\mathcal{F}_{\mu,L}$-interpolable ($0\leq\mu\leq L\leq\infty$) if and only if the following set of conditions holds for every pair of indices $i\in I$ and $j\in I$:
$f_{i}-f_{j}-\mathbf{g}_{j}^{\top}(\mathbf{x}_{i}-\mathbf{x}_{j})\geq\frac{1}{2(1-\mu/L)}\left(\frac{1}{L}\|\mathbf{g}_{i}-\mathbf{g}_{j}\|^{2}+\mu\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}-\frac{2\mu}{L}(\mathbf{g}_{j}-\mathbf{g}_{i})^{\top}(\mathbf{x}_{j}-\mathbf{x}_{i})\right).$ (5)
The corresponding discrete PEP model is
[…]
$\operatorname{s.t.}\quad$ $\mathcal{S}=\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in\{*,0,1,\cdots,K\}}$
$\mathbf{g}_{*}=0,\quad\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|\leq R^{2}$
$f_{i}-f_{j}-\mathbf{g}_{j}^{\top}(\mathbf{x}_{i}-\mathbf{x}_{j})\geq\frac{1}{2L}\left\|\mathbf{g}_{i}-\mathbf{g}_{j}\right\|^{2}\quad\forall i,j\in\{*,0,1,\cdots,K\}$
$\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\mathbf{g}_{k}\quad k=1,\cdots,K.$
The theoretical equivalence of the discrete PEP model and the worst-case performance of GFO over the given class of functions is proven in [64, 63] with convex interpolation definition that guarantees PEP to generate the tight worst-case performance.
The PEP model can also be used to find algorithms that minimize the worst-case function. In Drori and Teboulle [22], numerical results indicate a better algorithm than the well-known FGM. [39] follows this idea and provides an explicit form of the generated algorithm called the Optimized Gradient Method (OGM), which was proven to match the lower bound with a smaller $\mathcal{O}(1)$ constant [20]. [21] further reveals the equivalence between OGM and the conjugate gradient method–see also [40, 42, 43, 54].
PEP technique has been successfully applied to various optimization methods, including gradient descent [31], algorithms with line search [12, 21], proximal gradient methods [41], proximal point algorithms [38], algorithms with inexact oracles [13, 5], and many others [4, 58, 23]. Besides the deterministic smooth regimes, PEP has also been incorporated for stochastic problems [62] and for problems satisfying relative smoothness [65].
When optimizing only over the function class given an algorithm, PEP can be equivalently written as a Semidefinite Programming (SDP) and solved efficiently. However, when optimizing over both algorithm and function, the problem is nonconvex, and finding its global solution is NP-hard. Hence, to find an optimal algorithm a relaxed PEP formulation is generally solved which may not provide an optimal algorithm. Das Gupta et al. [10] proposes a method to solve the original nonconvex PEP and provides a numerical guide to design algorithms.
A closely related topic is the analysis based on Integral Quadratic Constraints (IQC), originally studied in control theory [48]. Similar to PEP, IQC aims to find worst-case functions through optimization, but it is not an exact formulation of the function class; hence, it can only provide an upper bound. This technique has been used to analyze first-order methods in [46, 33, 65] and gradient method with inexact oracles in [34, 9].
2.2 PEP for optimization with inexact oracle
For inexact gradient oracles, the GFO is generalized to inexact GFO (iGFO) as
[…]
and the discrete PEP model is adjusted as
[…]
$\operatorname{s.t.}\quad$ $\mathcal{S}=\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in\{*,0,1,\cdots,K\}}$
$\mathbf{g}_{*}=0,\quad\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}\leq R^{2}$
$f_{i}-f_{j}-\mathbf{g}_{j}^{\top}(\mathbf{x}_{i}-\mathbf{x}_{j})\geq\frac{1}{2L}\left\|\mathbf{g}_{i}-\mathbf{g}_{j}\right\|^{2}\quad\forall i,j\in\{*,0,1,\cdots,K\}$
$\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}(\mathbf{g}_{k}+\mathbf{e}_{k})\quad k=1,\cdots,K$
$\left\|\mathbf{e}_{k}\right\|^{2}\leq b_{k}^{2}\quad k=0,\cdots K-1,$
which is the generalization of the exact model. If $b_{k}\equiv 0$ then $\mathbf{e}_{k}\equiv\mathbf{0}$, iGFO reduces to GFO, and (PEP) reduces to (PEP-Exact).
This formulation can be further simplified by introducing a Gram matrix. Define
[…]
$\mathbf{X}$ $\triangleq\begin{bmatrix}\mathbf{x}_{0}&\mathbf{e}_{0}&\cdots&\mathbf{e}_{K-1}&\mathbf{g}_{0}&\cdots&\mathbf{g}_{K}\end{bmatrix}\in\mathbb{R}^{d\times(2K+2)},$ (8)
and the Gram matrix $\mathbf{G}\triangleq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{S}^{2K+2}$ which is a symmetric positive semidefinite (PSD) matrix. Also, we define standard basis vectors $\vec{\mathbf{f}}_{k}\in\mathbb{R}^{K+1}$ and $\vec{\mathbf{x}}_{k},\vec{\mathbf{g}}_{k},\vec{\mathbf{e}}_{k}\in\mathbb{R}^{2K+2}$ to select corresponding columns in $\mathbf{F}$ or $\mathbf{X}$ to recover the needed element, which satisfy
$f_{k}-f_{*}=\mathbf{F}\vec{\mathbf{f}}_{k},\quad\vec{\mathbf{g}}_{*}=\mathbf{0},\quad\mathbf{g}_{k}=\mathbf{X}\vec{\mathbf{g}}_{k},\quad\mathbf{e}_{k}=\mathbf{X}\vec{\mathbf{e}}_{k},$
$\mathbf{x}_{0}=\mathbf{X}\vec{\mathbf{x}}_{0},\quad\mathbf{x}_{k}=\mathbf{X}\vec{\mathbf{x}}_{k}=\mathbf{X}\left(\vec{\mathbf{x}}_{0}-\frac{1}{L}\sum_{i=0}^{k-1}\theta_{k,i}(\vec{\mathbf{g}}_{i}+\vec{\mathbf{e}}_{i})\right).$
Note. We note that the subscript $k$ on the standard basis vectors does not mean that the $k$-th element of that vector is equal to 1. For instance, based on the definitions of $\mathbf{F}$ and $\mathbf{X}$ in (7), $\vec{\mathbf{f}}_{k}$ is a standard basis vector with $k+1$-th element equal to 1, $\vec{\mathbf{e}}_{k}$ is a standard basis vector with $k+2$-th element equal to 1, $\vec{\mathbf{g}}_{k}$ is a standard basis vector with $K+k+2$-th element equal to 1, and $\vec{\mathbf{x}}_{0}$ is a standard basis vector with the first element equal to 1. With a slight abuse of the notation, $\vec{\mathbf{x}}_{k}$ (which is not a standard basis vector) is the vector used to represent $\mathbf{x}_{k}$ by a linear combination of the columns of $\mathbf{X}$ and it is a function of $\theta_{k,i}$.
Without loss of generality, we assume $f_{*}=0$ and $\mathbf{x}_{*}=0$, hence $\vec{\mathbf{f}}_{*}=\mathbf{0},\vec{\mathbf{x}}_{*}=\mathbf{0}$. With this new notation, the PEP problem is written as
[…]
$\operatorname{s.t.}\quad$ $\mathbf{F}(\vec{\mathbf{f}}_{j}-\vec{\mathbf{f}}_{i})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{ij})\leq 0,\quad\forall i,j\in\{*,0,1,\cdots,K\}$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top})-R^{2}\leq 0$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top})-b_{i}^{2}\leq 0,\quad i=0,\cdots,K-1$
$\operatorname{rank}(\mathbf{G})\leq d,$
where
$\mathbf{A}^{ij}\triangleq\frac{1}{2}((\vec{\mathbf{x}}_{i}-\vec{\mathbf{x}}_{j})\vec{\mathbf{g}}_{j}^{\top}+\vec{\mathbf{g}}_{j}(\vec{\mathbf{x}}_{i}-\vec{\mathbf{x}}_{j})^{\top})+\frac{1}{2L}(\vec{\mathbf{g}}_{i}-\vec{\mathbf{g}}_{j})(\vec{\mathbf{g}_{i}}-\vec{\mathbf{g}}_{j})^{\top}.$
When dealing with large-scale problems, i.e., when $2K+2\leq d$, we can drop the rank constraint $\operatorname{rank}{(G)}\leq d$ without changing the optimal value (see Theorem 5 in [64]), and the equivalent problem will be a Semidefinite Programming problem (SDP) written as
[…]
$\operatorname{s.t.}\quad$ $\mathbf{F}(\vec{\mathbf{f}}_{j}-\vec{\mathbf{f}}_{i})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{ij})\leq 0,\quad\forall i,j\in\{*,0,1,\cdots,K\}$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top})-R^{2}\leq 0$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top})-b_{i}^{2}\leq 0,\quad i=0,\cdots,K-1.$
The optimal solution of (SDP-PEP) will provide an exact worst-case performance of a given algorithm for convex problems. It is also possible to find the optimal algorithm with parameter ${\boldsymbol{\theta}}$ if we consider the minimax problem as
$\min_{\theta_{k,i}}\max_{\mathbf{G},\mathbf{F}\in\mathcal{S}(\text{SDP-PEP})}\mathbf{F}\vec{\mathbf{f}}_{K},$ (9)
[…]
Recall that for FGM, there are two sequences, the primary sequence $\{\mathbf{y}_{k}\}$ and the secondary one $\{\mathbf{x}_{k}\}$. The OGM is optimized for minimizing $f(\mathbf{x}_{k})-f_{*}$ while the commonly used measure is $f(\mathbf{y}_{K})-f_{*}$. Due to this fact, [40] propose a new measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ since for smooth function and with OGM update rule, $f(\mathbf{y}_{K+1})=f(\mathbf{x}_{K}-\frac{1}{L}\nabla f(\mathbf{x}_{K}))\leq f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$. The solution of this new measure results in OGM (Algorithm 4). These two algorithms only differ in the last iteration’s $\alpha_{K}$ update. For simplicity, we only focus on this measure and its corresponding relaxed PEP.
Algorithm 3 OGM with last adjustment
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
4:     $\alpha_{k+1}=\begin{cases}\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2},k\leq K-2\\ \frac{1+\sqrt{1+8\alpha_{k}^{2}}}{2},k=K-1\end{cases}$
5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
6:  end for
Algorithm 4 OGM
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
4:     $\alpha_{k+1}=\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
6:  end for
The relaxed optimization problem is
[…]
$\operatorname{s.t.}\quad$ $\mathbf{F}(\vec{\mathbf{f}}_{i+1}-\vec{\mathbf{f}}_{i})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{i,i+1})\leq 0,\quad i=0,\cdots K-1,$ $\qquad(v_{i,i+1})$
$\mathbf{F}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{*})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{*,i})\leq 0,\quad i=0,\cdots K,\qquad$ $(v_{*,i})$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top})-R^{2}\leq 0,$ $\qquad(\tau)$
$\operatorname{Tr}(\mathbf{G}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top})-b_{i}^{2}\leq 0,\quad i=0,\cdots,K-1.$ $(u_{i})$
One can solve this problem by solving its dual problem. To write the dual of (P), we define the dual variables as $\mathbf{v},\mathbf{v}_{*},\tau$, and $\mathbf{u}$ as indicated in the parenthesis following each constraint. The Lagrangian function is
$\mathcal{L}(\mathbf{G},\mathbf{F},\mathbf{v},\mathbf{v}_{*},\mathbf{u},\tau)$
$=$ $\mathbf{F}\vec{\mathbf{f}}_{K}-\frac{1}{2L}\mathbf{G}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}-\sum_{i=0}^{K-1}v_{i,i+1}\left(\mathbf{F}(\vec{\mathbf{f}}_{i+1}-\vec{\mathbf{f}}_{i})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{i,i+1})\right)-\sum_{i=0}^{K}v_{*,i}\left(\mathbf{F}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{*})+\operatorname{Tr}(\mathbf{G}\mathbf{A}^{*,i})\right)$
$-\tau\left(\operatorname{Tr}(\mathbf{G}\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top})-R^{2}\right)-\sum_{i=0}^{K-1}u_{i}\left(\operatorname{Tr}(\mathbf{G}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top})-b_{i}^{2}\right)$
$=$ $\tau R^{2}+\sum_{k=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
$-\operatorname{Tr}\left(\mathbf{G}\left(\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K-1}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\right)\right).$
For the problem $\max_{\mathbf{G}\in\mathbb{S}_{+},\mathbf{F}}\mathcal{L}(\mathbf{G},\mathbf{F},\mathbf{v},\mathbf{v}_{*},\mathbf{u},\tau)$ to have a bounded solution, we can write the corresponding dual problem as
[…]
$\operatorname{s.t.}\quad$ $\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}_{i}}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}v_{*,i}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})=0$
$\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\succeq 0.$
Here we omit the process of relaxing the minimax problem and use the solution in [42] directly to explain the property of this feasible solution with the inexact oracle. Define $A_{k}=\sum_{i=0}^{k}\alpha_{i}$ and $A_{0}=1,A_{-1}=0$, and assume $\alpha_{k}^{2}\leq A_{k}$, the feasible solution of the minimax problem is
[…]
and
[…]
Remark 2.3.
The above feasible solution works for both OGM and Generalized OGM (discussed in the next section), which means for any stepsize $\{\alpha_{k}\}$ and $\{A_{k}\}$ that satisfy $\alpha_{k}^{2}\leq A_{k}$, (10) and (11) satisfies the constraints of (D).
It is easy to verify that the equality constraint holds with this solution. We still need to find the feasible solution for $\mathbf{u}$ to satisfy the positive semidefinite (PSD) constraint. After substitution by (10) and (11), the left-hand-side matrix of the PSD constraint is
$\mathbf{M}=\begin{pmatrix}[l]\frac{L}{4A_{K}}&[\mathbf{0}]_{1\times K}&[\mathbf{p}]_{1\times(K+1)}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times(K+1)}\\ [\mathbf{p}^{\top}]_{(K+1)\times 1}&[\mathbf{B}^{\top}]_{(K+1)\times K}&[\mathbf{C}]_{(K+1)\times(K+1)}\end{pmatrix},$
where $\mathbf{U}=\operatorname{diag}(u_{0}\cdots u_{K-1})$, $\mathbf{p}_{i}=-\frac{\alpha_{i-1}}{2A_{K}}$,
$\mathbf{B}_{i,j}=\begin{cases}\frac{2\alpha_{i-1}\alpha_{i}+A_{i-1}}{2LA_{K}}\quad&j=i+1\\ \frac{\alpha_{i-1}\alpha_{j-1}}{LA_{K}}\quad&i+2\leq j\leq K+1\end{cases},\quad\text{and}\quad\mathbf{C}_{i,j}=\begin{cases}\frac{A_{i-1}}{LA_{K}}\quad&i=j\\ \frac{\alpha_{i-1}\alpha_{j-1}}{LA_{K}}\quad&i\neq j\\ \end{cases}.$
For $K=3$, one example of $\mathbf{M}$ is listed below
$\mathbf{M}=\left(\begin{array}[]{c|ccc|cccc}\frac{L}{4A_{3}}&0&0&0&-\frac{1}{2A_{3}}&-\frac{{\alpha}_{1}}{2A_{3}}&-\frac{{\alpha}_{2}}{2A_{3}}&-\frac{{\alpha}_{3}}{2A_{3}}\\[2.0pt] \hline\cr 0&{u}_{0}&0&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{1}&0&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&0&{u}_{2}&0&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr-\frac{1}{2A_{3}}&0&0&0&\frac{1}{LA_{3}}&\frac{{\alpha}_{1}}{LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ -\frac{{\alpha}_{1}}{2A_{3}}&\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&\frac{{\alpha}_{1}}{LA_{3}}&\frac{A_{1}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ -\frac{{\alpha}_{2}}{2A_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{2}}{LA_{3}}&\frac{A_{2}}{LA_{3}}&\frac{{\alpha}_{2}{\alpha}_{3}}{LA_{3}}\\ -\frac{{\alpha}_{3}}{2A_{3}}&\frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{2}{\alpha}_{3}}{LA_{3}}&\frac{1}{L}\end{array}\right)$
From the structure of $\mathbf{M}$, we can observe that all the rows of $\mathbf{B}$ have at least one positive element, hence to make $\mathbf{M}\succeq 0$ we must have $\mathbf{u}>\mathbf{0}$. Following the decomposition of [40], i.e., $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$ with $\mathbf{w}_{\tau}$ being the first column of $\mathbf{M}$, with the condition $A_{k}=\sum_{i=0}^{k}\alpha_{i}$, the residual matrix $\mathbf{R}$ has the form
$\mathbf{R}=\begin{pmatrix}[l]0&[\mathbf{0}]_{1\times K}&[\mathbf{0}]_{1\times K+1}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times K+1}\\ [\mathbf{0}]_{K+1\times 1}&[\mathbf{B}^{\top}]_{K+1\times K}&[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]_{K+1\times K+1}\end{pmatrix},$
where $\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}=\operatorname{diag}(0,\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})$ for any $\{A_{k}\}$, and equals to $\mathbf{0}$ with $\{\alpha_{k}\}$ defined in OGM. $\mathbf{R}$ in the example of $K=3$ is equal to
$\mathbf{R}=\left(\begin{array}[]{c|ccc|cccc}0&0&0&0&0&0&0&0\\[2.0pt] \hline\cr 0&{u}_{0}&0&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{1}&0&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&0&{u}_{2}&0&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr 0&0&0&0&0&0&0&0\\ 0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&0&\frac{A_{1}-\alpha_{1}^{2}}{LA_{3}}&0&0\\ 0&\frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&0&0&\frac{A_{2}-\alpha_{2}^{2}}{LA_{3}}&0\\ 0&\frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&0&0&0&\frac{A_{3}-\alpha_{3}^{2}}{LA_{3}}\end{array}\right).$
Recalling that $\mathbf{R}^{\backslash[1]}$ denotes the matrix resulting from dropping the first column and row of $\mathbf{R}$, this submatrix is exactly the Shur complement of the block $[\tau]$ of the matrix $\mathbf{M}$. Since $\tau>0$, we know that $\mathbf{M}\succeq 0$ if and only if $\mathbf{R}^{\backslash[1]}\succeq 0$. In the following lemma, we will show that $\mathbf{u}=\infty$ is the unique feasible solution when $A_{i}-\alpha_{i}^{2}=0$.
Lemma 2.4.
With the given solution in (10) and (11), and $A_{i}-\alpha_{i}^{2}=0$, $\mathbf{u}=\infty$ is the unique solution that makes $\mathbf{R}^{\backslash[1]}\succeq 0$. When $A_{i}-\alpha_{i}^{2}>0$, $\mathbf{u}=\infty$ is still feasible but not unique.
Proof.
First, we know that all $u_{i}$ are positive numbers. Define $\vec{\mathbf{v}}_{k}$ be the unit vector that only the $k$-th element is 1. If $\mathbf{R}^{\backslash[1]}\succeq 0$, then for any $c\in\mathbb{R}$ and defining $\mathbf{v}\triangleq\vec{\mathbf{v}}_{i}+c\vec{\mathbf{v}}_{j}$, the condition
$0\leq$ $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=c^{2}\mathbf{R}^{\backslash[1]}_{j,j}+\mathbf{R}^{\backslash[1]}_{i,i}+2c\mathbf{R}^{\backslash[1]}_{i,j}$
holds. To show that, there are four different scenarios:
  1. 1.
    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=0$,
  2. 2.
    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$,
  3. 3.
    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=0$,
  4. 4.
    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$.
For the first and second scenarios, it is obvious that $\mathbf{R}^{\backslash[1]}_{i,j}=0$, with diagonal elements being non-negative, the condition holds. For the third scenario, $\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$. Consider the case $\mathbf{R}^{\backslash[1]}_{i,j}>0$, since $c$ can be any negative number, the condition $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=u_{k}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only when $u_{k}=\infty$. For the fourth scenario, the situation is similar. $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=c^{2}u_{l}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only with $u_{l}=\infty$. We get the final result by combining all these scenarios. ∎
The exact OGM with feasible solution (10) and $\mathbf{u}=\infty$ provides the convergence rate of
$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{LR^{2}}{4A_{K}}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}=\frac{LR^{2}}{4A_{K}}.$
The unique feasible solution $\mathbf{u}=\infty$ has no effect on the final convergence rate since $\ b_{i}\equiv 0$. However, with the inexact gradient oracle, such a solution will ruin the convergence rate as the accumulated error approaches $\infty$. Hence, we need to change either the feasible solution given in (10) or the algorithm’s structure defined in (11). The details will be discussed in the next section.
3 Generalized OGM with inexact gradient oracle
The properties of OGM and its generalization have been thoroughly studied. However, the convergence analysis of OGM with inexact gradient oracle has not been performed yet. Furthermore, the existing analysis of FGM with inexact oracles with the absolute error assumption depends on the trajectory of sequences. This section aims at closing these gaps.
It is possible to optimize the algorithm with the same procedure as OGM, but the solution depends on iteration number $K$, initial condition $R$, and inexactness levels $\{b_{i}\}_{i=1}^{K}$. It is difficult to find a recursive structure of the algorithm that captures the numerical solution (see the experiment results in Appendix B). Hence, we focus on Generalized OGM (GOGM) with inexact oracle and try to find its convergence rate with $b_{i}>0$. Furthermore, the convergence bound of Generalized FGM (GFGM) is also analyzed in Section 3.3.
The GOGM is a generalization of OGM (Algorithm 4) as it only requires the stepsize to satisfy $\alpha_{k}^{2}\leq A_{k}$. The inexact GOGM (iGOGM) is provided in Algorithm 5, in which the gradients in steps 2 and 3 are replaced by their inexact alternatives.
Algorithm 5 Inexact Generalized Optimized Gradient Method (iGOGM)
0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for
The update rule in step 4 guarantees $\alpha_{k}^{2}\leq A_{k}$ when $\lambda_{k}\in[0,1]$ as
$\alpha_{k+1}^{2}=(A_{k+1}-A_{k})^{2}=\lambda_{k+1}A_{k+1}\leq A_{k+1}.$
Furthermore, when $\lambda=1$, the $\mathbf{x}_{k+1}$ update (step 5) reduce to
$\mathbf{x}_{k+1}=\left(1-\frac{1}{\alpha_{k+1}}\right)\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},$
the common update step in both FGM and OGM.
Remark 3.1.
The optimality measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ used in the analysis of exact GOGM is also a meaningful one for the inexact problem. For iGOGM, we have
$f(\mathbf{y}_{K+1})=$ $f\left(\mathbf{x}_{K}-\frac{1}{L}(\nabla f(\mathbf{x}_{k})+\mathbf{e}_{K})\right)$
$\leq$ $f(\mathbf{x}_{K})-\frac{1}{L}\left\langle\nabla f(\mathbf{x}_{K}),\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\rangle+\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\|^{2}$
$=$ $f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}+\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}.$
Comparing to the measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$, the extra term $\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}$ can be discarded since the constraint $\left\|\mathbf{e}_{K}\right\|^{2}\leq b_{K}^{2}$ is the only one that involves $\mathbf{e}_{K}$ and the PEP solution sets $\left\|\mathbf{e}_{K}\right\|^{2}=b_{K}^{2}$. Since the algorithm only generates $\{\mathbf{x}_{k}\}$ and $\{\mathbf{y}_{k}\}$ up to $k=K$, $\mathbf{y}_{K+1}$ is a dummy variable used for the theoretical guarantee, and setting $b_{K}=0$ does not affect the complexity of the algorithm.
3.1 PEP approach for the convergence analysis of iGOGM
The goal of this part is to find an analytical feasible solution of (D) with iGOGM stepsize. The feasible solution provides an upper bound of the worst-case convergence rate. To do so, first notice that the values of $\{v_{*,i}\}$ are uniquely determined by $\{v_{i,i+1}\}$ through the equality constraint in (D) and, hence, can be removed from the decision variables. Next, consider the decision variable $\tau$. Recall the objective of the problem (D), $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, contains two components: the first one is the convergence rate and the second one is the accumulated error. Given the convergence rate of the exact GOGM, i.e., $\frac{LR^{2}}{4A_{K}}$, we adjust the objective of (D) as
$\frac{L}{4A_{K}}R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2},$
which is the summation of the convergence rate of exact GOGM and the accumulated error $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$. With these replacements, the decision variables we need to consider are $\{v_{i,i+1}\}$ and $\{u_{i}\}$. To simplify the notation, in the following text, we will use $\mathbf{v}$ and $\mathbf{u}$ to represent the sequences $\{v_{i,i+1}\}$ and $\{u_{i}\}$, respectively.
Solving (D) with this new objective, we observe from the numerical solution that $\mathbf{v}$ is almost equal to its analytical counterpart for problems with exact oracles given in (10). In this experiment, we set $K=\{10,20,\cdots,100\}$, $R=\{1,10,1000\}$, and $L=\{1,10,1000\}$ which result in 90 different scenarios in total. For each scenario, we generate 10 instances with random stepsize parameter $\{\lambda_{k}\}$ and inexactness levels $\{b_{i}\}$. In Figure 1, three examples of the relative difference between the numerical solution of inexact GOGM $v^{num.}_{k}$ and the analytical solution of exact GOGM $v^{anly.}_{k}$ are illustrated. The solid line is an average of 90 instances. Furthermore, Figure 2 illustrates the histogram of relative error for all 900 instances which also supports this conjecture.
Figure 1: Difference between the numerical solution and the analytical solution (for the problem with exact oracle), i.e., (10), for $\mathbf{v}$. Solid lines are the average of 10 instances and the shaded areas represent the variation of individual instances.
Figure 2: Difference between numerical and analytical solution for $\mathbf{v}$
As we claimed in Lemma 2.4, OGM with (10) has a unique feasible solution with $\mathbf{u}=\infty$. To make $\mathbf{u}$ to be bounded, we have to guarantee that the stepsize condition holds with strict inequality, i.e. $A_{k}-\alpha_{k}^{2}>0$. Recall the matrix $\mathbf{R}^{\backslash[1]}$
$\mathbf{R}^{\backslash[1]}=\begin{pmatrix}[l][\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times K+1}\\ [\mathbf{B}^{\top}]_{K+1\times K}&[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]_{K+1\times K+1}\end{pmatrix},$
where the elements of the first column of $\mathbf{B}$ and $\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}$ are zero. So we can discard this zero row and column and write the $\mathbf{R}^{\backslash[1,K+2]}$ as
$\mathbf{R}^{\backslash[1,K+2]}=\begin{pmatrix}[l][\mathbf{U}]_{K\times K}&[\mathbf{B}^{\backslash[1]}]_{K\times K}\\ [\mathbf{B}^{\backslash[1]\top}]_{K\times K}&[\mathbf{D}]_{K\times K}\end{pmatrix},$
with $\mathbf{D}\triangleq[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]^{\backslash[1]}$.
Example of $\mathbf{R}^{\backslash[1,K+2]}$ with $K=3$ is
$\mathbf{R}^{\backslash[1,K+2]}=\left(\begin{array}[]{ccc|cccc}{u}_{0}&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&{u}_{1}&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{2}&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&\frac{A_{1}-\alpha_{1}^{2}}{LA_{3}}&0&0\\ \frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&0&\frac{A_{2}-\alpha_{2}^{2}}{LA_{3}}&0\\ \frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&0&0&\frac{A_{3}-\alpha_{3}^{2}}{LA_{3}}\end{array}\right).$
We need to find $\mathbf{u}$ such that $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$. Here we notice that $\mathbf{D}=\operatorname{diag}(\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})\succ 0$, hence, it is invertible with all elements strictly bigger than 0. We take the Shur complement of the block $\mathbf{D}$ and define the result matrix as $\mathbf{S}$ with
$\mathbf{S}_{i,j}=\begin{cases}u_{i-1}-\frac{(A_{i-1}+2\alpha_{i-1}\alpha_{i})^{2}}{4LA_{K}(A_{i}-\alpha_{i}^{2})}-\sum_{k=i}^{K-1}\frac{\alpha_{i-1}^{2}\alpha_{k+1}^{2}}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\quad i=j\\ -\frac{\alpha_{i}\alpha_{j-1}(A_{i-1}+2\alpha_{i-1}\alpha_{i})}{2LA_{K}(A_{i}-\alpha_{i}^{2})}-\sum_{k=i}^{K-1}\frac{\alpha_{i-1}\alpha_{j-1}\alpha_{k+1}^{2}}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\quad i>j\\ \mathbf{S}_{ji}\quad j>i\end{cases}.$
Example with $K=3$ is
$\mathbf{S}=\begin{pmatrix}{u}_{0}-\frac{\left(1+2\alpha_{1}\right)^{2}}{4LA_{3}\left({A}_{1}-{\alpha}_{1}^{2}\right)}-\frac{{\alpha}_{2}^{2}}{LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{1}-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right)^{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}^{2}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{2}-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right)^{2}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\end{pmatrix}.$
Then, $\mathbf{S}\succeq 0$ iff $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$. A nice property of $\mathbf{S}$ is that $\{u_{i}\}$ only shows in the diagonal elements and all the off-diagonal elements are negative constants. It is known that a symmetric diagonally dominated matrix with non-negative diagonal entries is PSD, with the definition of the diagonally dominated matrix as
$\left|\mathbf{S}_{i,i}\right|\geq\sum_{i\neq j}\left|\mathbf{S}_{i,j}\right|.$
Based on the special structure of $\mathbf{S}$, we can simplify the above inequality as
$\mathbf{S}_{i,i}\geq-\sum_{i\neq j}\mathbf{S}_{i,j}.$
Inserting the definition of $\mathbf{S}_{i,j}$, we can calculate the theoretical feasible solution for $\mathbf{u}$ as
[…]
The solution above also shows that when $A_{k}-\alpha_{k}^{2}=0$, $u=\infty$, which recovers the unique feasible solution we proved in Lemma 2.4.
Surprisingly, this solution matches the numerical solution when $b_{i}\equiv b$. In Figure 3, we follow the same settings as for Figure 2 with $\mathbf{v}$ equals to the values defined in Equation 10. The error between the optimal solution and the analytical $\hat{\mathbf{u}}$ given in (12) is under $10^{-3}$, validating our theoretical result.
Figure 3: Difference of the numerical and analytical solution $\hat{\mathbf{u}}$ given in (12) for different settings with fixed $b_{i}$ along iterations.
Finally, in Lemma 3.2, we provide an analytical solution for $\mathbf{u}$ that together with the analytical solutions for $\mathbf{v}$ and $\mathbf{v}_{*}$ (from the problem with exact oracle) provide a feasible solution to (D).
From the above discussion, we can summarize the results in the following lemma. The lemma provides a feasible solution to problem (D) with iGOGM algorithm.
Lemma 3.2 (Feasible solution to (D) with iGOGM).
For Algorithm 5, the following solution is a feasible solution to the problem (D). $\forall i=0,\cdots,K-1$
$\tau$ $=\frac{L}{4A_{K}},\quad v_{i,i+1}=\frac{A_{i}}{A_{K}},\quad v_{*,0}=v_{0,1},\quad v_{*,i}=v_{i,i+1}-v_{i-1,i},\quad v_{*,K}=1-v_{K-1,K},$
$u_{i}$ $=\frac{A_{i}(1+2\alpha_{i+1})(A_{i}+2\alpha_{i}\alpha_{i+1})}{4LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}+\sum_{k=i+1}^{K-1}\frac{A_{k}(1+2\alpha_{k+1})\alpha_{i}\alpha_{k+1}}{2LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}.$
3.2 PEP-inspired convergence analysis of iGOGM
As explained by [30], the feasible solution of PEP provides the proof of the convergence bound for the first-order algorithms. In the following theorem and the corresponding proof, we will show a PEP-inspired approach that can be verified without any knowledge of PEP or even SDP.
Theorem 3.3.
The sequence generated by Algorithm 5 satisfies
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
with $u_{k}=\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}$.
Proof.
Below are the inequalities used in PEP primal constraints and their corresponding dual variables:
[…]
Multiply inequalities with their corresponding dual variable and sum up, we have the inequality
$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mua\mkern-1.5mu)}}{{=}}$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mui\mkern-1.5mu)}}{{=}}$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mub\mkern-1.5mu)}}{{=}}$ $-\frac{4}{L}\left\|\frac{L}{4}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{4}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle,$
where (a) replaces $\theta_{k,i}$ with (11), and (b) uses $\tau=L/4$ and the fact that $A_{0}=\alpha_{0}=1$. If $\mathbf{e}_{i}=\mathbf{0}$, then the above inequality gives the convergence rate of iGOGM with exact oracle.
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|}{4A_{K}}$
Next, we will bound the error terms when $\mathbf{e}_{i}\neq\mathbf{0}$. For any $k=0,\cdots,K-1$,
$-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5muii\mkern-1.5mu)}}{{=}}$ $\frac{-1}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}$
$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}^{2}\alpha_{i}^{2}\left\|\mathbf{e}_{i}\right\|^{2}+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k}\sum_{j=0}^{i-1}2\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle$
$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}A_{k}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle+\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}.$
Summing over $k$, and changing the summation indices, we have
$\sum_{k=0}^{K-1}\left(-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle\right)$
$\leq$ $\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
For notation simplicity, we use $P_{k,i}$ to replace the coefficients of $\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$ as
$P_{k,k}\triangleq$ $\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})},$
$P_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}.$
Hence, the above equation can be rewritten as
$\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}P_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}(\left\|\mathbf{e}_{k}\right\|^{2}+\left\|\mathbf{e}_{i}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=k+1}^{K-1}\frac{1}{2}P_{i,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\sum_{k=0}^{K-1}\left(P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
With simple calculations, we can show
$P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}$
$=$ $\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2L(A_{i+1}-\alpha_{i+1}^{2})}.$
Inserting these results into the original inequality, we get the final bound. ∎
The key steps of the proof are the equalities $(i)$, $(ii)$, and inequality $(iii)$. The equality $(i)$ can be seen as the step of $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$ if one uses the Gram matrix $\mathbf{G}$ and standard basis vectors to rewrite this equality. The equality $(ii)$ can be seen as the reformulation of the Shur complement with the block $[\frac{A_{k+1}-\alpha_{k+1}^{2}}{LA_{K}}]$ of $\mathbf{R}^{\backslash[1,K+2]}$. The inequality (iii) is indeed the step of finding $\{u_{k}\}$ such that the matrix $\mathbf{S}$ is diagonally dominant. This observation shows that the steps we use to find and verify the feasible solution ($\mathbf{M}\succeq 0$) can be used to generate the theoretical proof.
When $\left\|\mathbf{e}_{k}\right\|\equiv 0$, the error term will be zero and the upper bound recovers the result of the exact GOGM. When $\left\|\mathbf{e}_{k}\right\|\neq 0$, this term is only determined by the total iteration number $K$ and the stepsize $\{\alpha_{k}\}$. It is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ and can be calculated when $\{\alpha_{k}\}$ is fixed.
Remark 3.4.
The inexactness assumption and corresponding analysis in [49] can also handle unbounded feasible regions. However, their inexactness assumption for convex problems requires the gradient estimate to be a subgradient, which is not a requirement in our work. Furthermore, their inexactness level along iterations has to be the same. Finally, in their FGM analysis, the accumulated error depends on the initial condition $R$. In contrast, our approach provides a pure separation of the convergence rate with the exact gradient oracle and the accumulated error. Our error term is independent of $R$ and holds for any arbitrary error along iterations.
3.3 PEP-inspired convergence analysis of iGFGM
Inspired by the approach for the analysis of iGOGM, we also find a convergence rate of inexact Generalized FGM (iGFGM) with the algorithm shown in Algorithm 6 in Section A.2. Similar to the result of iGOGM, the bound of iGFGM is a summation of the exact convergence rate and accumulated error.
Theorem 3.5.
The sequence generated by Algorithm 6 satisfies
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{2A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
with $u_{k}=\frac{A_{k}^{2}(1+\alpha_{k+1})}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2LA_{K}(2A_{i}-\alpha_{i}^{2})}$.
4 Exploiting the rate-error tradeoff and finding the optimal inexactness schedule
In Theorem 3.3, we provided the convergence bound of iGOGM with the explicit expression for the accumulated error. This section first looks into the tradeoff between the convergence rate and accumulated error through the stepsize with a fixed gradient inexactness level along iterations. Second, we aim to optimize the inexactness levels across iterations, for a given stepsize, so that the total effort to obtain the gradient oracles is minimized.
4.1 Convergence rate and the accumulated error tradeoff
Considering the convergence bound $\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$, with $\{A_{k}\},\{u_{k}\}$ being functions of $\{\alpha_{k}\}$, one can propose to minimize the bound with respect to $\{\alpha_{k}\}$. However, we note that such a problem is a high-order polynomial optimization and its solution is intractable in general.
Even if one is satisfied with the numerical solution to find the optimized algorithm, then it is preferable to solve the minimax problem (9) following the relaxation approach in [22] which we also used in Appendix B. This procedure generates the optimized stepsizes and the resulting convergence bound is better than the solution of the polynomial optimization problem as the derived theoretical $\{u_{i}\}$ in Lemma 3.2 is not necessarily the optimal solution.
Here we narrow down the iGOGM to iOGM-a, with the stepsize $\alpha_{k}=\frac{i+a}{a}$ and $A_{k}=\sum_{i=0}^{k}\frac{i+a}{a}=\frac{(k+2a)(k+1)}{2a}$. The condition $A_{k}>\alpha_{k}^{2}$ holds for any $k$ when $a>2$. Note that when $K\gg a$, the convergence rate is still $\mathcal{O}(K^{-2})$ in the exact case. Furthermore, [42] shows that the exact algorithm OGM-a has the asymptotic worst-case bound for function value as $\frac{a}{2}K^{-2}$ and the smallest gradient norm as $\frac{a\sqrt{6}}{2\sqrt{a-2}}K^{-1.5}$; it achieves the best performance of these two measures simultaneously when $a=4$. Both the simplicity of the stepsize and the good performances over function value and gradient norm indicate that OGM-a is a good selection under exact oracles and is worth further study for the inexact oracles too.
Replacing the stepsizes of OGM-$a$ in Theorem 3.3, we get the coefficients of the error as
$u_{k}=\frac{1}{4La(K+1)(K+2a)((a-2)k+2a^{2}-2a-2)}\left((2a+8)k^{4}+(11a^{2}+48a+16)k^{3}\right.$
$\left.+(20a^{3}+107a^{2}+70a+8)k^{2}+(12a^{4}+104a^{3}+100a^{2}+24a)k+36a^{4}+48a^{3}+16a^{2})\right)$
$+\sum_{i=k+1}^{K-1}\frac{(k+a)(2i^{3}+(9a+4)i^{2}+(13a^{2}+13a+2)i+6a^{3}+10a^{2}+4a}{La(K+1)(K+2a)((a-2)i+2a^{2}-2a-2)},$
which can be further simplified when $a$ is assigned with a value.
Take $a=4$ as an example, which has both best convergence rates for function value and the gradient norm under exact oracle, the $u_{k}$ is
$u_{k}=$ $\frac{16k^{4}+384k^{3}+3280k^{2}+11424k+12544}{16L(K+1)(K+8)(2k+22)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(2i^{3}+40i^{2}+262i+560)}{4L(K+8)(K+1)(2i+22)}$
$=$ $\frac{(k+7)^{2}(k+8)(k+2)}{2L(K+8)(K+1)(k+11)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(i+8)(i+7)(i+5)}{4L(K+1)(K+8)(i+11)}.$
Replacing $k+11$ with $k+8$ and $i+11$ with $i+8$ in the two denominators, we have a simpler expression with slightly bigger $u_{k}$ as
$u_{k}=\frac{(k+7)^{2}(k+2)}{2L(K+8)(K+1)}+\frac{(k+4)(K-k-1)(2K^{2}+2Kk+35K+2k^{2}+37k+210)}{24L(K+8)(K+1)}.$
Assume $b_{k}\equiv b$, then the accumulated error is calculated as
$b^{2}\sum_{k=0}^{K-1}u_{k}=\frac{b^{2}K(12K^{3}+303K^{2}+2687K+8758)}{480L(K+8)}.$
The above results indicate that the accumulated error is in $\mathcal{O}(\frac{b^{2}K^{3}}{L})$.
For any arbitrary $a$ values, it is difficult to derive a closed form of $u_{k}$ since the denominator can only be simplified as $k+c$ structure when $a$ is given as a specific number. Hence, we provide a numerical result of $u_{k}$ for given $K$ with different $a$ in Figure 4. For the same $a$ and $k$, bigger $K$ results in bigger $u_{k}$. For all $a$ values, $\{u_{k}\}$ first increases and then decreases. For a fixed $k$ and $K$, bigger $a$ results in smaller $u_{k}$ values.
Figure 4: Values of $u_{k}$ for OGM-$a$ with different values of $a$ and $K$.
Figure 5: Convergence rate and accumulated error for OGM-$a$ for different $a$ and $K$ values.
Based on this result, to reduce the accumulated error, one can consider increasing $a$. However, it will adversely slow down the convergence rate as $A_{K}=\frac{(K+2a)(K+1)}{2a}$ and decreases when $a$ is increased. Under this circumstance, we provide the comparison of convergence rate and accumulated error with different iteration numbers $K$, and $a$ in Figure 5, where $\tau=\frac{L}{4A_{K}}$. The convergence rate $\tau$ is smaller when $a$ is smaller while the accumulated error is bigger. When $a=4$, $\sum_{k=0}^{K-1}u_{k}$ is between $K^{2}$ and $K^{3}$, better than our relaxed analysis result. Another interesting fact is that when $K\ll a$ ($a=10^{5}$ and $a=10^{6}$), the convergence rate $\tau=\mathcal{O}(K^{-1})$ and the accumulate error is also $\mathcal{O}(K^{1})$. If $K\ll a$, we have $\alpha_{k}=1$, and the proposition below provides the convergence bound of this extreme scenario.
Proposition 4.1.
With the stepsize defined as $\alpha_{k}\equiv 1$, the sequence generated by iGOGM (Algorithm 5) satisfies
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\sum_{k=0}^{K-1}\frac{3(2K-k+1)}{4L(K+1)}\left\|\mathbf{e}_{k}\right\|^{2}.$
Furthermore, when $\left\|\mathbf{e}_{k}\right\|\leq\bar{b},\forall k$, the above inequality becomes
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\frac{9K}{8L}\bar{b}^{2}.$
4.2 Optimal inexactness schedule
Since the derived $u_{k}$ is not constant, it will be beneficial to set $b_{k}$ based on the value of $u_{k}$. Recall the definition of $\mathbb{\eta}$ from (2). When $r=1$, we can aim to optimize over the total $\eta$-complexity, i.e., $\sum_{k=0}^{K-1}\eta_{k}$, so that the accumulated error will not deteriorate the convergence rate. The problem is given as
[…]
$\operatorname{s.t.}\quad$ $\sum_{k=0}^{K-1}u_{k}b_{k}^{2}\leq\frac{LR^{2}}{4A_{K}}.$
This problem is a convex optimization when $h$ is convex and decreasing over linear constraint and at its optimal solution, the constraint is binding. Hence, the inequality constraint can be replaced with the equality one. Let $\lambda$ be the dual multiplier of the constraint, the Lagrangian function is
$\mathcal{L}(\mathbf{b},\lambda)=\sum_{k=0}^{K-1}h^{-1}(b_{k})+\lambda\left(\sum_{k=0}^{K-1}u_{k}b_{k}^{2}-\frac{LR^{2}}{4A_{k}}\right).$
At the optimal solution $\mathbf{b}^{*}$ and $\lambda^{*}$, from the optimality conditions, we have
[…]
Solving the above equations, we have a closed-form solution of $\mathbf{b}^{*}$.
In the following, we give the closed-form of $\mathbf{b}^{*}$ for two examples of $h(\eta)$ [47], which are two common structures of the inexact gradient oracles.
Lemma 4.2 ($h(\eta)$ with power law decay).
Let $h(\eta)=c_{1}\eta^{-c_{2}}$ with $c_{1},c_{2}>0$ to be two constants. The inverse function and its gradient are
$h^{-1}(b)=\left(\frac{c_{1}}{b}\right)^{1/c_{2}}\quad\text{and}\quad\nabla_{b}h^{-1}(b)=-\frac{c_{1}^{1/c_{2}}b^{-(1+c_{2})/c_{2}}}{c_{2}}.$
Furthermore, the optimal primal and dual solutions of (OPT-b) are
$b_{k}^{*}=\frac{\sqrt{L}R}{2\sqrt{A_{K}\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}}u_{k}^{d/(1+2c_{2})}},\quad\lambda^{*}=\frac{c^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
Proof.
With power law decay, (13) is written as
$-\frac{c^{1/c_{2}}(b_{k}^{*})^{-(1+c_{2})/c_{2}}}{c_{2}}+2\lambda^{*}b_{k}^{*}u_{k}=0,\quad k=0,1,\cdots,K-1.$
From this equation, $b_{k}^{*}$ can be calculated as
[…]
Inserting this equation into (13), and solve for $\lambda^{*}$ we get
$\lambda^{*}=\frac{c_{1}^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
Replacing this solution into (15), the final result can be derived. ∎
Lemma 4.3 ($h(\eta)$ with exponential decay).
Let $h(\eta)=q_{1}q_{2}^{-\eta}$ with $q_{1}>0$ and $q_{2}>1$ be two constants. Then
$h^{-1}(b)=\frac{\log q_{1}-\log b}{\log q_{2}},\quad\nabla_{b}h^{-1}(b)=-\frac{1}{b\log q_{2}},$
and, the optimal primal and dual solutions are
$b_{k}^{*}=\sqrt{\frac{LR^{2}}{4KA_{K}u_{k}}},\quad\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
Proof.
Similar to the proof of Lemma 4.2, from (13) we have
$u_{k}(b_{k}^{*})^{2}=\frac{1}{2\lambda^{*}\log q_{2}}.$ (16)
Inserting this into (14),
$\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
By the previous two equations, we can derive the final solution. ∎
To quantify the improvement after the optimized inexactness level, we calculate the total $\eta$-complexity for iOGM-4 algorithm for the two structures of $h(\eta)$ functions. We fix $L=1$ and $R=1$, then solve the constant $b$ through
$\bar{b}^{2}\sum_{k=0}^{K-1}\leq\frac{LR^{2}}{4A_{K}},$
and calculate the optimized $b$ by the solutions we derived in Lemmas 4.2 and 4.3. In Figure 6, we report the result for power law decay with four different values of $c_{2}$. Note that the $\eta$-complexity decreases with increased $c_{2}$. If we focus on the improvement of $\eta$-complexity for the optimized $\{b_{k}\}$, the improvement is in the same order as the complexity of constant $b$, which shows optimizing over $b$ can decrease the $\eta$-complexity significantly. As for the exponential decay in Figure 7, the improvement is not as large as that of the power law decay. Although the improvement is not as large as that of the power law decay, considering its absolute value, it still represents a considerable improvement.
Figure 6: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with power law decay with $c_{1}=1$ and $L=1,R=1$.
Figure 7: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with exponential decay with $q_{1}=1,q_{2}=e$ and $L=1,R=1$.
5 Conclusions
In this paper, we analyze the Generalized Optimal Gradient Method (GOGM) with the inexact gradient oracle under the absolute error assumption. By utilizing the Performance Estimation Problem (PEP) analysis tool, a new upper bound for the convergence rate of GOGM is derived. This bound is composed of two independent components, the convergence rate obtained from the exact gradient oracle and the accumulated error resulting from the gradient inexactness. Such a bound demonstrates the effect of the inexactness and does not require boundedness of the feasible region. Furthermore, this bound is independent of the initial condition.
From this convergence bound, the optimal schedule to set oracle inexactness along iterations is proposed. Such a study aims to minimize the sampling or computational effort for gradient estimation while maintaining the convergence rate.
Appendix A Supplementary lemmas and proofs
A.1 Proof of composition optimization
Lemma A.1.
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla f(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Also, $f(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$ and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Assume $\left|\mathcal{S}_{\nabla}\right|$ and $\left|\mathcal{S}_{\mathbf{g}}\right|$ are properly chosen so $\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\leq 1$, then the error of the estimated gradient is bounded as
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq b^{2},$
with probability at least $1-\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}$.
Proof.
By Young’s inequality,
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$=$ ~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))+\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\right\|^{2}\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
$\leq$ $2C_{f}^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2C_{g}^{2}L_{f}^{2}\left\|\tilde{\mathbf{g}}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}.$
Taking the expectation on both sides and with the bounded variance assumptions, we have
$\mathbb{E}[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}]$
$\leq$ $\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}.$
By Markov’s inequality, we have
$\Pr\left(\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\geq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)\right)$
$\leq$ $\frac{1}{\epsilon}\mathbb{E}\left[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\right]/\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
$\leq$ $\frac{1}{\epsilon}.$
Assuming $\epsilon>1$, i.e., the bound holds non-triviality, we can conclude that
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
holds with probability at least $1-\frac{1}{\epsilon}$.
A.2 Convergence rate for inexact generalized FGM
Algorithm 6 Inexact Generalized Fast Gradient Method (iGFGM)
0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{1}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for
Lemma A.2.
The sequence $\{\mathbf{x}_{k}\}$ generated by Algorithm 6 is the same sequence as the one generated by iGFO ($\hat{\mathbf{x}}_{k}=\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k-1}\theta_{k,i}\tilde{\nabla}f(\hat{\mathbf{x}}_{i})$) with
$\theta_{k,i}=\frac{\alpha_{i}(A_{k}-A_{i})+A_{i}}{A_{k}}.$
Proof.
The equivalence of these two sequences is shown through induction. For $k=0$ and $k=1$, the equivalence holds trivially. Assume $\mathbf{x}_{i}=\hat{\mathbf{x}}_{i}$ for all $i\leq k$ with $k>1$, then
$\mathbf{x}_{k+1}=$ $\left(1-\frac{\alpha_{k+1}}{A_{k+1}}\right)\left(\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})\right)+\frac{\alpha_{k+1}}{A_{k+1}}\left(\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k}\alpha_{i}\tilde{\nabla}f(\mathbf{x}_{i})\right)$
$=$ $\left(1-\frac{\alpha_{k+1}}{A_{k+1}}\right)\left(\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k-1}\theta_{k,i}\tilde{\nabla}f(\mathbf{x}_{i})-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})\right)+\frac{\alpha_{k+1}}{A_{k+1}}\left(\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k}\alpha_{i}\tilde{\nabla}f(\mathbf{x}_{i})\right)$
$=$ $\mathbf{x}_{0}-\frac{1}{L}\frac{A_{k+1}-\alpha_{k+1}}{A_{k+1}}\sum_{i=0}^{k-1}\frac{\alpha_{i}(A_{k}-A_{i})+A_{i}}{A_{k}}\tilde{\nabla}f(\mathbf{x}_{i})-\frac{1}{L}\frac{\alpha_{k+1}}{A_{k+1}}\sum_{i=0}^{k-1}\alpha_{i}\tilde{\nabla}f(\mathbf{x}_{i})$
$-\frac{1}{L}\left(1-\frac{\alpha_{k+1}}{A_{k+1}}+\frac{\alpha_{k+1}\alpha_{k}}{A_{k+1}}\right)\tilde{\nabla}f(\mathbf{x}_{k})$
$=$ $\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k-1}\frac{\alpha_{i}(A_{k}-A_{i})+A_{i}+(A_{k+1}-A_{k})\alpha_{i}}{A_{k+1}}\tilde{\nabla}f(\mathbf{x}_{i})-\frac{1}{L}\frac{\alpha_{k}\alpha_{k+1}+A_{k}}{A_{k+1}}\tilde{\nabla}f(\mathbf{x}_{k})$
$=$ $\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k-1}\frac{\alpha_{i}(A_{k+1}-A_{i})+A_{i}}{A_{k+1}}-\frac{1}{L}\frac{\alpha_{k}(A_{k+1}-A_{k})+A_{k}}{A_{k+1}}\tilde{\nabla}f(\mathbf{x}_{k})$
$=$ $\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k}\frac{\alpha_{i}(A_{k+1}-A_{i})+A_{i}}{A_{k+1}}\tilde{\nabla}f(\mathbf{x}_{i})$
$=$ $\mathbf{x}_{0}-\frac{1}{L}\sum_{i=0}^{k}\theta_{k+1,i}\tilde{\nabla}f(\mathbf{x}_{i})$
$=$ $\hat{\mathbf{x}}_{k+1}.$
Hence the sequences $\{\mathbf{x}_{k}\}$ and $\{\hat{\mathbf{x}}_{k}\}$ are equivalent to each other. ∎
Proof of Theorem 3.5
Proof.
Similar to the first step of the proof of Theorem 3.3,
$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{L}\sum_{k=1}^{K}\alpha_{k}\alpha_{k-1}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k-1})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\frac{2A_{0}-\alpha_{0}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{0})\right\|^{2}$
$-\sum_{k=0}^{K-1}\frac{1}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|(2A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+(\alpha_{k}\alpha_{k+1}+A_{k})\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}^{2}\alpha_{i}^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=0}^{i-1}\frac{\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle$
$\leq$ $\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
Following the same steps in the proof of iGOGM, define
$\hat{P}_{k,k}\triangleq$ $\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}$
$\hat{P}_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k+1}\alpha_{k}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})},$
we can bound the error terms as
$\sum_{k=0}^{K-1}\hat{P}_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\hat{P}_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
$\leq$ $\sum_{k=0}^{K-1}\left(\hat{P}_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}\hat{P}_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}\hat{P}_{i,k}\right)\left\|\mathbf{e}_{k}\right\|^{2}$
$=$ $\sum_{k=0}^{K-1}\left(\frac{A_{k}^{2}(1+\alpha_{k+1})}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2L(2A_{i}-\alpha_{i}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
Dividing both sides of the inequality by $A_{K}$, the final result is obtained. ∎
Appendix B Numerical solutions for optimized algorithm with inexact oracle
We provide the optimized stepsize for different $K$ with $L=1$ and $R=1$. We report three results, the first one is for $\mathbf{b}=\mathbf{0}$, the second one is for $b_{k}^{2}\equiv\bar{b}^{2}=0.01$ and the third one takes $\{b_{k}\}$ to be a decreasing sequence. Note that $\bar{b}^{2}$ in the second scenario is set equal to the average value of $b_{k}^{2}$ in the third scenario.
$K$ $\theta^{b=0}$
1 $\begin{pmatrix}1.6180\end{pmatrix}$
2 $\begin{pmatrix}1.6180&0\\ 1.7921&2.0193\end{pmatrix}$
3 $\begin{pmatrix}1.6180&0&0\\ 1.7921&2.0193&0\\ 1.8677&2.4618&2.2316\end{pmatrix}$
4 $\begin{pmatrix}1.6180&0&0&0\\ 1.7921&2.0193&0&0\\ 1.8677&2.4617&2.2316&0\\ 1.9078&2.6966&2.8856&2.3654\end{pmatrix}$
5 $\begin{pmatrix}1.6180&0&0&0&0\\ 1.7921&2.0193&0&0&0\\ 1.8676&2.4617&2.2315&0&0\\ 1.9078&2.6966&2.8855&2.3653&0\\ 1.9318&2.8373&3.2771&3.1828&2.4580\end{pmatrix}$
Table 1: Numerical solution of optimized first-order algorithm with $b_{i}\equiv 0$
$K$ $\theta^{b=\bar{b}}$
1 $\begin{pmatrix}1.5509\end{pmatrix}$
2 $\begin{pmatrix}1.5537&0\\ 1.7028&1.7642\end{pmatrix}$
3 $\begin{pmatrix}1.5552&0&0\\ 1.7075&1.7828&0\\ 1.7668&2.0875&1.7590\end{pmatrix}$
4 $\begin{pmatrix}1.5561&0&0&0\\ 1.7101&1.7926&0&0\\ 1.7719&2.1112&1.7962&0\\ 1.7984&2.2476&2.1371&1.6578\end{pmatrix}$
5 $\begin{pmatrix}1.3714&0&0&0&0\\ 1.4958&1.4457&0&0&0\\ 1.5444&1.6859&1.3960&0&0\\ 1.5675&1.8115&1.6837&1.3303&0\\ 1.5644&1.8444&1.7931&1.5876&1.2111\end{pmatrix}$
Table 2: Numerical solution of optimized first-order algorithm with $b_{i}\equiv\bar{b}$
$K$ $\theta^{b_{k}\downarrow}$
1 $\begin{pmatrix}1.5509\end{pmatrix}$
2 $\begin{pmatrix}1.5478&0\\ 1.7047&1.7954\end{pmatrix}$
3 $\begin{pmatrix}1.5495&0&0\\ 1.7043&1.7875&0\\ 1.7690&2.1161&1.8192\end{pmatrix}$
4 $\begin{pmatrix}0.0335&0&0&0\\ 0.0276&1.5575&0&0\\ 0.0259&1.7218&1.8158&0\\ 0.0252&1.7925&2.1663&1.8749\end{pmatrix}$
5 $\begin{pmatrix}0.0111&0&0&0&0\\ 0.0014&0.5086&0&0&0\\ 0.0005&0.4713&1.6643&0&0\\ 0.0002&0.4588&1.8883&1.8380&0\\ 0.0000&0.4531&1.9905&2.2202&1.8669\end{pmatrix}$
Table 3: Numerical solution of optimized first-order algorithm with $b_{i}$ decreasing
We can observe that the stepsize at each iteration with the exact gradient oracle remains the same while this is not the case with the inexact oracle. Even for the fixed inexactness level (second scenario), the optimized algorithm changes with the total iteration number. Another fact is that the optimized algorithm can identify a bad gradient estimate (see the first iteration of $\theta^{b_{k}\downarrow}$). It almost discards the first iteration since its inexactness is beyond some unknown threshold. Apart from the stepsize result, the optimized objective function value indicates that the initial condition $R$ plays a role since $\tau$ for the three scenarios has different values. This is different from the exact case [22, 39] as optimized stepsize is independent of $R$ and $L$.
Table 4: Objective values for the optimized algorithm
$K$ $\tau^{b=0}$ $\tau^{b=\bar{b}}$ $\tau^{b\downarrow}$ $\bar{b}^{2}\sum u^{b=\bar{b}}_{k}$ $\sum u^{b_{k}\downarrow}_{k}b^{2}_{k}$
1 0.0955 0.1123 0.0185 0.0185 0.0185
2 0.0520 0.0743 0.0271 0.0273 0.0271
3 0.0331 0.0583 0.0340 0.0339 0.0340
4 0.0230 0.0504 0.0310 0.0396 0.0310
5 0.0170 0.0500 0.0320 0.0400 0.0320

### v2

@@
 Abstract
-Performance analysis of first-order algorithms with inexact oracle has gained recent attention due to various emerging applications in which obtaining exact gradients is impossible or computationally expensive.
+Performance analysis of first-order algorithms with inexact oracles has gained recent attention due to various emerging applications in which obtaining exact gradients is impossible or computationally expensive.
 Previous research has demonstrated that the performance of accelerated first-order methods is more sensitive to gradient errors compared with non-accelerated ones.
@@
 Performance Estimation Problem (PEP) is used as the primary tool to analyze the convergence bounds of the underlying algorithms.
-By finding an analytical solution to PEP, we derive novel convergence bounds for Inexact Optimized Gradient Method (OGM) [39] and Inexact Fast Gradient Method (FGM) [16] with variable inexactness along iterations.
-Under the absolute error assumption, we derive bounds in which the accumulated errors are independent of the initial conditions and the trajectory of the sequences generated by the algorithms.
-Furthermore, we analyze the tradeoff between the convergence rate and accumulated error that guides finding the optimal stepsize.
-Finally, we determine the optimal strategy to set the gradient inexactness along iterations (if possible), ensuring that the accumulated error remains subordinate to the convergence rate.
+By finding an analytical solution to PEP, we derive novel convergence bounds of Generalized Optimized Gradient Method (GOGM) and Generalized Fast Gradient Method (GFGM) with inexact gradient oracles following the absolute error bound.
+The derived bounds allow varying oracle inexactness along the iterations; furthermore, their accumulated error terms are independent of the initial condition and any unknown parameters.
+Furthermore, we analyze the tradeoff between the vanishing term and the accumulated error in the convergence bound that guides finding the optimal stepsize.
+Finally, we determine the optimal strategy to set the gradient inexactness along iterations (if possible in a given application), ensuring that the accumulated error remains subordinate to the vanishing term.
 1 Introduction
@@
 This paper focuses on the class of merely convex functions with Lipschitz continuous gradient $f\in\mathcal{F}_{0,L}$.
-For this class of functions, the gradient descent (GD) method has the convergence rate $f(\mathbf{x}_{k})-f^{*}=\mathcal{O}(k^{-1})$.
-This rate can be improved to $\mathcal{O}(k^{-2})$ with Nesterovs’s fast gradient method (FGM) [51, 52, 50, 53].
-Recently, using computer-aided analysis, the Optimized Gradient Method (OGM) has been proposed [22, 39] that improves the convergence rate of FGM by $\sqrt{2}$ constant.
-The above rates are based on having access to the exact gradient at any point, which is not the case in many applications.
-The convergence rate deteriorates when there is an error in the gradient estimate.
-This paper considers gradient error in the absolute form as
+For this class of functions, the gradient descent (GD) method has the iteration complexity $f(\mathbf{x}_{k})-f^{*}=\mathcal{O}(k^{-1})$.
+This result can be improved to $\mathcal{O}(k^{-2})$ with Nesterov’s fast gradient method (FGM) [43, 44, 42, 45].
+Recently, using computer-aided analysis, the Optimized Gradient Method (OGM) has been proposed [18, 32] that improves the complexity of FGM by $\sqrt{2}$ constant.
+The above results are based on having access to the exact gradient at any point, which is not the case in many applications.
+The performance of the algorithms deteriorates when there is an error in the gradient estimate.
+We consider the gradient error as
 […]
 where the error $\mathbf{e}$ can be controlled by a parameter $\boldsymbol{\eta}\in\mathbb{R}^{r}$ through a positive function $h(\cdot)$ and the bound holds either deterministically, which is the focus of this work, or with high probability in stochastic settings.
-The parameter $\boldsymbol{\eta}$ represents, e.g., sampling size, computational effort, or accuracy of an approximation to achieve a desired gradient inexactness.
-Below, the gradient oracle inexactness is illustrated through a few examples.
-1.1 Examples of inexact gradient oracles
-Gradient inexactness is generally the result of the oracle not being evaluated at the desired point.
-For instance, if some or all coordinates of the desired point are the solution of an optimization problem that is only solved to a suboptimal solution, the resulting gradient will be inexact, e.g., bilevel optimization.
-In some other scenarios, the desired point results from an expectation that can only be estimated resulting in gradient inexactness see, e.g., nested composition optimization.
-Example 1.1 (Bilevel optimization [29]).
+The parameter $\boldsymbol{\eta}$ represents, e.g., computational effort, the accuracy of an approximation, or sampling size to achieve a desired gradient accuracy.
+1.1 Applications with inexact gradient oracles.
+In some applications, gradient inexactness is the result of the oracle not being evaluated at the desired point.
+For instance, in bilevel optimization, one block of coordinates of the upper-level problem is the solution of the lower-level optimization problem.
+However, in practice, the lower-level problem can be solved to a suboptimal solution, which results in upper-level gradient inexactness.
+Similarly, in composition optimization, the desired point results from an expectation that in many scenarios can only be estimated, resulting in gradient inexactness.
+These two applications are discussed below.
+Bilevel optimization.
 Consider the bilevel optimization problem
@@
 $\operatorname{s.t.}$ $\quad\mathbf{y}^{*}(\mathbf{x})=\operatorname*{argmin}_{\mathbf{y}}g(\mathbf{x},\mathbf{y}),$
-with $f$ being continuously differentiable and $g$ being two times continuously differentiable and strongly convex functions.
-To solve the problem with a gradient-based method, we first obtain the gradient with respect to $\mathbf{x}$ as
-$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$ $=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
-$=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla\mathbf{y}^{*}(\mathbf{x})^{\top}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x})).$
-This requires solving the lower-level problem to optimality, which is unattainable in many scenarios.
-When the lower-level problem is solved to a suboptimal point $\tilde{\mathbf{y}}(\mathbf{x})$ and the gradient is evaluated at this point instead of $\mathbf{y}^{*}(\mathbf{x})$, the resulting gradient is inexact:
-$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\neq\tilde{\nabla}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))\triangleq\nabla_{\mathbf{x}}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x})).$
-Under some continuity assumptions, the gradient error is shown to be bounded as
-~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left\|\tilde{\mathbf{y}}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|,$
-where $C\triangleq L_{f_{\mathbf{x}}}+\frac{L_{f_{\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}}+C_{f_{\mathbf{y}}}\left(\frac{L_{g_{\mathbf{x}}\mathbf{y}}}{\mu_{g}}+\frac{L_{g_{\mathbf{y}\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}^{2}}\right)$, with $L_{f_{\mathbf{x}}},L_{f_{\mathbf{y}}}$ are the Lipschitz continuity constants of $\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y})$ and $\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})$, respectively; $L_{g_{\mathbf{x}\mathbf{y}}},L_{g_{\mathbf{y}\mathbf{y}}}$ are the Lipschitz constants of $\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}),\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y})$ with respect to (w.r.t.) $\mathbf{y}$, respectively; $C_{f_{\mathbf{y}}}$, $C_{g_{\mathbf{x}\mathbf{y}}}$ are the upper bounds on $\left\|\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})\right\|$ and $\left\|\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y})\right\|$, respectively; and $\mu_{g}$ is the strong convexity constant of $g(\mathbf{x},\mathbf{y})$ w.r.t.
-$\mathbf{y}$– see [29] for details.
-Now let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations.
-From the convergence rate of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
-~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
-where $Q_{g}\triangleq L_{g}/\mu_{g}$. The above bound follows an exponential decay and represents the error bound in (2) with $\eta$ being the iteration number $k$.
-Example 1.2 (Composition optimization).
-Consider the nested composition optimization in the form of
-$\min_{\mathbf{x}}\ f(\mathbf{x})\triangleq h(\mathbf{g}(\mathbf{x}))\quad\text{where}\quad\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{N}\mathbf{g}_{i}(\mathbf{x}),$
+with $f$ being continuously differentiable and $g$ being two times continuously differentiable and strongly convex.
+To solve the problem with a gradient-based method, the gradient of the upper-level problem with respect to $\mathbf{x}$ is
+$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x})),$
+which requires solving the lower-level problem to optimality.
+When the lower-level problem is solved to a suboptimal point $\tilde{\mathbf{y}}(\mathbf{x})$ and the gradient is evaluated at this point instead of $\mathbf{y}^{*}(\mathbf{x})$, the resulting gradient is inexact.
+Let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations.
+From the iteration complexity of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
+$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
+where $Q_{g}\triangleq L_{g}/\mu_{g}$.
+The above bound follows an exponential decay and represents the error bound in (2) with $\boldsymbol{\eta}$ being the iteration number $k$ – For more details see A.1.
+Composition optimization.
+Consider the nested composition optimization problem
+$\min_{\mathbf{x}}\ f(\mathbf{x})\triangleq h(\mathbf{g}(\mathbf{x}))\quad\text{with}\quad\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{N}\mathbf{g}_{i}(\mathbf{x}),$
 where $h:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\mathbf{g}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ are continuously differentiable, and $N$ is a large positive integer.
 From the chain rule, the exact gradient is $\nabla f(\mathbf{x})=\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))$.
-However, assume one can only approximate $\mathbf{g}$ and $\nabla\mathbf{g}$ through sampling as
-$\tilde{\mathbf{g}}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\mathbf{g}}\right|}\mathbf{g}_{\mathcal{S}_{\mathbf{g}}[i]}(\mathbf{x})\quad\tilde{\nabla}\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\nabla}\right|}\nabla\mathbf{g}_{\mathcal{S}_{\nabla}[i]}(\mathbf{x}),$
-where $\mathcal{S}_{\mathbf{g}}$, $\mathcal{S}_{\nabla}$ are two subsets of $\{1,2,\cdots,N\}$.
-Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$.
-Furthermore, assume $h(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$, and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$.
-Then with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
-~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right).$
-The proof of the above bound is presented in Lemma A.1.
-This bound represents the error bound in (2) with $\boldsymbol{\eta}=(\left|\mathcal{S}_{\nabla}\right|,\left|\mathcal{S}_{\mathbf{g}}\right|)$.
+However, assume one can only approximate $\mathbf{g}$ and $\nabla\mathbf{g}$ through their minibach samples $\mathcal{S}_{\mathbf{g}}\subseteq[N]$ and $\mathcal{S}_{\nabla}\subseteq[N]$, respectively, where $N]\triangleq\{1,2,\cdots,N\}$.
+Then, under some Lipschitz continuity and bounded variance assumptions, as shown in Lemma A.1, with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
+$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{C_{1}}{\left|\mathcal{S}_{\nabla}\right|}+\frac{C_{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\right).$
+This bound represents the error bound in (2) with $\boldsymbol{\eta}=\left(\left|\mathcal{S}_{\nabla}\right|,\left|\mathcal{S}_{\mathbf{g}}\right|\right)$.
 Furthermore, if $\mathcal{S}_{\nabla}=\mathcal{S}_{\mathbf{g}}$, for any fixed $\epsilon$, the bound follows a power law decay.
-Another common scenario is when $\nabla f$ is unattainable, but, they can be approximated by (inexact) function values, i.e., zero-order information, denoted by $\tilde{f}$.
-In many such scenarios, it can be assumed that $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq b_{f}$.
-When $f(\mathbf{x})$ has Lipschitz continuous gradient, the zero-order methods generate inexact gradient estimates with quantifiable error bounds (see [6] and references therein).
-Example 1.3 (Gradient estimation via forward finite difference).
-Let $\vec{\mathbf{v}}_{i}\in\mathbb{R}^{d}$ denote the unit vector with the i-th element equal to 1, and $l>0$ be the finite difference interval. Define the i-th element of the gradient estimator as
-[~$[\tilde{\nabla}f(\mathbf{x})]_{i}=\frac{\tilde{f}(\mathbf{x}+l\vec{\mathbf{v}}_{i})-\tilde{f}(\mathbf{x})}{l}.$
-The error of the gradient estimate is shown by [6] to be bounded as
-~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\frac{\sqrt{d}Ll}{2}+\frac{2\sqrt{d}b_{f}}{l},$
+Another scenario is when $\nabla f$ is unattainable, but it can be approximated by inexact function values (zero-order information), denoted by $\tilde{f}$.
+In many such scenarios, we can assume $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq b_{f}$.
+Below, we discuss a deterministic and stochastic zero-order method that generates inexact gradient estimates with quantifiable error bounds [4].
+Gradient estimation via forward finite difference.
+Let $\mathbf{u}_{i}\in\mathbb{R}^{d}$ denote the unit vector with the $i$-th element equal to 1, and $l>0$ be the finite difference interval. Define the $i$-th element of the gradient estimator as
+$\tilde{\nabla}f(\mathbf{x})]_{i}=\frac{\tilde{f}(\mathbf{x}+l\mathbf{u}_{i})-\tilde{f}(\mathbf{x})}{l}.$
+When $f(\mathbf{x})$ has Lipschitz continuous gradient, the error of the gradient estimate is shown by [4] to be bounded as
+$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\frac{\sqrt{d}Ll}{2}+\frac{2\sqrt{d}b_{f}}{l},$
 which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$.
 Note that the forward finite difference requires querying the zero-order oracle $d+1$ times to approximate the gradient at $\mathbf{x}$.
-Another common approach for gradient approximation from (inexact) function values is based on Gaussian smoothing.
-This method randomly samples directions for gradient estimation; hence, the sample size can be some arbitrary number.
-Example 1.4 (Gradient estimation via Gaussian smoothing).
-Let $\{\mathbf{v}_{i}\}_{i=1}^{n}$ be i.i.d. random directions following the standard multivariate normal distribution, and define the gradient estimator as
-~$\tilde{\nabla}f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{f}(\mathbf{x}+l\mathbf{v}_{i})-\tilde{f}(\mathbf{x})}{l}\mathbf{v}_{i}.$
+Gradient estimation via Gaussian smoothing.
+Let $\{\mathbf{v}_{i}\}_{i=1}^{n}$ be i.i.d. random directions following the standard multivariate normal distribution, and define
+$\tilde{\nabla}f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{f}(\mathbf{x}+l\mathbf{v}_{i})-\tilde{f}(\mathbf{x})}{l}\mathbf{v}_{i}.$
 It is shown in [4] that
 $\left\|\mathbb{E}_{\{\mathbf{v}_{i}\}_{i=1}^{n}}[\tilde{\nabla}f(\mathbf{x})]-\nabla f(\mathbf{x})\right\|\leq\sqrt{d}Ll+\frac{\sqrt{d}b_{f}}{l},$
-which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$. Furthermore,
-~$\left\|\tilde{\nabla}f(\mathbf{x})-\mathbb{E}_{\{\mathbf{v}_{i}\}_{i=1}^{n}}[\tilde{\nabla}f(\mathbf{x})]\right\|^{2}\leq\frac{3d}{\epsilon n}\left(3\left\|\nabla f(\mathbf{x})\right\|^{2}+\frac{L^{2}l^{2}}{4}(d+2)(d+4)+\frac{4b_{f}^{2}}{l^{2}}\right)$
-holds with probability at least $1-\frac{1}{\epsilon}$.
-In the later probabilistic bound, $h(\boldsymbol{\eta})$ is determined by $l$, $n$ and $b_{f}$ as $\boldsymbol{\eta}=(l,b_{f},n)$.
-This bound contains a constant, a relative error, and an absolute error terms.
-1.2 Related works
-The study on inexact gradient methods (without acceleration) dates back to 1987 [56], where two notions of inexactness were defined, i.e.,
-~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|$ $\leq\delta,\qquad$ (absolute error), (3)
+which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$.
+1.2 Effect of oracle inexactness on acceleration
+A careful review of the literature shows that the effect of oracle inexactness on acceleration is addressed through 1) development of a specific error condition, and 2) analysis of a common (or customized) accelerated method under that error condition with (or in a few cases without) extra assumption.
+A detailed literature review on accelerated methods under different gradient error conditions is relegated to Section 1.4.
+This paper considers gradient inexactness in the form of the absolute error defined as [48]
 […]
-Below, we review the literature in stochastic setups for non-accelerated methods and (in more detail) in deterministic setups for accelerated gradient-based methods.
-Stochastic non-accelerated methods under biased oracles.
-The above gradient inexactness assumptions have specifically been considered in stochastic setups where ~$\tilde{\nabla}f(\mathbf{x})$ is a biased stochastic estimate of $\nabla f(\mathbf{x})$.
-[1] studies the performance of the biased stochastic gradient descent (SGD) method under both absolute and relative bias.
-[36] considers setups where lower bias and variance require higher computational cost and investigates their tradeoff in stochastic optimization.
-Similarly, [35] analyzes bias-variance tradeoff in the convex bandit setting.
-[7] analyzes biased stochastic gradient-based methods under the absolute bias oracles with and without bias-variance tradeoff.
-[14] (and references therein) summarizes different bias assumptions and introduces a hybrid approach integrating absolute and relative biases.
-[47] considers biased gradient estimates with absolute error and proposes adaptive algorithms that adaptively set the bias throughout the iterations of the algorithms.
-Accelerated gradient methods with inexact gradient oracles.
-Nesterov-type acceleration techniques [51] with inexact gradient oracles have been analyzed under different inexact assumptions.
-Below, these assumptions and their connections with the absolute error (which is the focus of this paper) are reviewed; moreover, the convergence performance of deterministic accelerated methods with inexact gradients are analyzed under the corresponding assumptions.
-An inexact generalization of Nesterov’s Fast Gradient Method (FGM), namely, iFGM, is presented in Algorithm 1 which obtains accelerated rates of convergence in convex setups.
-To simplify the exposition, we only consider the merely convex problems and we set the dual update step as gradient descent with projection.
-Algorithm 1 Inexact Fast Gradient Method (iFGM)
-0:  $\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $0<\alpha_{0}\leq 1$, $A_{0}=\alpha_{0}$.
-1:  for $k=0,\cdots,K-1$ do
-2:     $\mathbf{y}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k}))$
-3:     $\mathbf{z}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
-4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}^{2}\leq A_{k+1}$
-5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
-6:  end for
-d’Aspremont [11] analyzes the performance of iFGM to minimize Liptshitz smooth objective function over compact convex sets $\mathcal{Q}$, with the inexact gradient satisfying
-|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\delta,\quad\forall\mathbf{x},\mathbf{y},\mathbf{z}\in\mathcal{Q},$
-and shows the sequence generated by iFGM satisfies
-$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{A_{K}}+3\delta.$
-This inexactness assumption can be derived from the absolute error assumption for the bounded feasible region with diameter $D$ as
-|~$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq\bar{b}\left\|\mathbf{y}-\mathbf{z}\right\|\leq\bar{b}\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|=\bar{b}D\triangleq\delta,$
-which shows the dependence of the error term on the diameter of the feasible region.
-The notion of first-order $(\delta,L)$-oracle is proposed in [16], which removes the requirements of boundedness of the feasible set and differentiability of the objective function with a generalized definition of convexity and Lipschitz smoothness.
-More specifically, [16] assumes that the oracle provides an inexact function and subgradient pair $(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x}),\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}))$ that satisfies
-$0\leq f(\mathbf{y})-(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}+\delta_{\mathbf{x}}.$
-With $(\delta,L)$-oracle, [16] shows the sequence generated by iFGM satisfies
+Motivated by the application discussed in Section 1.1, (AE) can be considered as one of the most common forms to quantify the gradient error.
+Furthermore, as discussed in Section 1.4, obtaining other error conditions from (AE), e.g., BIE, IFO, etc., requires extra assumptions that limit their applications.
+Nesterov’s Fast Gradient Method (FGM) and its variations have been extensively studied under various inexact gradient assumptions, which we refer to as iFGM.
+The standard FGM update generates sequences $\{\mathbf{y}_{k}\}$ and $\{\mathbf{z}_{k}\}$, with the gradient evaluated at a convex combination of these points.
+The detailed update rule is
+$\begin{cases}\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k}),\\ \mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{1}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k}),\\ \mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},\end{cases}$
+where $\alpha_{0}=1$, $\alpha_{k+1}=\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2}$, and $\tilde{\nabla}$ denotes the inexact gradient.
+When the exact gradient is available, FGM achieves the optimal convergence rate of $\mathcal{O}(K^{-2})$.
+However, when the gradient is inexact, this fast convergence also accelerates error accumulation, undermining the method’s performance.
+To address this, a generalized version of iFGM (iGFGM, Algorithm 1) is often considered.
+This approach controls the increase of $\alpha_{k}$ to obtain a better tradeoff between the convergence rate and the accumulated error.
+For instance, under a specific assumption of the gradient inexactness (see details in Section 1.4), Devolder et al.
+[12] established the convergence result of iGFGM as
+$\mathcal{O}\left(\frac{1}{A_{K}}+\frac{\sum_{i=0}^{K}A_{i}\delta_{\mathbf{x}_{i}}}{A_{K}}\right),$
+where $\delta_{x_{i}}$ represents the gradient error at $x_{i}$ and $A_{i}=\sum_{j=1}^{i}\alpha_{j}$ with $\alpha_{i}^{2}\leq A_{i}$.
+While iGFGM has been widely investigated under error conditions other than (AE) (except for one work [58] which is discussed in Section 1.4), the Optimized Gradient Method (OGM) with an inexact gradient, i.e., iOGM, has received less attention.
+A similar tradeoff between convergence rate and accumulated error exists for iOGM.
+In this paper, we consider the generalized version [35], which we term iGOGM (Algorithm 2).
+We note that the difference between the iGFGM and iGOGM methods is in their step 3, where iGOGM’s stepsize is two times larger.
+When $\lambda_{k}=1$, the $\mathbf{x}_{k+1}$ updates (step 5) in both iGFGM and iGOGM are simplified as $\mathbf{x}_{k+1}=\left(1-\frac{1}{\alpha_{k+1}}\right)\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},$ and algorithms reduce to iFGM and iOGM, respectively.
+[…]
+1.3 Contributions
+In this paper, we analyze nonasymptotic convergence bounds of two accelerated gradient methods, namely Generalized Fast Gradient Method (iGFGM, Algorithm 1) and Generalized Optimized Gradient Method (iGOGM, Algorithm 2), under inexact gradient oracles satisfying the absolute error condition (AE).
+The (AE) error bound finds a range of applications, a few of which are discussed in Section 1.1.
+These bounds are established through the Performance Estimation Problem (PEP) technique and, unlike previous works, are independent of unknown quantities–see the first bullet below for more details.
+Furthermore, our analysis allows variable error along the iterations of the algorithms; hence, the established bounds allow exploiting the tradeoff between the per-iteration cost to control the bias and the total cost to obtain the optimal oracle inexactness schedule.
+Our key contributions can be summarized as follows:
+  • As summarized in Table 1, existing methods are limited by their reliance on strong assumptions (e.g., BIE, IFO, IFO-q, and RE) which are generally not easily verifiable in different applications.
+Furthermore, some previous analyses depend on unquantifiable terms, e.g., the dependence of the bound in [58] on unquantifiable $\tilde{R}_{K}$, the radius of the set containing the algorithm’s trajectory.
+Hence, it is not possible to evaluate the accumulated error in [58] and to determine the algorithm’s optimal parameter settings.
+Our work closes this gap by deriving a quantifiable expression for the accumulated error under the AE condition.
+  • We provide the convergence bound of iGOGM and iGFGM under the absolute error (AE) condition in Section 2.1.
+The derived convergence bounds consist of two components: the diminishing component independent of the oracle’s error and the component containing the accumulated error.
+Interestingly, the accumulated error is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$, and is determined solely by the Lipschitz constant and the stepsize.
+This result advances the convergence bound of [58], presented in (4), in the sense that it eliminates the dependence on the unknown parameter $\tilde{R}_{k}$.
+Furthermore, unlike Nabou et al.
+[41]’s inexactness condition for the gradient to be a subgradient, we do not require such a condition, and we yet achieve comparable convergence guarantees.
+  • In Section 2.2, we analyze the established convergence bounds from two different perspectives: i) we investigate the tradeoff between the vanishing term and the accumulated error; ii) we establish the optimal inexactness schedule, i.e., the minimal total cost to control the error, while preserving the accelerated convergence bound.
+The proofs of our main convergence results for iGOGM and iGFGM are presented in Sections 4 and (A.2), respectively.
+  • We derive an analytical feasible solution to the dual of the relaxed semidefinite programming formulation of the Performance Estimation Problem (PEP) for first-order algorithms with inexact (AE) oracles in Section 4.
+The approach used to find this solution inspires our theoretical proof of the convergence bound.
+Notably, this proof can be understood without prior knowledge of the PEP technique and could be of independent interest.
+The codes for our numerical experiments as well as those to verify the algebraic steps in our proofs are available at https://github.com/Yin-LIU/Inexact-Acceleration-PEP.
+Notations
+Vectors and matrices are denoted by bold and capitalized bold letters, respectively.
+Sets are denoted by calligraphic letters.
+In the absence of additional instruction, a bold letter coupled with an arrow, e.g., $\vec{\mathbf{u}}$, generally denotes a standard basis vector, which has exactly one element equal to 1 while other elements are 0.
+$\mathbf{M}_{i,j}$ stands for the $(i,j)$-th element of the matrix $\mathbf{M}$.
+The notation $\mathbf{M}^{\backslash[i]}$ refers to the submatrix of $\mathbf{M}$ obtained by removing its $i$-th column and row.
+$\mathbf{M}]_{n\times m}$ indicates the shape of the matrix and $\mathbf{M}]$ is used to emphasize the item inside the brackets is a matrix.
+$\left\|\cdot\right\|$ denotes the $l_{2}$ norm for vectors and Frobenius norm for matrices.
+$\log(\cdot)$ represents the natural logarithm.
+1.4 Related work
+Below, we discuss the literature related to accelerated methods under various oracle inexactness conditions.
+Bounded Inner Product Error (BIE)  d’Aspremont [8] focuses on smooth convex optimization over a compact convex set $\mathcal{Q}$, with the inexact gradient satisfying
+[…]
+The BIE condition can be derived from the absolute error assumption when the feasible set is bounded.
+Specifically, if the gradient error satisfies $\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq b$ for some constant $b$, then
+$\left|\left\langle\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{z}\right\rangle\right|\leq b\left\|\mathbf{y}-\mathbf{z}\right\|\leq b\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|=b\cdot D\triangleq\delta,$
+where $D=\max_{\mathbf{y},\mathbf{z}\in\mathcal{Q}}\left\|\mathbf{y}-\mathbf{z}\right\|$ represents the diameter of the feasible set.
+This derivation highlights that the error bound $\delta$ is directly proportional to the size of the feasible region.
+Under the BIE assumption, iFGM has the convergence rate of
+$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{A_{K}}+3\delta.$
+Note that the convergence rate is determined by two components: the first term decreases with the accumulation parameter $A_{K}$, and the second term is a constant that scales with the inner product error bound $\delta$.
+Inexact First-order $(\delta,L)$ Oracle (IFO)  Devolder et al.
+[12] introduce the $(\delta,L)$ first-order oracle condition for problems with unbounded feasible set or those with nonsmooth objective functions.
+The inexactness of the oracle pair $(\tilde{f}_{\delta_{\mathbf{x}}}(\mathbf{x}),\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x}))$ is quantified by
+[…]
+which is a relaxation of the first-order convexity and Lipschitz smoothness conditions.
+Under the IFO assumption, Devolder et al. [12] shows that the sequence generated by iFGM satisfies
 $f(\mathbf{y}_{K+1})-f_{*}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{A_{K}}+\frac{\sum_{i=0}^{K}A_{i}\delta_{\mathbf{x}_{i}}}{A_{K}}.$
-This result allows different inexactness levels for different iterations.
-However, to use the above results for an oracle that satisfies the absolute error assumption, the boundedness of the feasible region is needed [16], i.e., an oracles with absolute error assumption $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq\Delta_{1}$ and ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\Delta_{2}$, $\tilde{f}_{\delta_{\mathbf{x}}}=\tilde{f}(\mathbf{x})-\Delta_{1}-\Delta_{2}D$, ~$\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x})=\tilde{\nabla}f(\mathbf{x})$, and $\delta_{\mathbf{x}}=2\Delta_{1}+2\Delta_{2}D$ (with $D$ being the feasible region diameter) is a $(\delta,L)-$oracle.
-To overcome the boundedness drawback of the $(\delta,L)$-oracle, [49] proposes the inexact first-order oracle of degree $q$ as
-$0\leq f(\mathbf{y})-(f(\mathbf{x})+\left\langle\tilde{\nabla}f_{\delta}(\mathbf{x}),\mathbf{y}-\mathbf{x}\right\rangle\leq\frac{L}{2}\left\|\mathbf{y}-\mathbf{x}\right\|^{2}+\delta\left\|\mathbf{x}-\mathbf{y}\right\|^{q},$
-with $q\in[0,2)$.
-With the appropriate selection of the $q$ and $\delta$ parameters, any inexact gradient that satisfies the absolute error condition satisfies the right-hand-side inequality above without the requirement of the bounded feasible region.
-Under this new oracle assumption, the sequence generated by iFGM satisfies
-$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4(L+q\rho)\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{(K+3)(2-q)\delta^{\frac{2}{2-q}}}{2\rho^{\frac{q}{2-q}}},$
-with $\rho>0$ be any fixed number. Minimizing the right-hand-side with respect to $\rho$, the convergence rate becomes
-$f(\mathbf{y}_{K+1})-f_{*}\leq\frac{4L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+1)(K+2)}+\frac{8^{\frac{q}{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{q}(K+3)}{((K+1)(K+2)(K+3))^{\frac{q}{2}}}\delta.$
-This latter condition allows unbounded regions; however, its left-hand-side inequality requires the inexact gradient to be a subgradient, which can not be guaranteed under the absolute error assumption.
-Furthermore, in the convergence bound, the error term depends on the initial condition.
-Finally, the analysis is provided for fixed inexactness levels.
-Another accelerated algorithm studied under inexact oracles is the Similar Triangles Method (STM) [27]– see the inexact version in Algorithm 2.
-Compared to FGM, it only requires one projection for each iteration and, hence, it is preferable for constrained problems.
-The convergence properties of iSTM are studied under both absolute and relative errors.
-Algorithm 2 Inexact Similar Triangles Method (iSTM)
-0:  $\mathbf{z}_{-1}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $\alpha_{-1}=A_{-1}=0$.
-1:  for $k=0,\cdots,K$ do
-2:     $A_{k}=A_{k}+\alpha_{k}$ with $\alpha_{k}^{2}\leq A_{k}$
-3:     $\mathbf{x}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k-1})$
-4:     $\mathbf{z}_{k}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
-5:     $\mathbf{y}_{k}=\frac{1}{A_{k}}(A_{k-1}\mathbf{y}_{k-1}+\alpha_{k}\mathbf{z}_{k})$
-6:  end for
-When the feasible region $\mathcal{Q}=\mathbb{R}^{d}$, i.e., no projection is required, iSTM can be formulated as iFGM.
-From $A_{k}=A_{k-1}+\alpha_{k}$, we can see that the $\mathbf{x}_{k}$ updates in iFGM and iSTM are the same with different indices since iFGM updates $\mathbf{x}_{k}$ at the end of each iteration while iSTM updates it at the beginning of each iteration.
-The $\mathbf{z}_{k}$ updates for the two algorithms are the same.
-Hence, we only need to consider the $\mathbf{y}_{k}$ updates.
-From step 3 of iSTM, $A_{k-1}\mathbf{y}_{k-1}=A_{k}\mathbf{x}_{k}-\alpha_{k}\mathbf{z}_{k-1}$.
-Inserting this identity into step 5, and by the update of step 4, we have
-$\mathbf{y}_{k}=\mathbf{x}_{k}+\frac{\alpha_{k}}{A_{k}}(\mathbf{z}_{k}-\mathbf{z}_{k-1})=\mathbf{x}_{k}-\frac{\alpha_{k}^{2}}{A_{k}L}\tilde{\nabla}f(\mathbf{x}_{k}).$
-When $\alpha_{k}^{2}=A_{k}$, the above update is equivalent to the step 2 in iFGM.
-Furthermore, [2] analyzes iSTM under the inexact gradient with absolute error bound. The sequence generated by iSTM with $\alpha_{k}^{2}=A_{k},\forall k$, satisfies
-$f(\mathbf{y}_{K})-f_{*}\leq\frac{8L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{K^{2}}+3\tilde{R}_{K}\delta+\frac{K\delta^{2}}{2L},$
-with $\tilde{R}_{K}\triangleq\max_{0\leq j\leq K}\{\left\|\mathbf{z}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{x}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{y}_{j}-\mathbf{x}_{*}\right\|\}$.
-This convergence bound is only given under fixed inexactness levels and restricts the stepsize as $\alpha_{k}^{2}=A_{k}$.
-Another drawback is that $\tilde{R}_{K}$ is not exactly quantifiable; indeed, [2] only shows its boundedness by $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ under the exact gradient oracle.
-The boundedness of $\tilde{R}_{K}$ is shown by the same group under the relative error assumption with a relaxed stepsize [44].
-It is shown that the sequence generated by iSTM with $\alpha_{k}=\frac{(k+2)^{p-1}}{2\alpha}$, $p\in[1,2]$, and $\alpha=\mathcal{O}(\max\{1,K^{\frac{p}{4}}\sqrt{\delta},K^{\frac{p}{2}}\delta,K^{p}\delta^{2}\})$ satisfies
-$f(\mathbf{y}_{K})-f_{*}\leq\frac{16\alpha L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+2)^{p}}.$
-The paper shows $\tilde{R}_{K}$ is bounded by $2\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ when the stepsize is properly chosen and, hence, there is no separate error term in the convergence rate.
-Besides the abovementioned works, several researchers have proposed alternative error assumptions and results.
-The result of iFGM for the strongly convex setup is provided in [17].
-[8] considers a variation of iFGM with the absolute error setting.
-To exploit the tradeoff between the convergence rate and accumulated error, the intermediate gradient method is proposed in [18] which is closely related to iFGM.
-This method is further generalized for other setups in [25, 19, 60, 28, 61].
-[26] and [3] analyze nonaccelerated and accelerated algorithms under relative and absolute errors, respectively, with the latter focusing on the tradeoff between robustness to error and convergence rate–see also [32].
-[24, 45] assume the error to satisfy the ~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq B\left\|\mathbf{x}-\mathbf{x}_{*}\right\|$ condition, which can be derived from the relative error bound.
-[24] provides results for stochastic proximal inexact gradient methods with and without acceleration for strongly convex problems, while [45] studies generalized monotone variational inequality problems.
-Other notable contributions include the analysis of absolute and relative inaccuracy for proximal point methods in [57], the study of proximal gradient methods with absolute error in [59], the inexact proximal gradient method for weakly convex functions with absolute error in [37], and minimization of the accumulated error when the total allocation of error-control resources is fixed [15].
-1.3 Contributions
-This paper employs Performance Estimation Problem (PEP) techniques to primarily analyze the convergence rate of the Generalized Optimized Gradient Method with an inexact gradient oracle (iGOGM).
-Our key contributions are as follows:
-  • We derive an analytical solution to the relaxed PEP optimization problem for iGOGM under variable oracle inexactness levels.
-The approach used to find this solution inspires our theoretical proof of the convergence.
-Notably, this proof can be understood without prior knowledge of PEP techniques.
-  • The derived convergence bound consists of two components: the exact convergence rate and the accumulated error.
-Significantly, the error term is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$, and is determined solely by the Lipschitz constant and the step size.
-  • We provide a comprehensive analysis of the convergence bound from two perspectives: i) We investigate the tradeoff between the convergence rate and the accumulated error.
-ii) We establish optimal inexactness levels that preserve the method’s convergence rate.
-  • We also provide the PEP-inspired convergence bound for the Generalized Fast Gradient Method with inexact gradient oracle (iGFGM).
-We also provide the codes for our numerical experiments as well as the verification of the algebraic steps in our proofs.
-These codes are available at https://github.com/Yin-LIU/Inexact-Acceleration-PEP.
-Notations
+This result is similar to that of (BIE) condition, i.e., a term decreasing with iteration and an accumulated error term.
+However, (IFO) condition also allows varying inexactness levels along the iterations.
+While (IFO) condition allows an unbounded feasible set, obtaining it from the absolute error condition requires bounded feasible sets.
+Specifically, for an oracle with absolute errors $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq\Delta_{1}$ and $\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\Delta_{2}$.
+By defining $\tilde{f}_{\delta_{\mathbf{x}}}\triangleq\tilde{f}(\mathbf{x})-\Delta_{1}-\Delta_{2}D$ and $\tilde{\nabla}f_{\delta_{\mathbf{x}}}(\mathbf{x})\triangleq\tilde{\nabla}f(\mathbf{x})$, where $\delta_{\mathbf{x}}=2\Delta_{1}+2\Delta_{2}D$ (with $D$ representing the feasible region’s diameter), it can be shown the oracle with (AE) condition is a $(\delta,L)$-oracle [12].
+Inexact First-order $(\delta,L)$ Oracle of Degree $q$ (IFO-q) To be able to address oracle inexactness in minimization of smooth nonconvex or nonsmooth convex objectives, while sacrificing the inexact gradients to be subgradients, Nabou et al.
+[41] propose the IFO-q oracle with an additional degree parameter $q\in[0,2)$ as
 […]
-2 Performance Estimation Problems (PEP)
-2.1 PEP overview
-The tool used for analyzing the convergence rate of the inexact accelerated methods is known as Performance Estimation Problem (PEP).
-This technique is proposed by Drori and Teboulle [22] which uses an optimization problem to find the worst-case performance when a given algorithm is applied to a class of functions.
+With an appropriate selection of the parameters $q$ and $\delta$, any inexact gradient satisfying the absolute error condition also satisfies the upper bound IFO-q condition even on unbounded feasible sets.
+However, the lower bound of this inequality requires the inexact gradient to be a subgradient of the objective function.
+This subgradient condition is not inherently guaranteed by the absolute error assumption, limiting the applicability of the IFO-q oracle when solely relying on absolute gradient errors.
+The convergence rate of iFGM with an oracle satisfying IFO–q is given as
+[…]
+where both terms in the upper bound depend on the initial condition.
+Absolute Error (AE) The absolute error condition (AE), has recently been investigated in a couple of works.
+Notably, Vasin et al.
+[58] study a variant of the accelerated algorithm known as the Similar Triangles Method (STM) [22] with inexact gradients, i.e., (iSTM).
+In contrast to FGM, STM requires only one projection per iteration, making it preferable for constrained problems.
+For unconstrained optimization, it is shown by Vasin et al.
+[58] that iSTM achieves the convergence rate of
+[…]
+where $\tilde{R}_{K}\triangleq\max_{0\leq j\leq K}\big\{\left\|\mathbf{z}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{x}_{j}-\mathbf{x}_{*}\right\|,\left\|\mathbf{y}_{j}-\mathbf{x}_{*}\right\|\big\}$.
+However, this convergence bound is only given under fixed inexactness level along the iterations and restricts the stepsize as $\alpha_{k}^{2}=A_{k}$.
+Furthermore, $\tilde{R}_{K}$ is not explicitly quantifiable under the absolute error assumption, since it is only shown to be bounded by $\|\mathbf{x}_{0}-\mathbf{x}_{*}\|$ under the exact gradient oracle.
+When the feasible region is bounded with diameter $D$, one can set $\tilde{R}_{K}=D$, thereby recovering the rate similar to previous cases.
+This coincides with the observation that both the BIE and IFO-($\delta,L$) conditions can be derived from (AE) assumption under a bounded feasible region.
+Relative Error (RE) [48] The relative error (RE) is a stronger assumption as it requires the error to decrease with the gradient norm and enforces the gradient oracle to be more accurate near the stationary points in smooth unconstrained optimization.
+[…]
+Under this condition, Kornilov et al.
+[37] establish the boundedness of $\tilde{R}_{K}$ in (4) for the iSTM algorithm.
+Specifically, they show that with appropriate step size selection, $\tilde{R}_{K}$ can be bounded by $2\|\mathbf{x}_{0}-\mathbf{x}_{*}\|$, eliminating the need for separate error terms in the convergence rate.
+The resulting convergence rate is
+$f(\mathbf{y}_{K})-f_{*}\leq\frac{16aL\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{(K+2)^{p}},$
+where $p\in[1,2]$ and $a=\mathcal{O}\left(\max\left\{1,K^{\frac{p}{4}}\sqrt{\delta},K^{\frac{p}{2}}\delta,K^{p}\delta^{2}\right\}\right)$ is a parameter that depends on the relative error level $\delta$ and the iteration count $K$.
+The inexact assumptions discussed above and their corresponding convergence rates are summarized in Table 1.
+Table 1: Summary of gradient inexactness error conditions and the corresponding complexity bounds for accelerated methods.
+(BFS): Bounded Feasible Set; (BG): Bounded Gradient; (SG): inexact gradient being a SubGradient.
+The $u_{k}$ in “this work” are functions of the stepsize sequence $\{\alpha_{i}\}_{i=1}^{K}$.
+Error condition (AE) Allows changing error Iteration complexity
+along iterations?
+(BIE) [8] $\Leftarrow$ (BFS)+ $\times$ $\mathcal{O}\left(\frac{1}{A_{K}}+\delta\right)$
+(IFO) [12] $\Leftarrow$(BFS)+ $\mathcal{O}\left(\frac{1}{A_{K}}+\frac{\sum_{i=0}^{K}A_{i}\delta_{\mathbf{x}_{i}}}{A_{K}}\right)$
+(IFO-q[41] $\Leftarrow$(SG)+ $\times$ $\mathcal{O}\left(\frac{1}{K^{2}}+\frac{\delta}{K^{3q/2-1}}\right)$
+[…]
+(AE) [58] / $\times$ $\mathcal{O}\left(\frac{1}{K^{2}}+\tilde{R}_{K}\delta+K\delta^{2}\right)$
+(AE)(this work) / $\mathcal{O}\left(\frac{1}{A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}\right)$
+Besides the aforementioned works, accelerated methods with inexact gradient oracles have also been discussed in other settings.
+The iFGM for the strongly convex setup is presented in [13], while [5] explores a variation of iFGM under the absolute error condition.
+To leverage the tradeoff between convergence rate and accumulated error, the “Intermediate Gradient Method” is introduced in [14], which is closely related to iFGM.
+This approach was subsequently generalized for various settings, see [20, 15, 52, 23, 53].
+To analyze the effect of gradient inexactness on the convergence bound, Gannot [21] and Aybat et al.
+[1] examine nonaccelerated and accelerated algorithms under relative and absolute errors, respectively.
+The latter specifically addresses the balance between robustness to error and convergence rate–see also [27].
+Other notable contributions include the analysis of absolute and relative inaccuracy for proximal point methods in [49], the study of proximal gradient methods with absolute error in [51], the inexact proximal gradient method for weakly convex functions with absolute error in [30], and the minimization of accumulated (controllable) error for optimal computational resource allocation in [11].
+2 Main results
+This section contains our main theoretical results.
+Section 2.1 provides nonasymptotic convergence bounds of GOGM and GFGM algorithms with inexact gradient oracles.
+Section 2.2 exploits the established convergence bounds to obtain the optimal tradeoff between the convergence rate and accumulated error and to obtain the optimal inexactness schedule along the algorithm iterations to minimize the total $\eta$-complexity, i.e., $\sum_{k=1}^{K}\eta_{k}$ in (2) when $r=1$.
+2.1 Main results for iGOGM and iGFGM
+The primary assumptions of the paper are summarized below:
+Assumption 2.1.
+The objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup+\infty$ is proper, closed, convex, and $L$-Lipschitz smooth.
+The gradient estimate defined as $\tilde{\nabla}f(\mathbf{x})=\nabla f(\mathbf{x})+\mathbf{e}_{\mathbf{x}}$ satisfies the absolute error (AE) condition
+$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|^{2}\leq b_{\mathbf{x}}^{2},$
+where $b_{\mathbf{x}}$ is an arbitrary constant for each $\mathbf{x}$.
+The theorem below provides the nonasymptotic convergence bound for the iGOGM algorithm. The optimality measure is discussed in Remark 2.3.
+Theorem 2.2.
+Under Assumption 2.1, the sequence generated by iGOGM (Algorithm 2) satisfies
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
+with $u_{k}=\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}$.
+Remark 2.3.
+We use the same optimality measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ used in the analysis of exact GOGM, as it is still a meaningful one for the inexact problem.
+For iGFGM and iGOGM, we have
+$f(\mathbf{y}_{K+1})=$ $f\left(\mathbf{x}_{K}-\frac{1}{L}(\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K})\right)$
+$\leq$ $f(\mathbf{x}_{K})-\frac{1}{L}\left\langle\nabla f(\mathbf{x}_{K}),\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\rangle+\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\|^{2}$
+$=$ $f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}+\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}.$
+Comparing to the measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$, the extra term $\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}$ can be discarded since the constraint $\left\|\mathbf{e}_{K}\right\|^{2}\leq b_{K}^{2}$ is the only one that involves $\mathbf{e}_{K}$ and the PEP solution sets $\left\|\mathbf{e}_{K}\right\|^{2}=b_{K}^{2}$.
+Since the algorithm only generates $\{\mathbf{x}_{k}\}$ and $\{\mathbf{y}_{k}\}$ up to $k=K$, $\mathbf{y}_{K+1}$ is a dummy variable used for the theoretical guarantee, and setting $b_{K}=0$ does not affect the complexity of the algorithm.
+The proof sketch of Theorem 2.2 is provided below.
+The proof contains a couple of crucial but non-intuitive steps, motivated by the PEP analysis, provided in Section 3.2.
+The detailed proof is provided in Section 4.1.
+Proof sketch.
+The core strategy is to combine inequalities derived from Lipschitz smoothness and convexity.
+By rearranging terms, applying specific multipliers, and creating a telescoping sum, we can isolate and bound the desired optimality measure.
+This process is guided by the analytical feasible solution identified in Section 4.
+The key steps are as follows.
+First, we select two fundamental inequalities and assign their respective coefficients:
+[…]
+[…]
+Summing these inequalities weighted by their multipliers and rearranging the terms to move $f(\mathbf{x}_{K})-f_{*}$ on the left-hand side yields a key intermediate inequality.
+Subtracting $\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ from both sides gives us the optimality measure we seek to bound, which has the form
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $-\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
+$-\frac{1}{2L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}.$
+By substituting the update rule of the iGOGM algorithm for $\mathbf{x}_{k}$, the inequality becomes
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
+$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle.$
+Next, our goal is to eliminate the inner products with the negative quadratic terms.
+To achieve this, we express the inner products as a difference of squared norms.
+For instance, the first inner-product can be bounded using the equality
+$\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle$
+$=$ $\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\frac{1}{4\tau}\left\|\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}.$
+After some algebraic manipulation and setting $\tau=\frac{L}{4A_{K}}$, we obtain
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle.$
+We continue this procedure for the remaining inner-product terms involving the error vectors $\mathbf{e}_{k}$. This is accomplished by introducing the quadratic term
+$\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}.$
+As this step is algebraically intensive, we omit the detailed derivation and present the final inequality as
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}\left(\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
+The theorem below provides the nonasymptotic convergence bound for the iGFGM algorithm. The proof is provided in Section A.2.
+Theorem 2.4.
+Under Assumption 2.1, the sequence generated by iGFGM (Algorithm 1) satisfies
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{2A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
+with $u_{k}=\frac{A_{k}^{2}(1+\alpha_{k+1})}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2LA_{K}(2A_{i}-\alpha_{i}^{2})}$.
+2.2 Exploiting the rate-error tradeoff and finding the optimal inexactness schedule
+Based on the convergence bound of the iGOGM algorithm, this section first looks into the tradeoff between the convergence rate and accumulated error by changing the stepsize given a fixed gradient inexactness along iterations.
+Second, we aim to optimize the inexactness levels along the iterations, for a given stepsize, so the total cost of the oracle that controls the inexactness is minimized.
+Similar analysis follows for the iGFGM algorithm, but it is not included for brevity.
+2.2.1 Convergence rate and the accumulated error tradeoff
+Considering the bound $\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$, with ${A_{k}},{u_{k}}$ being functions of $\alpha_{k}$}, one can propose to minimize the bound with respect to $\alpha_{k}$}.
+However, we note that such a problem is a high-order polynomial optimization and its solution is intractable in general.
+Even if one is satisfied with the numerical solution to find the optimized algorithm, it is preferable to solve the minimax problem (13) numerically following the relaxation approach in [18] as calculated in Appendix B.
+The resulting optimized stepsize sequence and convergence bound are better than the solution of the polynomial optimization problem as the derived theoretical $u_{i}$} in Lemma 4.1 is not necessarily the optimal solution.
+We narrow the iGOGM down to iOGM-a, with the stepsize $\alpha_{i}=\frac{i+a}{a}$ and $A_{k}=\sum_{i=0}^{k}\frac{i+a}{a}=\frac{(k+2a)(k+1)}{2a}$.
+The condition $A_{k}>\alpha_{k}^{2}$ holds for any $k$ when $a>2$.
+Note that when $K\gg a$, the convergence rate is still $\mathcal{O}(K^{-2})$ in the exact case.
+Furthermore, [35] shows that the exact algorithm OGM-a has the asymptotic worst-case bound for function value as $\frac{a}{2}K^{-2}$ and the smallest gradient norm as $\frac{a\sqrt{6}}{2\sqrt{a-2}}K^{-1.5}$; it achieves the best performance for both measures with $a=4$.
+The simplicity of the stepsize as well as its good performances over function value and gradient norm makes OGM-a a proper choice under exact oracles, and worth further study under inexact oracles as well.
+Replacing the stepsizes of OGM-$a$ in Theorem 2.2, we have
+$u_{k}=\frac{1}{4La(K+1)(K+2a)((a-2)k+2a^{2}-2a-2)}\left((2a+8)k^{4}+(11a^{2}+48a+16)k^{3}\right$.
+$\left.+(20a^{3}+107a^{2}+70a+8)k^{2}+(12a^{4}+104a^{3}+100a^{2}+24a)k+36a^{4}+48a^{3}+16a^{2})\right)$
+$+\sum_{i=k+1}^{K-1}\frac{(k+a)(2i^{3}+(9a+4)i^{2}+(13a^{2}+13a+2)i+6a^{3}+10a^{2}+4a}{La(K+1)(K+2a)((a-2)i+2a^{2}-2a-2)}.$
+With $a=4$, as argued above, we have
+$u_{k}=$ $\frac{16k^{4}+384k^{3}+3280k^{2}+11424k+12544}{16L(K+1)(K+8)(2k+22)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(2i^{3}+40i^{2}+262i+560)}{4L(K+8)(K+1)(2i+22)}$
+$=$ $\frac{(k+7)^{2}(k+8)(k+2)}{2L(K+8)(K+1)(k+11)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(i+8)(i+7)(i+5)}{4L(K+1)(K+8)(i+11)}.$
+Replacing $k+11$ with $k+8$ and $i+11$ with $i+8$ in the two denominators, we have a simpler expression with slightly bigger $u_{k}$ as
+$u_{k}=\frac{(k+7)^{2}(k+2)}{2L(K+8)(K+1)}+\frac{(k+4)(K-k-1)(2K^{2}+2Kk+35K+2k^{2}+37k+210)}{24L(K+8)(K+1)}.$
+Assume $b_{k}\equiv\bar{b}$, then the accumulated error is calculated as
+$\bar{b}^{2}\sum_{k=0}^{K-1}u_{k}=\frac{\bar{b}^{2}K(12K^{3}+303K^{2}+2687K+8758)}{480L(K+8)}.$
+The above results indicate that the accumulated error is in $\mathcal{O}(\frac{\bar{b}^{2}K^{3}}{L})$.
+For any arbitrary $a$ values, it is difficult to derive a closed form of $u_{k}$ since the denominator can only be simplified as $k+c$ structure given a specific value for $a$.
+Hence, we provide a numerical result of $u_{k}$ for a given $K$ with different $a$ in Figure 1.
+For the same $a$ and $k$, bigger $K$ results in bigger $u_{k}$.
+For each $a$ value, $u_{k}$} first increases and then decreases with $k$.
+For a fixed $k$ and $K$, bigger $a$ results in smaller $u_{k}$ values.
+Figure 1: Values of $u_{k}$ for OGM-$a$ with different values of $a$ and $K$.
+Figure 2: Convergence rate and accumulated error for OGM-$a$ for different $a$ and $K$ values.
+Based on this result, to reduce the accumulated error, one can consider increasing $a$.
+However, this will adversely slow down the convergence rate as $A_{K}=\frac{(K+2a)(K+1)}{2a}$ and decreases when $a$ is increased.
+We provide the comparison of convergence rate and accumulated error with different iteration numbers $K$, and $a$ in Figure 2, where $\tau=\frac{L}{4A_{K}}$.
+The convergence rate $\tau$ is smaller for smaller $a$ while the accumulated error is bigger.
+When $a=4$, $\sum_{k=0}^{K-1}u_{k}$ is between $K^{2}$ and $K^{3}$, better than our relaxed analysis result.
+Another interesting fact is that when $K\ll a$ ($a=10^{5}$ and $a=10^{6}$), the convergence rate $\tau=\mathcal{O}(K^{-1})$ and the accumulate error is $\mathcal{O}(K^{1})$.
+If $K\ll a$, we have $\alpha_{k}=1$, and the proposition below provides the convergence bound of this extreme scenario.
+Proposition 2.5.
+With the stepsize defined as $\alpha_{k}\equiv 1$, the sequence generated by iGOGM (Algorithm 2) satisfies
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\sum_{k=0}^{K-1}\frac{3(2K-k+1)}{4L(K+1)}\left\|\mathbf{e}_{k}\right\|^{2}.$
+Furthermore, when $\left\|\mathbf{e}_{k}\right\|\leq\bar{b},\forall k$, the above inequality becomes
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\frac{9K}{8L}\bar{b}^{2}.$
+2.2.2 Optimal inexactness schedule
+Since the derived $u_{k}$ is not constant, it will be beneficial to set $b_{k}$ based on the value of $u_{k}$.
+Recall the definition of $\mathbb{\eta}$ from (2), when $r=1$, we can minimize the total $\eta$-complexity, i.e., $\sum_{k=0}^{K-1}\eta_{k}$, so that the accumulated error does not exceed the convergence rate as
+[…]
+This is a convex optimization problem if $h$ is convex and decreasing over a convex quadratic constraint, and at its optimal solution, the constraint is binding.
+Hence, the inequality constraint can be replaced with the equality.
+Let $\lambda$ be the dual multiplier of the constraint, the Lagrangian function is
+$\mathcal{L}(\mathbf{b},\lambda)=\sum_{k=0}^{K-1}h^{-1}(b_{k})+\lambda\left(\sum_{k=0}^{K-1}u_{k}b_{k}^{2}-\frac{LR^{2}}{4A_{k}}\right).$
+At the optimal solution $\mathbf{b}^{*}$ and $\lambda^{*}$, from the optimality conditions, we have
+[…]
+[…]
+Solving the above equations, we have a closed-form solution of $\mathbf{b}^{*}$.
+In the following, we give the closed-form of $\mathbf{b}^{*}$ for two examples of $h(\eta)$ [39], which are two common structures of the inexact gradient oracles.
+Lemma 2.6 ($h(\eta)$ with power law decay).
+Let $h(\eta)=c_{1}\eta^{-c_{2}}$ with $c_{1},c_{2}>0$ to be two constants. The inverse function and its gradient are
+$h^{-1}(b)=\left(\frac{c_{1}}{b}\right)^{1/c_{2}}\quad\text{and}\quad\nabla_{b}h^{-1}(b)=-\frac{c_{1}^{1/c_{2}}b^{-(1+c_{2})/c_{2}}}{c_{2}}.$
+Furthermore, the optimal primal and dual solutions of (OPT-b) are
+$b_{k}^{*}=$ $\frac{\sqrt{L}R}{2\sqrt{A_{K}\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}}u_{k}^{d/(1+2c_{2})}},$
+$\lambda^{*}=$ $\frac{c^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
+Proof.
+With power law decay, (5) is written as
+$\frac{c^{1/c_{2}}(b_{k}^{*})^{-(1+c_{2})/c_{2}}}{c_{2}}+2\lambda^{*}b_{k}^{*}u_{k}=0,\quad k=0,1,\cdots,K-1.$
+From this equation, $b_{k}^{*}$ can be calculated as
+[…]
+Inserting this equation into (5), and solve for $\lambda^{*}$ we get
+$\lambda^{*}=\frac{c_{1}^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
+Replacing this solution into (7), the final result can be derived. ∎
+Lemma 2.7 ($h(\eta)$ with exponential decay).
+Let $h(\eta)=q_{1}q_{2}^{-\eta}$ with $q_{1}>0$ and $q_{2}>1$ be two constants. Then
+$h^{-1}(b)=\frac{\log q_{1}-\log b}{\log q_{2}},\quad\nabla_{b}h^{-1}(b)=-\frac{1}{b\log q_{2}},$
+and, the optimal primal and dual solutions are
+$b_{k}^{*}=\sqrt{\frac{LR^{2}}{4KA_{K}u_{k}}},\quad\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
+Proof.
+Similar to the proof of Lemma 2.6, from (5) we have
+$u_{k}(b_{k}^{*})^{2}=\frac{1}{2\lambda^{*}\log q_{2}}.$ (8)
+Inserting this into (6), we have
+$\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
+By the previous two equations, we can derive the final solution. ∎
+To quantify the improvement after the optimized inexactness level, we calculate the total $\eta$-complexity for iOGM-4 algorithm for the two structures of $h(\eta)$ functions.
+We fix $L=1$ and $R=1$, then solve the constant $\bar{b}$ through
+$\bar{b}^{2}\sum_{k=0}^{K-1}\leq\frac{LR^{2}}{4A_{K}},$
+and calculate the optimal $\bar{b}$ by the solutions we derived in Lemmas 2.6 and 2.7.
+In Figure 3, we illustrate the result for power law decay with four different values of $c_{2}$.
+Note that the $\eta$-complexity decreases by increasing $c_{2}$.
+If we focus on the improvement of $\eta$-complexity for the optimized $\{b_{k}\}$, the improvement is in the same order as the complexity of constant $b$, which shows optimizing over $b$ can decrease the $\eta$-complexity significantly.
+As for the exponential decay in Figure 4, the improvement is not as large as that of the power law decay.
+Although the improvement is not as large as that of the power law decay, considering its absolute value, it shows a considerable improvement.
+Figure 3: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with power law decay with $c_{1}=1$ and $L=1,R=1$.
+Figure 4: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with exponential decay with $q_{1}=1,q_{2}=e$ and $L=1,R=1$.
+3 Optimized Gradient Method with inexact gradient oracle
+In this section, we introduce the analytical tool used to derive our quantifiable convergence bounds, i.e., Performance Estimation Problem (PEP).
+Proposed by Drori and Teboulle [18], PEP is a technique that formulates an optimization problem to find the worst-case performance of an algorithm on a given class of functions.
+Inspired by the derivation of the Optimized Gradient Method (OGM), which relies on this tool, our analysis will first focus on iGOGM and then be extended to iGFGM.
+3.1 Performance Estimation Problems (PEP) overview
 To solve the convex and Lipschitz smooth problems with the general first-order method using an exact gradient oracle, the PEP is defined as
-$\max_{\begin{subarray}{c}d,f\\ \mathbf{x}_{*},\mathbf{x}_{0},\cdots,\mathbf{x}_{k}\in\mathbb{R}^{d}\\ \nabla f(\mathbf{x}_{0}),\cdots,\nabla f(\mathbf{x}_{K})\in\mathbb{R}^{d}\end{subarray}}\quad$ $f(\mathbf{x}_{K})-f(\mathbf{x}_{*})$
+$\max_{\begin{subarray}{c}d,f\\ \mathbf{x}_{*},\mathbf{x}_{0},\cdots,\mathbf{x}_{k}\in\mathbb{R}^{d}\end{subarray}}\quad$ $f(\mathbf{x}_{K})-f(\mathbf{x}_{*})$
 $\operatorname{s.t.}\qquad\quad$ $f\in\mathcal{F}_{0,L},\quad\mathbf{x}_{*}\in\operatorname*{argmin}_{\mathbf{x}}f(\mathbf{x})$
@@
 $\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\nabla f(\mathbf{x}_{i})\quad k=1,\cdots,K,$
-where $\mathcal{F}_{0,L}$ denotes the class of merely convex functions with Lipschitz continuous gradients. The general first-order algorithm (GFO) performs the update
-$\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\nabla f(\mathbf{x}_{i})\quad k=1,\cdots,K,$
-which covers a wide range of first-order methods, including GD, FGM, Polyak’s Heavy Ball Method [55], and Optimized Gradient Method (see Algorithms 4 and 3[22].
-The above problem is an infinite dimensional optimization as the variable $f$ is a function.
-[22] proposed a notion of $\mathcal{F}_{\mu,L}$-interpolation that only considers function values and gradients at discrete points $\{\mathbf{x}_{k}\}$, which are the terms that play a role in the optimization problem.
+where $\mathcal{F}_{0,L}$ denotes the class of merely convex functions with Lipschitz continuous gradients.
+Note that $\mathbf{x}_{k}=\mathbf{x}_{0}-\sum_{i=0}^{k-1}\frac{1}{L}\theta_{k,i}\nabla f(\mathbf{x}_{i}),\ k=1,\cdots,K$, is the general first-order algorithm (GFO) performs which covers a wide range of first-order methods, including GD, FGM, Polyak’s Heavy Ball Method [47], and Optimized Gradient Method [18].
+The above problem is an infinite-dimensional optimization as the variable $f$ is a function.
+Drori and Teboulle [18] proposed a notion of $\mathcal{F}_{\mu,L}$-interpolation that only considers function values and gradients at discrete points $\{\mathbf{x}_{k}\}$, which are the terms that play a role in the optimization problem.
 Definition 3.1 ($\mathcal{F}_{\mu,L}$-interpolation (Definition 2 in [55])).
 Let $\mathcal{I}$ be an index set, and consider the set of triples $\mathcal{S}=\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in\mathcal{I}}$ where $\mathbf{x}_{i},\mathbf{g}_{i}\in\mathbb{R}^{d}$ for all $i\in\mathcal{I}$.
-Set $\mathcal{S}$ is $\mathcal{F}_{\mu,L}$-interpolatable if and only if there exists a function $f\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ such that we have both $\mathbf{g}_{i}=\nabla f(\mathbf{x}_{i})$ and $f(x_{i})=f_{i}$ for all $i\in\mathcal{I}$.
-Theorem 2.2 ($\mathcal{F}_{\mu,L}$-interpretability (Theorem 4 in [63]) ).
-Set $\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in I}$ is $\mathcal{F}_{\mu,L}$-interpolable ($0\leq\mu\leq L\leq\infty$) if and only if the following set of conditions holds for every pair of indices $i\in I$ and $j\in I$:
+Set $\mathcal{S}$ is $\mathcal{F}_{\mu,L}$-interpolatable if and only if there exists a function $f\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ such that we have both $\mathbf{g}_{i}=\nabla f(\mathbf{x}_{i})$ and $f(\mathbf{x}_{i})=f_{i}$ for all $i\in\mathcal{I}$.
+Theorem 3.2 ($\mathcal{F}_{\mu,L}$-interpolable (Theorem 4 in [55]) ).
+Set $\{(\mathbf{x}_{i},\mathbf{g}_{i},f_{i})\}_{i\in\mathcal{I}}$ is $\mathcal{F}_{\mu,L}$-interpolable ($0\leq\mu\leq L\leq\infty$) if and only if the following set of conditions holds for every pair of indices $i\in\mathcal{I}$ and $j\in\mathcal{I}$:
 $f_{i}-f_{j}-\mathbf{g}_{j}^{\top}(\mathbf{x}_{i}-\mathbf{x}_{j})\geq\frac{1}{2(1-\mu/L)}\left(\frac{1}{L}\|\mathbf{g}_{i}-\mathbf{g}_{j}\|^{2}+\mu\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}-\frac{2\mu}{L}(\mathbf{g}_{j}-\mathbf{g}_{i})^{\top}(\mathbf{x}_{j}-\mathbf{x}_{i})\right).$ (9)
@@
 The PEP model can also be used to find algorithms that minimize the worst-case function.
-In Drori and Teboulle [22], numerical results indicate a better algorithm than the well-known FGM.
-[39] follows this idea and provides an explicit form of the generated algorithm called the Optimized Gradient Method (OGM), which was proven to match the lower bound with a smaller $\mathcal{O}(1)$ constant [20].
-[21] further reveals the equivalence between OGM and the conjugate gradient method–see also [40, 42, 43, 54].
+In Drori and Teboulle [18], numerical results indicate existence of a better algorithm than the well-known FGM.
+Kim and Fessler [32] follow this idea and provide an explicit form of the generated algorithm called the Optimized Gradient Method (OGM), which is proven to match the lower bound with a smaller $\mathcal{O}(1)$ constant [16].
+Drori and Taylor [17] further reveal the equivalence between OGM and the conjugate gradient method–see also [33, 35, 36, 46].
 PEP technique has been successfully applied to various optimization methods, including gradient descent [26], algorithms with line search [9, 17], proximal gradient methods [34], proximal point algorithms [31], algorithms with inexact oracles [10, 3], and many others [2, 50, 19].
 Besides the deterministic smooth regimes, PEP has also been incorporated for stochastic problems [54] and for problems satisfying relative smoothness [57].
-When optimizing only over the function class given an algorithm, PEP can be equivalently written as a Semidefinite Programming (SDP) and solved efficiently.
-However, when optimizing over both algorithm and function, the problem is nonconvex, and finding its global solution is NP-hard.
+Given an algorithm, if $d$ is large enough, PEP can be equivalently written as a Semidefinite Programming (SDP) and solved efficiently.
+However, when optimizing over both algorithm and function, the underlying minmax problem (see (13)) is nonconvex, and finding its global solution is NP-hard.
 Hence, to find an optimal algorithm a relaxed PEP formulation is generally solved which may not provide an optimal algorithm.
 Das Gupta et al.
-[10] proposes a method to solve the original nonconvex PEP and provides a numerical guide to design algorithms.
+[7] propose a method to solve the original nonconvex PEP and provides a numerical guide to design algorithms.
+Remark 3.3.
 A closely related topic is the analysis based on Integral Quadratic Constraints (IQC), originally studied in control theory [40].
@@
 and the Gram matrix $\mathbf{G}\triangleq\mathbf{X}^{\top}\mathbf{X}\in\mathbb{S}^{2K+2}$ which is a symmetric positive semidefinite (PSD) matrix.
-Also, we define standard basis vectors $\vec{\mathbf{f}}_{k}\in\mathbb{R}^{K+1}$ and $\vec{\mathbf{x}}_{k},\vec{\mathbf{g}}_{k},\vec{\mathbf{e}}_{k}\in\mathbb{R}^{2K+2}$ to select corresponding columns in $\mathbf{F}$ or $\mathbf{X}$ to recover the needed element, which satisfy
+Also, we define sparse vectors $\vec{\mathbf{f}}_{k}\in\mathbb{R}^{K+1}$ and $\vec{\mathbf{x}}_{k},\vec{\mathbf{g}}_{k},\vec{\mathbf{e}}_{k}\in\mathbb{R}^{2K+2}$ to select corresponding columns in $\mathbf{F}$ or $\mathbf{X}$ to recover the needed element, which satisfy
 $f_{k}-f_{*}=\mathbf{F}\vec{\mathbf{f}}_{k},\quad\vec{\mathbf{g}}_{*}=\mathbf{0},\quad\mathbf{g}_{k}=\mathbf{X}\vec{\mathbf{g}}_{k},\quad\mathbf{e}_{k}=\mathbf{X}\vec{\mathbf{e}}_{k},$
@@
 Note.
-We note that the subscript $k$ on the standard basis vectors does not mean that the $k$-th element of that vector is equal to 1.
-For instance, based on the definitions of $\mathbf{F}$ and $\mathbf{X}$ in (7), $\vec{\mathbf{f}}_{k}$ is a standard basis vector with $k+1$-th element equal to 1, $\vec{\mathbf{e}}_{k}$ is a standard basis vector with $k+2$-th element equal to 1, $\vec{\mathbf{g}}_{k}$ is a standard basis vector with $K+k+2$-th element equal to 1, and $\vec{\mathbf{x}}_{0}$ is a standard basis vector with the first element equal to 1.
-With a slight abuse of the notation, $\vec{\mathbf{x}}_{k}$ (which is not a standard basis vector) is the vector used to represent $\mathbf{x}_{k}$ by a linear combination of the columns of $\mathbf{X}$ and it is a function of $\theta_{k,i}$.
+We note that the subscript $k$ on these 0-1 vectors does not mean their $k$-th elements are equal to 1.
+For instance, based on the definitions of $\mathbf{F}$ and $\mathbf{X}$ in (11), $\vec{\mathbf{f}}_{k}$, $\vec{\mathbf{e}}_{k}$, and $\vec{\mathbf{g}}_{k}$ are standard basis vectors with $k+1$-th, $k+2$-th, $K+k+2$-th elements equal to 1, respectively, and $\vec{\mathbf{x}}_{0}$ is a standard basis vector with its first element equal to 1.
+With a slight abuse of the notation, $\vec{\mathbf{x}}_{k}$ (which is not a basis vector) is the vector used to represent $\mathbf{x}_{k}$ by a linear combination of the columns of $\mathbf{X}$, and it is a function of $\theta_{k,i}$.
 Without loss of generality, we assume $f_{*}=0$ and $\mathbf{x}_{*}=0$, hence $\vec{\mathbf{f}}_{*}=\mathbf{0},\vec{\mathbf{x}}_{*}=\mathbf{0}$. With this new notation, the PEP problem is written as
@@
 The optimal solution of (SDP-PEP) will provide an exact worst-case performance of a given algorithm for convex problems.
-It is also possible to find the optimal algorithm with parameter ${\boldsymbol{\theta}}$ if we consider the minimax problem as
-$\min_{\theta_{k,i}}\max_{\mathbf{G},\mathbf{F}\in\mathcal{S}(\text{SDP-PEP})}\mathbf{F}\vec{\mathbf{f}}_{K},$ (9)
+It is possible to find the optimal algorithm with parameter $\boldsymbol{\theta}$} by solving the minimax problem
 […]
+where $\mathbf{G},\mathbf{F}\in\mathcal{S}(\text{SDP-PEP})$ refers to $\mathbf{G}$ and $\mathbf{F}$ satisfy the constraints of (SDP-PEP).
+Since $\mathbf{A}^{i,j}$ is indeed a function of $\{\theta_{k,i}\}$, this minimax problem has bilinear terms in the constraints of the inner problem, and solving it for the global solution is NP-hard.
+OGM’s [18] idea is to select a subset of the constraints and upper bound the inner maximization by its dual minimization problem.
+By wisely selecting the constraints, the hard minimax problem can be relaxed as a bilinear minimization problem which can be further relaxed as a linear SDP.
+Solving this final problem numerically inspires a theoretical solution and results in the first version of OGM when gradients are exact.
 Recall that for FGM, there are two sequences, the primary sequence $\{\mathbf{y}_{k}\}$ and the secondary one $\{\mathbf{x}_{k}\}$.
 The OGM is optimized for minimizing $f(\mathbf{x}_{k})-f_{*}$ while the commonly used measure is $f(\mathbf{y}_{K})-f_{*}$.
-Due to this fact, [40] propose a new measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ since for smooth function and with OGM update rule, $f(\mathbf{y}_{K+1})=f(\mathbf{x}_{K}-\frac{1}{L}\nabla f(\mathbf{x}_{K}))\leq f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$.
-The solution of this new measure results in OGM (Algorithm 4).
-These two algorithms only differ in the last iteration’s $\alpha_{K}$ update.
-For simplicity, we only focus on this measure and its corresponding relaxed PEP.
-Algorithm 3 OGM with last adjustment
-0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
-1:  for $k=0,\cdots,K-1$ do
-2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
-3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
-4:     $\alpha_{k+1}=\begin{cases}\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2},k\leq K-2\\ \frac{1+\sqrt{1+8\alpha_{k}^{2}}}{2},k=K-1\end{cases}$
-5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
-6:  end for
-Algorithm 4 OGM
-0:  $\mathbf{z}_{0}=\mathbf{x}_{0}$, $\alpha_{0}=1$.
-1:  for $k=0,\cdots,K-1$ do
-2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\nabla f(\mathbf{x}_{k})$
-3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\nabla f(\mathbf{x}_{k})$
-4:     $\alpha_{k+1}=\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2}$
-5:     $\mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1}$
-6:  end for
-The relaxed optimization problem is
+Due to this fact, Kim and Fessler [33] propose a new measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ since for smooth function and with OGM update rule, $f(\mathbf{y}_{K+1})=f(\mathbf{x}_{K}-\frac{1}{L}\nabla f(\mathbf{x}_{K}))\leq f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$.
+Based on the constraint selection and objective modification discussed above, the relaxed optimization problem is
 […]
@@
 $-\tau\left(\operatorname{Tr}(\mathbf{G}\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top})-R^{2}\right)-\sum_{i=0}^{K-1}u_{i}\left(\operatorname{Tr}(\mathbf{G}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top})-b_{i}^{2}\right)$
-$=$ $\tau R^{2}+\sum_{k=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
+$=$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
 $-\operatorname{Tr}\left(\mathbf{G}\left(\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K-1}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\right)\right).$
@@
 $\operatorname{s.t.}\quad$ $\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}_{i}}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}v_{*,i}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})=0$
-$\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\succeq 0.$
-Here we omit the process of relaxing the minimax problem and use the solution in [42] directly to explain the property of this feasible solution with the inexact oracle.
-Define $A_{k}=\sum_{i=0}^{k}\alpha_{i}$ and $A_{0}=1,A_{-1}=0$, and assume $\alpha_{k}^{2}\leq A_{k}$, the feasible solution of the minimax problem is
+[…]
+Kim and Fessler [32] analytically solved (D) nested inside the minimization over the algorithm parameters $\boldsymbol{\theta}$}, in the absence of gradient inexactness, i.e., when $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}=0$ and $\sum_{i=0}^{K}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}=0$, and derived OGM with the following solution,
 […]
@@
 […]
+with $A_{k}=\sum_{i=0}^{k}\alpha_{i}$, $A_{0}=1,A_{-1}=0$.
+Further, they showed that this solution is also a feasible solution to GOGM [35] with $\alpha_{k}^{2}\leq A_{k}$.
+We will follow this process, use the feasible solution above, and derive the value of $\mathbf{u}$ for GOGM with the inexact gradients.
 Remark 3.4.
-The above feasible solution works for both OGM and Generalized OGM (discussed in the next section), which means for any stepsize $\{\alpha_{k}\}$ and $\{A_{k}\}$ that satisfy $\alpha_{k}^{2}\leq A_{k}$, (10) and (11) satisfies the constraints of (D).
-It is easy to verify that the equality constraint holds with this solution.
-We still need to find the feasible solution for $\mathbf{u}$ to satisfy the positive semidefinite (PSD) constraint.
-After substitution by (10) and (11), the left-hand-side matrix of the PSD constraint is
-$\mathbf{M}=\begin{pmatrix}[l]\frac{L}{4A_{K}}&[\mathbf{0}]_{1\times K}&[\mathbf{p}]_{1\times(K+1)}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times(K+1)}\\ [\mathbf{p}^{\top}]_{(K+1)\times 1}&[\mathbf{B}^{\top}]_{(K+1)\times K}&[\mathbf{C}]_{(K+1)\times(K+1)}\end{pmatrix},$
+We attempted to derive an optimized inexact first-order method by extending the procedure used for OGM [32].
+However, this approach revealed a significant practical limitation.
+The core issue is that the derived optimal step sizes for the inexact method lack a recursive structure.
+As shown in Appendix B, in the exact gradient case, the step size for a given iteration is independent of the total number of iterations $K$.
+This allows for a practical, recursive implementation where past iterates remain fixed.
+In the inexact setting, however, this property does not hold.
+The optimal step sizes for all previous iterations depend on the pre-specified total number of iterations $K$.
+This means that if one were to extend a run by increasing $K$, all previously computed steps would no longer be optimal and would need to be recomputed.
+This phenomenon makes such an algorithm computationally impractical for implementation.
+It is easy to verify that the equality constraint in (D) holds with this solution.
+We still need to find a feasible solution for $\mathbf{u}$ to satisfy the PSD constraint in (14).
+After substitution by (15) and (16), the left-hand-side matrix of the PSD constraint (denoted as $\mathbf{M}$) in (14) is
+$\mathbf{M}=\begin{pmatrix}[l]\tau&[\mathbf{0}]_{1\times K}&[\mathbf{p}]_{1\times(K+1)}\\ [\mathbf{0}]_{K\times 1}&[\mathbf{U}]_{K\times K}&[\mathbf{B}]_{K\times(K+1)}\\ [\mathbf{p}^{\top}]_{(K+1)\times 1}&[\mathbf{B}^{\top}]_{(K+1)\times K}&[\mathbf{C}]_{(K+1)\times(K+1)}\end{pmatrix}$,
 where $\mathbf{U}=\operatorname{diag}(u_{0}\cdots u_{K-1})$, $\mathbf{p}_{i}=-\frac{\alpha_{i-1}}{2A_{K}}$,
@@
 $\mathbf{R}=\left(\begin{array}[]{c|ccc|cccc}0&0&0&0&0&0&0&0\\[2.0pt] \hline\cr 0&{u}_{0}&0&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{1}&0&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&0&{u}_{2}&0&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr 0&0&0&0&0&0&0&0\\ 0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&0&\frac{A_{1}-\alpha_{1}^{2}}{LA_{3}}&0&0\\ 0&\frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&0&0&\frac{A_{2}-\alpha_{2}^{2}}{LA_{3}}&0\\ 0&\frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&0&0&0&\frac{A_{3}-\alpha_{3}^{2}}{LA_{3}}\end{array}\right).$
-Recalling that $\mathbf{R}^{\backslash[1]}$ denotes the matrix resulting from dropping the first column and row of $\mathbf{R}$, this submatrix is exactly the Shur complement of the block $[\tau]$ of the matrix $\mathbf{M}$.
-Since $\tau>0$, we know that $\mathbf{M}\succeq 0$ if and only if $\mathbf{R}^{\backslash[1]}\succeq 0$.
-In the following lemma, we will show that $\mathbf{u}=\infty$ is the unique feasible solution when $A_{i}-\alpha_{i}^{2}=0$.
+Recalling that $\mathbf{R}^{\backslash[1]}$ denotes the matrix resulting from dropping the first column and row of $\mathbf{R}$, this submatrix is exactly the Shur complement of the block $\tau]$ of the matrix $\mathbf{M}$.
+Since $\tau>0$ (as $\tau=\frac{L}{4A_{K}}$), we know that $\mathbf{M}\succeq 0$ if and only if $\mathbf{R}^{\backslash[1]}\succeq 0$.
+In the following lemma, we will show that $\mathbf{u}=\infty$ is the unique feasible solution when $A_{i}=\alpha_{i}^{2},\forall i=1,\cdots,K$.
 Lemma 3.5.
-With the given solution in (10) and (11), and $A_{i}-\alpha_{i}^{2}=0$, $\mathbf{u}=\infty$ is the unique solution that makes $\mathbf{R}^{\backslash[1]}\succeq 0$.
-When $A_{i}-\alpha_{i}^{2}>0$, $\mathbf{u}=\infty$ is still feasible but not unique.
+With the given solution in (15) and (16), if $A_{i}-\alpha_{i}^{2}=0,\forall i=1,\cdots,K$, then $\mathbf{u}=\infty$ is the unique solution that makes $\mathbf{R}^{\backslash[1]}\succeq 0$.
+When $A_{i}-\alpha_{i}^{2}>0$, $\mathbf{u}=\infty$ is still a feasible solution but not the unique one.
 Proof.
@@
 holds. To show that, there are four different scenarios:
-  1. 1.
-    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=0$,
-  2. 2.
-    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$,
-  3. 3.
-    $\mathbf{R}^{\backslash[1]}_{i,i}=u_{k},\mathbf{R}^{\backslash[1]}_{j,j}=0$,
-  4. 4.
-    $\mathbf{R}^{\backslash[1]}_{i,i}=0,\mathbf{R}^{\backslash[1]}_{j,j}=u_{l}$.
+[…]
 For the first and second scenarios, it is obvious that $\mathbf{R}^{\backslash[1]}_{i,j}=0$, with diagonal elements being non-negative, the condition holds.
@@
 $\mathbf{v}^{\top}\mathbf{R}^{\backslash[1]}\mathbf{v}=c^{2}u_{l}+2c\mathbf{R}^{\backslash[1]}_{i,j}\geq 0$ holds only with $u_{l}=\infty$.
-We get the final result by combining all these scenarios.
 ∎
@@
 4 Generalized OGM with inexact gradient oracle
-The properties of OGM and its generalization have been thoroughly studied.
-However, the convergence analysis of OGM with inexact gradient oracle has not been performed yet.
-Furthermore, the existing analysis of FGM with inexact oracles with the absolute error assumption depends on the trajectory of sequences.
-This section aims at closing these gaps.
-It is possible to optimize the algorithm with the same procedure as OGM, but the solution depends on iteration number $K$, initial condition $R$, and inexactness levels $\{b_{i}\}_{i=1}^{K}$.
-It is difficult to find a recursive structure of the algorithm that captures the numerical solution (see the experiment results in Appendix B).
-Hence, we focus on Generalized OGM (GOGM) with inexact oracle and try to find its convergence rate with $b_{i}>0$.
-Furthermore, the convergence bound of Generalized FGM (GFGM) is also analyzed in Section 3.3.
-The GOGM is a generalization of OGM (Algorithm 4) as it only requires the stepsize to satisfy $\alpha_{k}^{2}\leq A_{k}$.
-The inexact GOGM (iGOGM) is provided in Algorithm 5, in which the gradients in steps 2 and 3 are replaced by their inexact alternatives.
-Algorithm 5 Inexact Generalized Optimized Gradient Method (iGOGM)
-0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
-1:  for $k=0,\cdots,K-1$ do
-2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
-3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{2}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
-4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
-5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
-6:  end for
-The update rule in step 4 guarantees $\alpha_{k}^{2}\leq A_{k}$ when $\lambda_{k}\in[0,1]$ as
-$\alpha_{k+1}^{2}=(A_{k+1}-A_{k})^{2}=\lambda_{k+1}A_{k+1}\leq A_{k+1}.$
-Furthermore, when $\lambda=1$, the $\mathbf{x}_{k+1}$ update (step 5) reduce to
-$\mathbf{x}_{k+1}=\left(1-\frac{1}{\alpha_{k+1}}\right)\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},$
-the common update step in both FGM and OGM.
-Remark 3.1.
-The optimality measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ used in the analysis of exact GOGM is also a meaningful one for the inexact problem.
-For iGOGM, we have
-$f(\mathbf{y}_{K+1})=$ $f\left(\mathbf{x}_{K}-\frac{1}{L}(\nabla f(\mathbf{x}_{k})+\mathbf{e}_{K})\right)$
-$\leq$ $f(\mathbf{x}_{K})-\frac{1}{L}\left\langle\nabla f(\mathbf{x}_{K}),\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\rangle+\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})+\mathbf{e}_{K}\right\|^{2}$
-$=$ $f(\mathbf{x}_{K})-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}+\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}.$
-Comparing to the measure $f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$, the extra term $\frac{1}{2L}\left\|\mathbf{e}_{K}\right\|^{2}$ can be discarded since the constraint $\left\|\mathbf{e}_{K}\right\|^{2}\leq b_{K}^{2}$ is the only one that involves $\mathbf{e}_{K}$ and the PEP solution sets $\left\|\mathbf{e}_{K}\right\|^{2}=b_{K}^{2}$.
-Since the algorithm only generates $\{\mathbf{x}_{k}\}$ and $\{\mathbf{y}_{k}\}$ up to $k=K$, $\mathbf{y}_{K+1}$ is a dummy variable used for the theoretical guarantee, and setting $b_{K}=0$ does not affect the complexity of the algorithm.
-3.1 PEP approach for the convergence analysis of iGOGM
-The goal of this part is to find an analytical feasible solution of (D) with iGOGM stepsize.
-The feasible solution provides an upper bound of the worst-case convergence rate.
+The goal of this part is to find an analytical feasible solution of (D) with the iGOGM stepsize.
+From the weak duality theorem, the feasible solution provides an upper bound on the worst-case convergence rate.
 To do so, first notice that the values of $\{v_{*,i}\}$ are uniquely determined by $\{v_{i,i+1}\}$ through the equality constraint in (D) and, hence, can be removed from the decision variables.
@@
 Recall the objective of the problem (D), $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, contains two components: the first one is the convergence rate and the second one is the accumulated error.
-Given the convergence rate of the exact GOGM, i.e., $\frac{LR^{2}}{4A_{K}}$, we adjust the objective of (D) as
-$\frac{L}{4A_{K}}R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2},$
-which is the summation of the convergence rate of exact GOGM and the accumulated error $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$.
+Given the convergence rate of the exact GOGM, i.e., $\frac{LR^{2}}{4A_{K}}$, we adjust the objective of (D) as $\frac{L}{4A_{K}}R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$, which is the summation of the convergence rate of exact GOGM and the accumulated error $\sum_{i=0}^{K-1}u_{i}b_{i}^{2}$.
 With these replacements, the decision variables we need to consider are $\{v_{i,i+1}\}$ and $\{u_{i}\}$.
-To simplify the notation, in the following text, we will use $\mathbf{v}$ and $\mathbf{u}$ to represent the sequences $\{v_{i,i+1}\}$ and $\{u_{i}\}$, respectively.
-Solving (D) with this new objective, we observe from the numerical solution that $\mathbf{v}$ is almost equal to its analytical counterpart for problems with exact oracles given in (10).
+To simplify the notation, below, we will use $\mathbf{v}$ and $\mathbf{u}$ to represent the sequences $\{v_{i,i+1}\}$ and $\{u_{i}\}$, respectively.
+Solving (D) with the new objective, we observe from the numerical solution that $\mathbf{v}$ is almost equal to its analytical counterpart for problems with exact oracles given in (15).
 In this experiment, we set $K=\{10,20,\cdots,100\}$, $R=\{1,10,1000\}$, and $L=\{1,10,1000\}$ which result in 90 different scenarios in total.
 For each scenario, we generate 10 instances with random stepsize parameter $\{\lambda_{k}\}$ and inexactness levels $\{b_{i}\}$.
-In Figure 1, three examples of the relative difference between the numerical solution of inexact GOGM $v^{num.}_{k}$ and the analytical solution of exact GOGM $v^{anly.}_{k}$ are illustrated.
+In Figure 5, three examples of the relative difference between the numerical solution of iGOGM $v^{num.}_{k}$ and the analytical solution of exact GOGM $v^{anly.}_{k}$ are illustrated.
 The solid line is an average of 90 instances.
@@
 Figure 6: Difference between numerical and analytical solution for $\mathbf{v}$
-As we claimed in Lemma 2.4, OGM with (10) has a unique feasible solution with $\mathbf{u}=\infty$.
-To make $\mathbf{u}$ to be bounded, we have to guarantee that the stepsize condition holds with strict inequality, i.e.
-$A_{k}-\alpha_{k}^{2}>0$.
+As we claimed in Lemma 3.5, OGM ($A_{k}=\alpha_{k}^{2}$) with (15) has a unique feasible solution with $\mathbf{u}=\infty$.
+To make $\mathbf{u}$ to be bounded, we have to guarantee that the stepsize condition holds with strict inequality, i.e., $A_{k}-\alpha_{k}^{2}>0$.
 Recall the matrix $\mathbf{R}^{\backslash[1]}$
@@
 where the elements of the first column of $\mathbf{B}$ and $\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}$ are zero.
-So we can discard this zero row and column and write the $\mathbf{R}^{\backslash[1,K+2]}$ as
+So we can discard this zero row and column and write the resulting matrix $\mathbf{R}^{\backslash[1,K+2]}$ as
 $\mathbf{R}^{\backslash[1,K+2]}=\begin{pmatrix}[l][\mathbf{U}]_{K\times K}&[\mathbf{B}^{\backslash[1]}]_{K\times K}\\ [\mathbf{B}^{\backslash[1]\top}]_{K\times K}&[\mathbf{D}]_{K\times K}\end{pmatrix}$,
-with $\mathbf{D}\triangleq[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]^{\backslash[1]}$.
-Example of $\mathbf{R}^{\backslash[1,K+2]}$ with $K=3$ is
-$\mathbf{R}^{\backslash[1,K+2]}=\left(\begin{array}[]{ccc|cccc}{u}_{0}&0&0&\frac{1+2{\alpha}_{1}}{2LA_{3}}&\frac{{\alpha}_{2}}{LA_{3}}&\frac{{\alpha}_{3}}{LA_{3}}\\ 0&{u}_{1}&0&0&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}\\ 0&0&{u}_{2}&0&0&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}\\[2.0pt] \hline\cr\frac{1+2{\alpha}_{1}}{2LA_{3}}&0&0&\frac{A_{1}-\alpha_{1}^{2}}{LA_{3}}&0&0\\ \frac{{\alpha}_{2}}{LA_{3}}&\frac{2{\alpha}_{1}{\alpha}_{2}+A_{1}}{2LA_{3}}&0&0&\frac{A_{2}-\alpha_{2}^{2}}{LA_{3}}&0\\ \frac{{\alpha}_{3}}{LA_{3}}&\frac{{\alpha}_{1}{\alpha}_{3}}{LA_{3}}&\frac{2{\alpha}_{2}{\alpha}_{3}+A_{2}}{2LA_{3}}&0&0&\frac{A_{3}-\alpha_{3}^{2}}{LA_{3}}\end{array}\right).$
+with $\mathbf{D}\triangleq[\mathbf{C}-\frac{1}{\tau}\mathbf{p}^{\top}\mathbf{p}]^{\backslash[1]}$. The example of $\mathbf{R}^{\backslash[1,K+2]}$ with $K=3$ is
+[…]
 We need to find $\mathbf{u}$ such that $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$.
-Here we notice that $\mathbf{D}=\operatorname{diag}(\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})\succ 0$, hence, it is invertible with all elements strictly bigger than 0.
-We take the Shur complement of the block $\mathbf{D}$ and define the result matrix as $\mathbf{S}$ with
+Here we notice that $\mathbf{D}=\operatorname{diag}(\{(A_{i-1}-\alpha_{i-1}^{2})/(LA_{K})\})$, hence, it is invertible as all diagonal elements are strictly bigger than 0.
+We take the Shur complement of the block $\mathbf{D}$ and define the result matrix as $\mathbf{S}$ where
 $\mathbf{S}_{i,j}=\begin{cases}u_{i-1}-\frac{(A_{i-1}+2\alpha_{i-1}\alpha_{i})^{2}}{4LA_{K}(A_{i}-\alpha_{i}^{2})}-\sum_{k=i}^{K-1}\frac{\alpha_{i-1}^{2}\alpha_{k+1}^{2}}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\quad i=j\\ -\frac{\alpha_{i}\alpha_{j-1}(A_{i-1}+2\alpha_{i-1}\alpha_{i})}{2LA_{K}(A_{i}-\alpha_{i}^{2})}-\sum_{k=i}^{K-1}\frac{\alpha_{i-1}\alpha_{j-1}\alpha_{k+1}^{2}}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\quad i>j\\ \mathbf{S}_{ji}\quad j>i\end{cases}.$
-Example with $K=3$ is
-$\mathbf{S}=\begin{pmatrix}{u}_{0}-\frac{\left(1+2\alpha_{1}\right)^{2}}{4LA_{3}\left({A}_{1}-{\alpha}_{1}^{2}\right)}-\frac{{\alpha}_{2}^{2}}{LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left({A}_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right){\alpha}_{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{1}-\frac{\left({A}_{1}+2{\alpha}_{1}{\alpha}_{2}\right)^{2}}{4LA_{3}\left({A}_{2}-{\alpha}_{2}^{2}\right)}-\frac{{\alpha}_{1}^{2}{\alpha}_{3}^{2}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\\ -\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{3}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right){\alpha}_{1}{\alpha}_{3}}{LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}&{u}_{2}-\frac{\left(A_{2}+2{\alpha}_{2}{\alpha}_{3}\right)^{2}}{4LA_{3}\left({A}_{3}-{\alpha}_{3}^{2}\right)}\end{pmatrix}.$
-Then, $\mathbf{S}\succeq 0$ iff $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$.
-A nice property of $\mathbf{S}$ is that $\{u_{i}\}$ only shows in the diagonal elements and all the off-diagonal elements are negative constants.
-It is known that a symmetric diagonally dominated matrix with non-negative diagonal entries is PSD, with the definition of the diagonally dominated matrix as
+The example with $K=3$ is
+[…]
+Then, we have $\mathbf{R}^{\backslash[1,K+2]}\succeq 0$ iff $\mathbf{S}\succeq 0$.
+A nice property of $\mathbf{S}$ is that $\{u_{i}\}$ only shows in the diagonal elements, and all the off-diagonal elements are negative.
+To ensure $\mathbf{S}\succeq 0$, we use the property of the diagonally dominated matrix, which is that a symmetric diagonally dominated matrix with non-negative diagonal entries is PSD, with the definition of the diagonally dominated matrix as
 $\left|\mathbf{S}_{i,i}\right|\geq\sum_{i\neq j}\left|\mathbf{S}_{i,j}\right|.$
-Based on the special structure of $\mathbf{S}$, we can simplify the above inequality as
-$\mathbf{S}_{i,i}\geq-\sum_{i\neq j}\mathbf{S}_{i,j}.$
+Based on the special structure of $\mathbf{S}$, we can simplify the above inequality as $\mathbf{S}_{i,i}\geq-\sum_{i\neq j}\mathbf{S}_{i,j}$.
 Inserting the definition of $\mathbf{S}_{i,j}$, we can calculate the theoretical feasible solution for $\mathbf{u}$ as
@@
 The solution above also shows that when $A_{k}-\alpha_{k}^{2}=0$, $u=\infty$, which recovers the unique feasible solution we proved in Lemma 3.5.
-Surprisingly, this solution matches the numerical solution when $b_{i}\equiv b$.
-In Figure 3, we follow the same settings as for Figure 2 with $\mathbf{v}$ equals to the values defined in Equation 10.
+Surprisingly, this solution is close to the numerical solution when $b_{i}\equiv b$.
+In Figure 7, we follow the same settings as for Figure 6 with $\mathbf{v}$ equals to the values defined in (15).
 The error between the optimal solution and the analytical $\hat{\mathbf{u}}$ given in (19) is under $10^{-3}$, validating our theoretical result.
 Figure 7: Difference of the numerical and analytical solution $\hat{\mathbf{u}}$ given in (19) for different settings with fixed $b_{i}$ along iterations.
-Finally, in Lemma 3.2, we provide an analytical solution for $\mathbf{u}$ that together with the analytical solutions for $\mathbf{v}$ and $\mathbf{v}_{*}$ (from the problem with exact oracle) provide a feasible solution to (D).
-From the above discussion, we can summarize the results in the following lemma. The lemma provides a feasible solution to problem (D) with iGOGM algorithm.
+We now summarize the above results in the following lemma.
+Specifically, Lemma 4.1 establishes a feasible solution to problem (D) for the iGOGM algorithm.
+This is achieved by providing an analytical solution for $\mathbf{u}$ which, when combined with the analytical solutions for $\mathbf{v}$ and $\mathbf{v}_{*}$ from the problem with the exact oracle, satisfies the problem’s constraints.
 Lemma 4.1 (Feasible solution to (D) with iGOGM).
-For Algorithm 5, the following solution is a feasible solution to the problem (D). $\forall i=0,\cdots,K-1$
+Under Assumption 2.1, for the iGOGM (Algorithm 2), the following solution is a feasible solution to the problem (D).
 $\tau$ $=\frac{L}{4A_{K}},\quad v_{i,i+1}=\frac{A_{i}}{A_{K}},\quad v_{*,0}=v_{0,1},\quad v_{*,i}=v_{i,i+1}-v_{i-1,i},\quad v_{*,K}=1-v_{K-1,K},$
-$u_{i}$ $=\frac{A_{i}(1+2\alpha_{i+1})(A_{i}+2\alpha_{i}\alpha_{i+1})}{4LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}+\sum_{k=i+1}^{K-1}\frac{A_{k}(1+2\alpha_{k+1})\alpha_{i}\alpha_{k+1}}{2LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}.$
-3.2 PEP-inspired convergence analysis of iGOGM
-As explained by [30], the feasible solution of PEP provides the proof of the convergence bound for the first-order algorithms.
-In the following theorem and the corresponding proof, we will show a PEP-inspired approach that can be verified without any knowledge of PEP or even SDP.
-Theorem 3.3.
-The sequence generated by Algorithm 5 satisfies
-$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
-with $u_{k}=\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}$.
-Proof.
-Below are the inequalities used in PEP primal constraints and their corresponding dual variables:
+$u_{i}$ $=\frac{A_{i}(1+2\alpha_{i+1})(A_{i}+2\alpha_{i}\alpha_{i+1})}{4LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}+\sum_{k=i+1}^{K-1}\frac{A_{k}(1+2\alpha_{k+1})\alpha_{i}\alpha_{k+1}}{2LA_{K}(A_{k+1}-\alpha_{k+1}^{2})},\forall i=0,\cdots,K-1.$
+4.1 PEP-inspired proof of the convergence analysis of iGOGM
+In this section, we provide the detailed proof of Theorem 2.2.
+As explained in [25], a feasible solution to the PEP’s dual problem (SDP-PEP), or its relaxed version (D), provides a direct proof for the algorithm’s convergence bound.
+We note that even though we use the PEP technique to find an upper bound on the convergence bound, the proof can be understood without any prior knowledge of the PEP or SDP’s duality theory.
+Following this idea, we will establish our proof for iGOGM based on the feasible solution we derive in Lemma 4.1.
+Recall the properties of Lagrangian duality, any such feasible solution of (D) yields the following upper bound on the optimization error
+$f(\mathbf{x}_{K})-f_{*}\leq$ $\mathcal{L}(\mathbf{G},\mathbf{F},\mathbf{v},\mathbf{v}_{*},\mathbf{u},\tau)$
+$=$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}+\mathbf{F}\left(\vec{\mathbf{f}}_{K}+\sum_{i=0}^{K-1}v_{i,i+1}(\vec{\mathbf{f}}_{i}-\vec{\mathbf{f}}_{i+1})+\sum_{i=0}^{K}(\vec{\mathbf{f}}_{*}-\vec{\mathbf{f}}_{i})\right)$
+$-\operatorname{Tr}\left(\mathbf{G}\left(\sum_{i=0}^{K-1}v_{i,i+1}\mathbf{A}^{i,i+1}+\sum_{i=0}^{K}v_{*,i}\mathbf{A}^{*,i}+\tau\vec{\mathbf{x}}_{0}\vec{\mathbf{x}}_{0}^{\top}+\sum_{i=0}^{K-1}u_{i}\vec{\mathbf{e}}_{i}\vec{\mathbf{e}}_{i}^{\top}+\frac{1}{2L}\vec{\mathbf{g}}_{K}\vec{\mathbf{g}}_{K}^{\top}\right)\right)$
+$\leq$ $\tau R^{2}+\sum_{i=0}^{K-1}u_{i}b_{i}^{2}.$
+Based on the above inequality, our proof is constructed by forming a linear combination of the inequalities derived from the function’s convexity and $L$-smoothness properties.
+The dual variables $\mathbf{v}$ and $\mathbf{u}$ act as the coefficients for this combination.
+The first step is to multiply the following inequalities by their corresponding dual variables
 […]
 […]
-Multiply inequalities with their corresponding dual variable and sum up, we have the inequality
-$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
-$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
-$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
-$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
-=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mua\mkern-1.5mu)}}{{=}}$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
-$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
-=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mui\mkern-1.5mu)}}{{=}}$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
-=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mub\mkern-1.5mu)}}{{=}}$ $-\frac{4}{L}\left\|\frac{L}{4}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{4}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle,$
-where (a) replaces $\theta_{k,i}$ with (11), and (b) uses $\tau=L/4$ and the fact that $A_{0}=\alpha_{0}=1$.
-If $\mathbf{e}_{i}=\mathbf{0}$, then the above inequality gives the convergence rate of iGOGM with exact oracle.
-$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|}{4A_{K}}$
-Next, we will bound the error terms when $\mathbf{e}_{i}\neq\mathbf{0}$. For any $k=0,\cdots,K-1$,
-$-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
-=$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5muii\mkern-1.5mu)}}{{=}}$ $\frac{-1}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}$
-$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}^{2}\alpha_{i}^{2}\left\|\mathbf{e}_{i}\right\|^{2}+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k}\sum_{j=0}^{i-1}2\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle$
-$+\frac{1}{L(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}A_{k}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle+\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}.$
-Summing over $k$, and changing the summation indices, we have
-$\sum_{k=0}^{K-1}\left(-\frac{A_{k+1}-\alpha_{k+1}^{2}}{L}\left\|\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{2}{L}\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle\right)$
-$\leq$ $\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
-$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
+[…]
+[…]
+Summing over $k$ and subtracting $\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$ on both sides and rearranging terms we get
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle$
+$-\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle-\frac{1}{2L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}$
+$-\frac{1}{2L}\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}.$
+Replacing $\mathbf{x}_{k}$ with the update rule of iGOGM defined by (10) and (16), the above inequality becomes
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle$
+$+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}.$
+To complete the proof, we must find $\tau$ and $\{u_{k}\}$ that ensures the right-hand-side of the above inequality is bounded by $\tau R^{2}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$.
+This is equivalent to finding a feasible solution for the dual problem (D), since the bound is guaranteed if the positive semidefinite constraint (14) holds.
+Our procedure of finding $\mathbf{u}$ is exactly the step of canceling the remaining terms in the right-hand-side.
+We will provide a detailed explanation below.
+First, we do the Shur complement step corresponding to $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$, which is also the step to bound $\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle$.
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\tau(R^{2}-\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2})+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\underbrace{\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}}_{\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}}$
+$+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau A_{K}^{2}}-\frac{A_{k}}{LA_{K}}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\frac{\alpha_{k}\alpha_{i}}{A_{K}^{2}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle$
+$+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
+$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\frac{4A_{K}}{L}\left\|\frac{L}{4A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle,$
+where in the last equality we replace $\tau=\frac{L}{4A_{K}}$ and $A_{0}=\alpha_{0}=1$.
+In the exact case, $b_{k}\equiv 0$ and $\mathbf{e}_{k}\equiv\mathbf{0}$, the inner-product terms in the above inequality are all zero, and all norm terms have negative coefficients.
+The convergence result of the exact GOGM can be recovered as
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|^{2}}{4A_{K}}.$
+In the inexact case, we need to continue to find a value of $\{u_{k}\}$ that cancels the non-zero inner-product terms.
+To achieve this goal, we will follow the step of the Shur complement applied to matrix $\mathbb{R}^{\backslash[1,K+2]}$ (17), i.e., $\mathbf{R}^{\backslash[1,K+2]}/\mathbf{D}=\mathbf{U}-\mathbf{B}^{\backslash[1]}\mathbf{D}^{-1}\mathbf{B}^{\backslash[1]^{\top}}$:
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=1}^{K}\frac{A_{k}-\alpha_{k}^{2}}{LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$-\frac{2}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
+$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}^{2}\alpha_{i}^{2}\left\|\mathbf{e}_{i}\right\|^{2}$
+$-\underbrace{\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|(A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+\frac{1}{2}A_{k}\mathbf{e}_{k}\right\|^{2}}_{\mathbf{B}^{\backslash[1]}\mathbf{D}^{-1}\mathbf{B}^{\backslash[1]^{\top}}}$
+$+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k}\sum_{j=0}^{i-1}2\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle$
+$+\sum_{k=0}^{K-1}\frac{1}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}A_{k}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle+\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
+$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
 For notation simplicity, we use $P_{k,i}$ to replace the coefficients of $\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$ as
-$P_{k,k}\triangleq$ $\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{L(A_{j+1}-\alpha_{j+1}^{2})},$
-$P_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(A_{j+1}-\alpha_{j+1}^{2})}.$
-Hence, the above equation can be rewritten as
-$\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}P_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
-$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}(\left\|\mathbf{e}_{k}\right\|^{2}+\left\|\mathbf{e}_{i}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
-$=$ $-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=k+1}^{K-1}\frac{1}{2}P_{i,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
-$\stackrel{{\scriptstyle(iii)}}{{\leq}}$ $\sum_{k=0}^{K-1}\left(P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
+$P_{k,k}\triangleq$ $\frac{(2\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})},$
+$P_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(A_{k}+2\alpha_{k+1}\alpha_{k})}{LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{2\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(A_{j+1}-\alpha_{j+1}^{2})}.$
+Hence, the above inequality can be rewritten as
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}P_{k,i}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle$
+$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}(\left\|\mathbf{e}_{k}\right\|^{2}+\left\|\mathbf{e}_{i}\right\|^{2})+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
+$=$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}(b_{k}^{2}-\left\|\mathbf{e}_{k}\right\|^{2})-\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}-\mathbf{e}_{i}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{1}{2}P_{k,i}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{i=k+1}^{K-1}\frac{1}{2}P_{i,k}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}P_{k,k}\left\|\mathbf{e}_{k}\right\|^{2}$
+$\leq$ $\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}+\sum_{k=0}^{K-1}\left(P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}-u_{k}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
 With simple calculations, we can show
 $P_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}P_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}P_{i,k}$
-$=$ $\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4L(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2L(A_{i+1}-\alpha_{i+1}^{2})}.$
-Inserting these results into the original inequality, we get the final bound. ∎
-The key steps of the proof are the equalities $(i)$, $(ii)$, and inequality $(iii)$.
-The equality $(i)$ can be seen as the step of $\mathbf{M}=\frac{1}{\tau}\mathbf{w}_{\tau}\mathbf{w}_{\tau}^{\top}+\mathbf{R}$ if one uses the Gram matrix $\mathbf{G}$ and standard basis vectors to rewrite this equality.
-The equality $(ii)$ can be seen as the reformulation of the Shur complement with the block $[\frac{A_{k+1}-\alpha_{k+1}^{2}}{LA_{K}}]$ of $\mathbf{R}^{\backslash[1,K+2]}$.
-The inequality (iii) is indeed the step of finding $\{u_{k}\}$ such that the matrix $\mathbf{S}$ is diagonally dominant.
-This observation shows that the steps we use to find and verify the feasible solution ($\mathbf{M}\succeq 0$) can be used to generate the theoretical proof.
-When $\left\|\mathbf{e}_{k}\right\|\equiv 0$, the error term will be zero and the upper bound recovers the result of the exact GOGM.
-When $\left\|\mathbf{e}_{k}\right\|\neq 0$, this term is only determined by the total iteration number $K$ and the stepsize $\{\alpha_{k}\}$.
-It is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$ and can be calculated when $\{\alpha_{k}\}$ is fixed.
-Remark 3.4.
-The inexactness assumption and corresponding analysis in [49] can also handle unbounded feasible regions.
-However, their inexactness assumption for convex problems requires the gradient estimate to be a subgradient, which is not a requirement in our work.
-Furthermore, their inexactness level along iterations has to be the same.
-Finally, in their FGM analysis, the accumulated error depends on the initial condition $R$.
-In contrast, our approach provides a pure separation of the convergence rate with the exact gradient oracle and the accumulated error.
-Our error term is independent of $R$ and holds for any arbitrary error along iterations.
-3.3 PEP-inspired convergence analysis of iGFGM
+$=$ $\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})}.$
+If each $u_{k}$ is large enough, in this case,
+$u_{k}\geq\frac{A_{k}(1+2\alpha_{k+1})(A_{k}+2\alpha_{k}\alpha_{k+1})}{4LA_{K}(A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K-1}\frac{A_{i}(1+2\alpha_{i+1})\alpha_{k}\alpha_{i+1}}{2LA_{K}(A_{i+1}-\alpha_{i+1}^{2})},$
+then the coefficients of $\left\|\mathbf{e}_{k}\right\|^{2}$ will be negative and we can derive the convergence result of iGOGM as
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{LR^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}.$
+The selection of the lower bound of $u_{k}$ is equivalent to establishing conditions for $\mathbf{S}$ in (18) to be a diagonally dominated matrix.
+Inserting these results into the original inequality, we get the final bound.
+∎
 Inspired by the approach for the analysis of iGOGM, we also find a convergence rate of inexact Generalized FGM (iGFGM) with the algorithm shown in Algorithm 1 in Section A.2.
 Similar to the result of iGOGM, the bound of iGFGM is a summation of the exact convergence rate and accumulated error.
-Theorem 3.5.
-The sequence generated by Algorithm 6 satisfies
-$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{2A_{K}}+\sum_{k=0}^{K-1}u_{k}\left\|\mathbf{e}_{k}\right\|^{2},$
-with $u_{k}=\frac{A_{k}^{2}(1+\alpha_{k+1})}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2LA_{K}(2A_{i}-\alpha_{i}^{2})}$.
-4 Exploiting the rate-error tradeoff and finding the optimal inexactness schedule
-In Theorem 3.3, we provided the convergence bound of iGOGM with the explicit expression for the accumulated error.
-This section first looks into the tradeoff between the convergence rate and accumulated error through the stepsize with a fixed gradient inexactness level along iterations.
-Second, we aim to optimize the inexactness levels across iterations, for a given stepsize, so that the total effort to obtain the gradient oracles is minimized.
-4.1 Convergence rate and the accumulated error tradeoff
-Considering the convergence bound $\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}}{4A_{K}}+\sum_{k=0}^{K-1}u_{k}b_{k}^{2}$, with $\{A_{k}\},\{u_{k}\}$ being functions of $\{\alpha_{k}\}$, one can propose to minimize the bound with respect to $\{\alpha_{k}\}$.
-However, we note that such a problem is a high-order polynomial optimization and its solution is intractable in general.
-Even if one is satisfied with the numerical solution to find the optimized algorithm, then it is preferable to solve the minimax problem (9) following the relaxation approach in [22] which we also used in Appendix B.
-This procedure generates the optimized stepsizes and the resulting convergence bound is better than the solution of the polynomial optimization problem as the derived theoretical $\{u_{i}\}$ in Lemma 3.2 is not necessarily the optimal solution.
-Here we narrow down the iGOGM to iOGM-a, with the stepsize $\alpha_{k}=\frac{i+a}{a}$ and $A_{k}=\sum_{i=0}^{k}\frac{i+a}{a}=\frac{(k+2a)(k+1)}{2a}$.
-The condition $A_{k}>\alpha_{k}^{2}$ holds for any $k$ when $a>2$.
-Note that when $K\gg a$, the convergence rate is still $\mathcal{O}(K^{-2})$ in the exact case.
-Furthermore, [42] shows that the exact algorithm OGM-a has the asymptotic worst-case bound for function value as $\frac{a}{2}K^{-2}$ and the smallest gradient norm as $\frac{a\sqrt{6}}{2\sqrt{a-2}}K^{-1.5}$; it achieves the best performance of these two measures simultaneously when $a=4$.
-Both the simplicity of the stepsize and the good performances over function value and gradient norm indicate that OGM-a is a good selection under exact oracles and is worth further study for the inexact oracles too.
-Replacing the stepsizes of OGM-$a$ in Theorem 3.3, we get the coefficients of the error as
-$u_{k}=\frac{1}{4La(K+1)(K+2a)((a-2)k+2a^{2}-2a-2)}\left((2a+8)k^{4}+(11a^{2}+48a+16)k^{3}\right.$
-$\left.+(20a^{3}+107a^{2}+70a+8)k^{2}+(12a^{4}+104a^{3}+100a^{2}+24a)k+36a^{4}+48a^{3}+16a^{2})\right)$
-$+\sum_{i=k+1}^{K-1}\frac{(k+a)(2i^{3}+(9a+4)i^{2}+(13a^{2}+13a+2)i+6a^{3}+10a^{2}+4a}{La(K+1)(K+2a)((a-2)i+2a^{2}-2a-2)},$
-which can be further simplified when $a$ is assigned with a value.
-Take $a=4$ as an example, which has both best convergence rates for function value and the gradient norm under exact oracle, the $u_{k}$ is
-$u_{k}=$ $\frac{16k^{4}+384k^{3}+3280k^{2}+11424k+12544}{16L(K+1)(K+8)(2k+22)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(2i^{3}+40i^{2}+262i+560)}{4L(K+8)(K+1)(2i+22)}$
-$=$ $\frac{(k+7)^{2}(k+8)(k+2)}{2L(K+8)(K+1)(k+11)}+\sum_{i=k+1}^{K-1}\frac{(k+4)(i+8)(i+7)(i+5)}{4L(K+1)(K+8)(i+11)}.$
-Replacing $k+11$ with $k+8$ and $i+11$ with $i+8$ in the two denominators, we have a simpler expression with slightly bigger $u_{k}$ as
-$u_{k}=\frac{(k+7)^{2}(k+2)}{2L(K+8)(K+1)}+\frac{(k+4)(K-k-1)(2K^{2}+2Kk+35K+2k^{2}+37k+210)}{24L(K+8)(K+1)}.$
-Assume $b_{k}\equiv b$, then the accumulated error is calculated as
-$b^{2}\sum_{k=0}^{K-1}u_{k}=\frac{b^{2}K(12K^{3}+303K^{2}+2687K+8758)}{480L(K+8)}.$
-The above results indicate that the accumulated error is in $\mathcal{O}(\frac{b^{2}K^{3}}{L})$.
-For any arbitrary $a$ values, it is difficult to derive a closed form of $u_{k}$ since the denominator can only be simplified as $k+c$ structure when $a$ is given as a specific number.
-Hence, we provide a numerical result of $u_{k}$ for given $K$ with different $a$ in Figure 4.
-For the same $a$ and $k$, bigger $K$ results in bigger $u_{k}$.
-For all $a$ values, $\{u_{k}\}$ first increases and then decreases.
-For a fixed $k$ and $K$, bigger $a$ results in smaller $u_{k}$ values.
-Figure 4: Values of $u_{k}$ for OGM-$a$ with different values of $a$ and $K$.
-Figure 5: Convergence rate and accumulated error for OGM-$a$ for different $a$ and $K$ values.
-Based on this result, to reduce the accumulated error, one can consider increasing $a$.
-However, it will adversely slow down the convergence rate as $A_{K}=\frac{(K+2a)(K+1)}{2a}$ and decreases when $a$ is increased.
-Under this circumstance, we provide the comparison of convergence rate and accumulated error with different iteration numbers $K$, and $a$ in Figure 5, where $\tau=\frac{L}{4A_{K}}$.
-The convergence rate $\tau$ is smaller when $a$ is smaller while the accumulated error is bigger.
-When $a=4$, $\sum_{k=0}^{K-1}u_{k}$ is between $K^{2}$ and $K^{3}$, better than our relaxed analysis result.
-Another interesting fact is that when $K\ll a$ ($a=10^{5}$ and $a=10^{6}$), the convergence rate $\tau=\mathcal{O}(K^{-1})$ and the accumulate error is also $\mathcal{O}(K^{1})$.
-If $K\ll a$, we have $\alpha_{k}=1$, and the proposition below provides the convergence bound of this extreme scenario.
-Proposition 4.1.
-With the stepsize defined as $\alpha_{k}\equiv 1$, the sequence generated by iGOGM (Algorithm 5) satisfies
-$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\sum_{k=0}^{K-1}\frac{3(2K-k+1)}{4L(K+1)}\left\|\mathbf{e}_{k}\right\|^{2}.$
-Furthermore, when $\left\|\mathbf{e}_{k}\right\|\leq\bar{b},\forall k$, the above inequality becomes
-$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}\leq\frac{L\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|}{4(K+1)}+\frac{9K}{8L}\bar{b}^{2}.$
-4.2 Optimal inexactness schedule
-Since the derived $u_{k}$ is not constant, it will be beneficial to set $b_{k}$ based on the value of $u_{k}$.
-Recall the definition of $\mathbb{\eta}$ from (2).
-When $r=1$, we can aim to optimize over the total $\eta$-complexity, i.e., $\sum_{k=0}^{K-1}\eta_{k}$, so that the accumulated error will not deteriorate the convergence rate.
-The problem is given as
-[…]
-$\operatorname{s.t.}\quad$ $\sum_{k=0}^{K-1}u_{k}b_{k}^{2}\leq\frac{LR^{2}}{4A_{K}}.$
-This problem is a convex optimization when $h$ is convex and decreasing over linear constraint and at its optimal solution, the constraint is binding.
-Hence, the inequality constraint can be replaced with the equality one.
-Let $\lambda$ be the dual multiplier of the constraint, the Lagrangian function is
-$\mathcal{L}(\mathbf{b},\lambda)=\sum_{k=0}^{K-1}h^{-1}(b_{k})+\lambda\left(\sum_{k=0}^{K-1}u_{k}b_{k}^{2}-\frac{LR^{2}}{4A_{k}}\right).$
-At the optimal solution $\mathbf{b}^{*}$ and $\lambda^{*}$, from the optimality conditions, we have
-[…]
-[…]
-Solving the above equations, we have a closed-form solution of $\mathbf{b}^{*}$.
-In the following, we give the closed-form of $\mathbf{b}^{*}$ for two examples of $h(\eta)$ [47], which are two common structures of the inexact gradient oracles.
-Lemma 4.2 ($h(\eta)$ with power law decay).
-Let $h(\eta)=c_{1}\eta^{-c_{2}}$ with $c_{1},c_{2}>0$ to be two constants. The inverse function and its gradient are
-$h^{-1}(b)=\left(\frac{c_{1}}{b}\right)^{1/c_{2}}\quad\text{and}\quad\nabla_{b}h^{-1}(b)=-\frac{c_{1}^{1/c_{2}}b^{-(1+c_{2})/c_{2}}}{c_{2}}.$
-Furthermore, the optimal primal and dual solutions of (OPT-b) are
-$b_{k}^{*}=\frac{\sqrt{L}R}{2\sqrt{A_{K}\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}}u_{k}^{d/(1+2c_{2})}},\quad\lambda^{*}=\frac{c^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
-Proof.
-With power law decay, (13) is written as
-$-\frac{c^{1/c_{2}}(b_{k}^{*})^{-(1+c_{2})/c_{2}}}{c_{2}}+2\lambda^{*}b_{k}^{*}u_{k}=0,\quad k=0,1,\cdots,K-1.$
-From this equation, $b_{k}^{*}$ can be calculated as
-[…]
-Inserting this equation into (13), and solve for $\lambda^{*}$ we get
-$\lambda^{*}=\frac{c_{1}^{1/c_{2}}}{2c_{2}}\left(\frac{LR^{2}}{4A_{K}}\right)^{-(1+2c_{2})/(2c_{2})}\left(\sum_{k=0}^{K-1}u_{k}^{1/(1+2c_{2})}\right)^{(1+2c_{2})/(2c_{2})}.$
-Replacing this solution into (15), the final result can be derived. ∎
-Lemma 4.3 ($h(\eta)$ with exponential decay).
-Let $h(\eta)=q_{1}q_{2}^{-\eta}$ with $q_{1}>0$ and $q_{2}>1$ be two constants. Then
-$h^{-1}(b)=\frac{\log q_{1}-\log b}{\log q_{2}},\quad\nabla_{b}h^{-1}(b)=-\frac{1}{b\log q_{2}},$
-and, the optimal primal and dual solutions are
-$b_{k}^{*}=\sqrt{\frac{LR^{2}}{4KA_{K}u_{k}}},\quad\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
-Proof.
-Similar to the proof of Lemma 4.2, from (13) we have
-$u_{k}(b_{k}^{*})^{2}=\frac{1}{2\lambda^{*}\log q_{2}}.$ (16)
-Inserting this into (14),
-$\lambda^{*}=\frac{2KA_{K}}{LR^{2}\log q_{2}}.$
-By the previous two equations, we can derive the final solution. ∎
-To quantify the improvement after the optimized inexactness level, we calculate the total $\eta$-complexity for iOGM-4 algorithm for the two structures of $h(\eta)$ functions.
-We fix $L=1$ and $R=1$, then solve the constant $b$ through
-$\bar{b}^{2}\sum_{k=0}^{K-1}\leq\frac{LR^{2}}{4A_{K}},$
-and calculate the optimized $b$ by the solutions we derived in Lemmas 4.2 and 4.3.
-In Figure 6, we report the result for power law decay with four different values of $c_{2}$.
-Note that the $\eta$-complexity decreases with increased $c_{2}$.
-If we focus on the improvement of $\eta$-complexity for the optimized $\{b_{k}\}$, the improvement is in the same order as the complexity of constant $b$, which shows optimizing over $b$ can decrease the $\eta$-complexity significantly.
-As for the exponential decay in Figure 7, the improvement is not as large as that of the power law decay.
-Although the improvement is not as large as that of the power law decay, considering its absolute value, it still represents a considerable improvement.
-Figure 6: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with power law decay with $c_{1}=1$ and $L=1,R=1$.
-Figure 7: Total $\eta$-complexity of iOGM-4 for $h(\eta)$ with exponential decay with $q_{1}=1,q_{2}=e$ and $L=1,R=1$.
+Remark 4.2.
+Note that Theorem 2.2 is valid only for $0<\alpha_{k}^{2}<A_{k}$, i.e., $\alpha_{k}^{2}$ cannot be equal to $A_{k}$, which reflects a limitation of the analytically feasible solution we derived.
+However, as demonstrated in Theorem 2.4, for iFGM, we can set $\alpha_{k}^{2}=A_{k}$.
+This particular choice allows iFGM to attain its fastest convergence rate if the oracle is exact.
 5 Conclusions
-In this paper, we analyze the Generalized Optimal Gradient Method (GOGM) with the inexact gradient oracle under the absolute error assumption.
-By utilizing the Performance Estimation Problem (PEP) analysis tool, a new upper bound for the convergence rate of GOGM is derived.
-This bound is composed of two independent components, the convergence rate obtained from the exact gradient oracle and the accumulated error resulting from the gradient inexactness.
+In this paper, we analyze the Generalized Optimized Gradient Method (GOGM) with the inexact gradient oracle, i.e., iGOGM, under the absolute error assumption.
+By utilizing the Performance Estimation Problem (PEP) analysis tool, a new upper bound for the convergence rate of the iGOGM algorithm is derived.
+This bound comprises two components: the convergence rate obtained from the exact gradient oracle and the accumulated error resulting from the gradient oracle’s inexactness.
 Such a bound demonstrates the effect of the inexactness and does not require boundedness of the feasible region.
-Furthermore, this bound is independent of the initial condition.
-From this convergence bound, the optimal schedule to set oracle inexactness along iterations is proposed.
+Furthermore, the accumulated error component of the bound is independent of the initial condition.
+Furthermore, from this convergence bound, the optimal schedule to set the oracle inexactness along iterations is proposed.
 Such a study aims to minimize the sampling or computational effort for gradient estimation while maintaining the convergence rate.
 Appendix A Supplementary lemmas and proofs
-A.1 Proof of composition optimization
+A.1 Details on motivational examples
+Bilevel optimization.
+Consider the bilevel optimization problem
+$\min_{\mathbf{x}}$ $\quad f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
+$\operatorname{s.t.}$ $\quad\mathbf{y}^{*}(\mathbf{x})=\operatorname*{argmin}_{\mathbf{y}}g(\mathbf{x},\mathbf{y}),$
+with $f$ being continuously differentiable and $g$ being two times continuously differentiable and strongly convex functions.
+To solve the problem with a gradient-based method, we first obtain the gradient with respect to $\mathbf{x}$ as
+$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$ $=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))$
+$=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla\mathbf{y}^{*}(\mathbf{x})^{\top}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x})).$
+This requires solving the lower-level problem to optimality, which is unattainable in many scenarios.
+When the lower-level problem is solved to a suboptimal point $\tilde{\mathbf{y}}(\mathbf{x})$ and the gradient is evaluated at this point instead of $\mathbf{y}^{*}(\mathbf{x})$, the resulting gradient is inexact:
+$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\neq\tilde{\nabla}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))\triangleq\nabla_{\mathbf{x}}f(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\tilde{\mathbf{y}}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x})).$
+Under some continuity assumptions, the gradient error is shown to be bounded as
+$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left\|\tilde{\mathbf{y}}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|,$
+where $C\triangleq L_{f_{\mathbf{x}}}+\frac{L_{f_{\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}}+C_{f_{\mathbf{y}}}\left(\frac{L_{g_{\mathbf{x}}\mathbf{y}}}{\mu_{g}}+\frac{L_{g_{\mathbf{y}\mathbf{y}}}C_{g_{\mathbf{x}\mathbf{y}}}}{\mu_{g}^{2}}\right)$, with $L_{f_{\mathbf{x}}},L_{f_{\mathbf{y}}}$ are the Lipschitz continuity constants of $\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y})$ and $\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})$, respectively; $L_{g_{\mathbf{x}\mathbf{y}}},L_{g_{\mathbf{y}\mathbf{y}}}$ are the Lipschitz constants of $\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}),\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y})$ with respect to (w.r.t.) $\mathbf{y}$, respectively; $C_{f_{\mathbf{y}}}$, $C_{g_{\mathbf{x}\mathbf{y}}}$ are the upper bounds on $\left\|\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y})\right\|$ and $\left\|\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y})\right\|$, respectively; and $\mu_{g}$ is the strong convexity constant of $g(\mathbf{x},\mathbf{y})$ w.r.t.
+$\mathbf{y}$– see [24] for details.
+Now let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations.
+From the iteration complexity of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
+$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
+where $Q_{g}\triangleq L_{g}/\mu_{g}$. The above bound follows an exponential decay and represents the error bound in (2) with $\eta$ being the iteration number $k$.
+Composition optimization.
+Consider the nested composition optimization
+$\min_{\mathbf{x}}\ f(\mathbf{x})\triangleq h(\mathbf{g}(\mathbf{x}))\quad\text{where}\quad\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{N}\mathbf{g}_{i}(\mathbf{x}),$
+where $h:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\mathbf{g}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ are continuously differentiable, and $N$ is a large positive integer.
+From the chain rule, the exact gradient is $\nabla f(\mathbf{x})=\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))$.
+However, assume one can only approximate $\mathbf{g}$ and $\nabla\mathbf{g}$ through sampling as
+$\tilde{\mathbf{g}}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\mathbf{g}}\right|}\mathbf{g}_{\mathcal{S}_{\mathbf{g}}[i]}(\mathbf{x})\quad\tilde{\nabla}\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{\left|\mathcal{S}_{\nabla}\right|}\nabla\mathbf{g}_{\mathcal{S}_{\nabla}[i]}(\mathbf{x}),$
+where $\mathcal{S}_{\mathbf{g}}$, $\mathcal{S}_{\nabla}$ are two subsets of $\{1,2,\cdots,N\}$.
+Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{h}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$.
+Furthermore, assume $h(\mathbf{v})$ is Lipschitz smooth with constant $L_{h}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$, and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$.
+Then with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
+$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right).$
+The proof of the above bound is presented in Lemma A.1.
+This bound represents the error bound in (2) with $\boldsymbol{\eta}=(\left|\mathcal{S}_{\nabla}\right|,\left|\mathcal{S}_{\mathbf{g}}\right|)$.
+Furthermore, if $\mathcal{S}_{\nabla}=\mathcal{S}_{\mathbf{g}}$, for any fixed $\epsilon$, the bound follows a power law decay.
 Lemma A.1.
-Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla f(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$.
-Also, $f(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$ and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$.
-Assume $\left|\mathcal{S}_{\nabla}\right|$ and $\left|\mathcal{S}_{\mathbf{g}}\right|$ are properly chosen so $\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\leq 1$, then the error of the estimated gradient is bounded as
-~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq b^{2},$
-with probability at least $1-\frac{2C_{f}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}$.
+Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{h}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$.
+Also, $f(\mathbf{v})$ is Lipschitz smooth with constant $L_{h}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$ and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$.
+Assume $\left|\mathcal{S}_{\nabla}\right|$ and $\left|\mathcal{S}_{\mathbf{g}}\right|$ are properly chosen so $\frac{2C_{h}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\leq 1$, then the error of the estimated gradient is bounded as
+$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq b^{2},$
+with probability at least $1-\frac{2C_{h}^{2}}{b^{2}\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{b^{2}\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}$.
 Proof.
 By Young’s inequality,
-~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
-$=$ ~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))+\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
-$\leq$ $2\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
-$\leq$ $2\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\right\|^{2}\left\|\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}$
-$\leq$ $2C_{f}^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2C_{g}^{2}L_{f}^{2}\left\|\tilde{\mathbf{g}}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}.$
+$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
+$=$ $\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))+\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
+$\leq$ $2\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
+$\leq$ $2\left\|\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))\right\|^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2\left\|\nabla\mathbf{g}(\mathbf{x})\right\|^{2}\left\|\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}$
+$\leq$ $2C_{h}^{2}\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}+2C_{g}^{2}L_{h}^{2}\left\|\tilde{\mathbf{g}}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}.$
 Taking the expectation on both sides and with the bounded variance assumptions, we have
-$\mathbb{E}[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}]$
-$\leq$ $\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}.$
+$\mathbb{E}[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}]$
+$\leq$ $\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}.$
 By Markov’s inequality, we have
-$\Pr\left(\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\geq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)\right)$
-$\leq$ $\frac{1}{\epsilon}\mathbb{E}\left[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\right]/\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
+$\Pr\left(\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\geq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)\right)$
+$\leq$ $\frac{1}{\epsilon}\mathbb{E}\left[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\right]/\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
 $\leq$ $\frac{1}{\epsilon}.$
 Assuming $\epsilon>1$, i.e., the bound holds non-triviality, we can conclude that
-~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
+$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
 holds with probability at least $1-\frac{1}{\epsilon}$.
-A.2 Convergence rate for inexact generalized FGM
-Algorithm 6 Inexact Generalized Fast Gradient Method (iGFGM)
-0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
-1:  for $k=0,\cdots,K-1$ do
-2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
-3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{1}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
-4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
-5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
-6:  end for
+A.2 Convergence bound of iGFGM
 Lemma A.2.
@@
 Similar to the first step of the proof of Theorem 2.2,
-$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
-$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
-$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
-$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
-$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
-$=$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
-$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$+\frac{1}{L}\sum_{k=1}^{K}\alpha_{k}\alpha_{k-1}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k-1})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle$
-$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
-$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
-$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
-$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\frac{2A_{0}-\alpha_{0}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{0})\right\|^{2}$
-$-\sum_{k=0}^{K-1}\frac{1}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|(2A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+(\alpha_{k}\alpha_{k+1}+A_{k})\mathbf{e}_{k}\right\|^{2}$
-$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}^{2}\alpha_{i}^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
-$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=0}^{i-1}\frac{\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle$
-$\leq$ $\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
-$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
+$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$\leq$ $-\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
+$-\frac{1}{2L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
+$=$ $\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}+A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
+$=$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau A_{K}^{2}}-\frac{A_{k}}{LA_{K}}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\frac{\alpha_{k}\alpha_{i}}{A_{K}^{2}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
+$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$+\frac{1}{L}\sum_{k=1}^{K}\frac{\alpha_{k}\alpha_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k-1})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle$
+$-\frac{1}{LA_{K}}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
+$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
+$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}+A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
+$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\frac{2A_{0}-\alpha_{0}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{0})\right\|^{2}$
+$-\sum_{k=0}^{K-1}\frac{1}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|(2A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+(\alpha_{k}\alpha_{k+1}+A_{k})\mathbf{e}_{k}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}^{2}\alpha_{i}^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=0}^{i-1}\frac{\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle$
+$\leq$ $\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
+$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
 Following the same steps in the proof of iGOGM, define
-$\hat{P}_{k,k}\triangleq$ $\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}$
-$\hat{P}_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k+1}\alpha_{k}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})},$
+$\hat{P}_{k,k}\triangleq$ $\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}$
+$\hat{P}_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k+1}\alpha_{k}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})},$
 we can bound the error terms as
@@
 $\leq$ $\sum_{k=0}^{K-1}\left(\hat{P}_{k,k}+\frac{1}{2}\sum_{i=0}^{k-1}\hat{P}_{k,i}+\frac{1}{2}\sum_{i=k+1}^{K-1}\hat{P}_{i,k}\right)\left\|\mathbf{e}_{k}\right\|^{2}$
-$=$ $\sum_{k=0}^{K-1}\left(\frac{A_{k}^{2}(1+\alpha_{k+1})}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2L(2A_{i}-\alpha_{i}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
-Dividing both sides of the inequality by $A_{K}$, the final result is obtained. ∎
+$=$ $\sum_{k=0}^{K-1}\left(\frac{A_{k}^{2}(1+\alpha_{k+1})}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2LA_{K}(2A_{i}-\alpha_{i}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
 Appendix B Numerical solutions for optimized algorithm with inexact oracle
@@
 5 $\begin{pmatrix}0.0111&0&0&0&0\\ 0.0014&0.5086&0&0&0\\ 0.0005&0.4713&1.6643&0&0\\ 0.0002&0.4588&1.8883&1.8380&0\\ 0.0000&0.4531&1.9905&2.2202&1.8669\end{pmatrix}$
-Table 3: Numerical solution of optimized first-order algorithm with $b_{i}$ decreasing
+Table 4: Numerical solution of the optimized first-order algorithm with $b_{i}$ decreasing
 We can observe that the stepsize at each iteration with the exact gradient oracle remains the same while this is not the case with the inexact oracle.